### v1

# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

Abstract.
Let $G$ be a $p$-adic analytic pro-$p$ group of dimension $d$, with lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. We produce an approximate series which descends regularly in strata and whose terms deviate from $P_{i}(G)$ in a uniformly bounded way. This brings to light a new set of rational invariants $\xi_{1},\ldots,\xi_{d}\in[\nicefrac{{1}}{{d}},1]$, canonically associated to $G$, such that
$\sup_{i\in\mathbb{N}}\;\Bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\Big{(}\sum_{k=1}^{d}\xi_{k}\Big{)}(i-1)\Bigr{\rvert}<\infty$
and such that the Hausdorff dimensions $\operatorname{hdim}_{G}^{\mathcal{L}}(H)$ of closed subgroups $H\leq_{\mathrm{c}}G$ with respect to $\mathcal{L}$ are confined to the set
$\Big{\{}\Big{(}\sum_{k=1}^{d}\varepsilon_{k}\xi_{k}\Big{)}/\Big{(}\sum_{k=1}^{d}\xi_{k}\Big{)}\mid\varepsilon_{1},\ldots,\varepsilon_{d}\in\{0,1\}\Big{\}}.$
In particular, the Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)$ is discrete and consists of at most $2^{d}$ rational numbers.
Furthermore, we show that, if $G$ is a countably based pro-$p$ group with an open subgroup mapping onto the free abelian pro-$p$ group $\mathbb{Z}_{p}\oplus\mathbb{Z}_{p}$, then for every prescribed finite set $\{0,1\}\subseteq X\subseteq[0,1]$ there is a filtration series $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(G)=X$; in particular, $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert$ is unbounded, as $\mathcal{S}$ runs through all filtration series of $G$ with $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert<\infty$.
Finally, we establish that finitely generated non-abelian direct products $G$ of free pro-$p$ groups have full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)=[0,1]$ with respect to the lower $p$-series $\mathcal{L}$.
Key words and phrases:
Lower $p$-series, $p$-adic analytic groups, Hausdorff dimension, Hausdorff spectrum, free pro-$p$ groups
2010 Mathematics Subject Classification:
Primary 20E18; Secondary 22E20, 28A78
Funding acknowledgements. The first author was supported by the Spanish Government, grant MTM2017-86802-P, partly with FEDER funds, and by the Basque Government, grant IT974-16; he received funding from the European Union’s Horizon 2021 research and innovation programme under the Marie Sklodowska-Curie grant agreement, project 101067088. The second author received support from the Knut and Alice Wallenberg Foundation. The third author acknowledges support from EPSRC, grant EP/T005068/1 and from the Folke Lannér’s Fund. The research was partially conducted in the framework of the DFG-funded research training group GRK 2240: Algebro-Geometric Methods in Algebra, Arithmetic and Topology.
1. Introduction
Throughout $p$ denotes a prime, and $\mathbb{Z}_{p}$ is the ring of $p$-adic integers. We are interested in the lower $p$-series of $p$-adic analytic pro-$p$ groups and questions about Hausdorff dimension with respect to the lower $p$-series. To put our results into context, we also consider specific situations involving more general pro-$p$ groups and other filtration series.
Let $G$ be a finitely generated pro-$p$ group. The lower $p$-series of $G$ – sometimes called the lower $p$-central series – is defined recursively as follows:
$P_{1}(G)=G,$ $\quad\text{and}\quad P_{i}(G)=\overline{P_{i-1}(G)^{p}\,[P_{i-1}(G),G]}\quad\text{for $i\in\mathbb{N}$ with $i\geq 2$,}$
where we adopt the notation from [7, Sec. 1.2]. It is the fastest descending central series of $G$ with elementary abelian sections and it consists of open normal subgroups forming a base of neighbourhoods for the element $1$. The lower $p$-series plays a significant role in the study of pro-$p$ groups. For instance, $\bigoplus_{i=1}^{\infty}P_{i}(G)/P_{i+1}(G)$ naturally carries the structure of an $\mathbb{F}_{p}$-Lie algebra and can frequently be used to analyse the structure of $G$ or its finite quotients; compare with [16, Chap. VIII], [24, Chap. II] and [23, Sec. 5].
In order to understand the lower $p$-series of $p$-adic analytic pro-$p$ groups, it is useful to consider an analogous series for $\mathbb{Z}_{p}$-lattices furnished with a group action. Let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a (continuous) right $G$-action, in other words, let $L$ be a right $\mathbb{Z}_{p}G$-module that is finitely generated and torsion-free as a $\mathbb{Z}_{p}$-module. The two-sided ideal
$\mathfrak{a}_{G}=\sum_{g\in G}(g-1)\mathbb{Z}_{p}G+p\mathbb{Z}_{p}G\,\trianglelefteq\,\mathbb{Z}_{p}G$
is the kernel of the natural epimorphism from $\mathbb{Z}_{p}G$ onto the finite field $\mathbb{F}_{p}$ that sends each group element to $1$; it is convenient to refer to $\mathfrak{a}_{G}$, somewhat informally, as the $p$-augmentation ideal of the group ring $\mathbb{Z}_{p}G$. The powers $\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the $p$-augmentation ideal intersect in $\{0\}$ and induce a topology on $\mathbb{Z}_{p}G$ which gives rise to the completed group ring $\mathbb{Z}_{p}[\![G]\!]$; see [7, Sec. 7.1]. The descending series of open $\mathbb{Z}_{p}G$-submodules
$\lambda_{i}(L)=L.\mathfrak{a}_{G}^{\,i}\quad\text{for $i\in\mathbb{N}_{0}$,}$ (1.1)
is the fastest descending $G$-central series of $L$ with elementary abelian sections. Accordingly we call the series (1.1) the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$, taking note of the notational shift in the index in comparison to the lower $p$-series of a group. As $G$ is a pro-$p$ group, it acts unipotently on every principal congruence quotient $L/p^{j}L$, $j\in\mathbb{N}$, and this implies that $\bigcap_{i\in\mathbb{N}_{0}}\lambda_{i}(L)=\{0\}$.
In general, a filtration series of $L$, regarded as a $\mathbb{Z}_{p}$-lattice, is a descending chain $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ of open $\mathbb{Z}_{p}$-sublattices $L_{i}\leq_{\mathrm{o}}L$ such that $\bigcap_{i\in\mathbb{N}_{0}}L_{i}=\{0\}$. For $c\in\mathbb{N}_{0}$, we say that two filtration series $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and $\mathcal{S}^{*}\colon L=L_{0}^{*}\geq L_{1}^{*}\geq\ldots$ are $c$-equivalent if for all $i\in\mathbb{N}$,
$p^{c}L_{i}\subseteq L_{i}^{*}\quad\text{and}\quad p^{c}L_{i}^{*}\subseteq L_{i}.$
We say that $\mathcal{S}$ and $\mathcal{S}^{*}$ are equivalent if they are $c$-equivalent for some $c\in\mathbb{N}_{0}$. This is but one of several ways to establish the concept or a closely related one. For instance, if $\log_{p}\lvert L_{i}+L_{i}^{*}:L_{i}\cap L_{i}^{*}\rvert\leq c$ for all $i\in\mathbb{N}$, then $\mathcal{S}$ and $\mathcal{S}^{*}$ are $c$-equivalent; conversely, if $\mathcal{S}$ and $\mathcal{S}^{*}$ are $c$-equivalent, then $\log_{p}\lvert L_{i}+L_{i}^{*}:L_{i}\cap L_{i}^{*}\rvert\leq 2c$ for all $i\in\mathbb{N}$. Thus $\mathcal{S}$ and $\mathcal{S}^{*}$ are equivalent if and only if $\lvert L_{i}+L_{i}^{*}:L_{i}\cap L_{i}^{*}\rvert$ is uniformly bounded as $i$ ranges over all positive integers.
Our aim is to show that the lower $p$-series of $L$ is equivalent to a filtration series that is descending very regularly, in a sense that we make precise now. Let $d=\dim_{\mathbb{Z}_{p}}L$ denote the dimension of the $\mathbb{Z}_{p}$-lattice $L$, and let $x_{1},\ldots,x_{d}$ be a $\mathbb{Z}_{p}$-basis for $L$. We say that a filtration series $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ of $L$ has a (vertical) stratification given by the frame $(x_{1},\ldots,x_{d})$ if there are positive real parameters $\xi_{1},\ldots,\xi_{d}$ with $\xi_{1}\leq\ldots\leq\xi_{d}$ such that $\mathcal{S}$ is equivalent to the filtration series
$\mathcal{S}^{*}\colon L_{i}^{*}=\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k};$
in this situation, $(\xi_{1},\ldots,\xi_{d})$ is uniquely determined by the frame and we refer to this $d$-tuple as the growth rate of the stratification. Informally, we call $\xi_{1},\ldots,\xi_{d}$ the ‘growth rates’ of the stratification. We say that the stratification $\mathcal{S}^{*}$ is given by $G$-invariant sublattices if all terms $L_{i}^{*}$ are $G$-invariant.
Theorem 1.1.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a $G$-action. Then the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification, and any two stratifications have the same growth rate.
Moreover, if $d=\dim_{\mathbb{Z}_{p}}L$ and if $(x_{1},\ldots,x_{d})$ is a frame for a stratification of the lower $p$-series of $L$ with growth rate $(\xi_{1},\ldots,\xi_{d})$, then
  1. (i)
    $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$ are rational numbers,
  2. (ii)
    for every $e\in\{1,\ldots,d-1\}$ with $\xi_{e}<\xi_{e+1}$ the sublattice $\bigoplus_{k=1}^{e}\mathbb{Z}_{p}x_{k}$ is $G$-invariant, and
  3. (iii)
    the entire stratification is given by $G$-invariant sublattices.
The above result yields an analogous description of the lower $p$-series of a $p$-adic analytic pro-$p$ group. We recall that, for every $p$-adic analytic pro-$p$ group $G$, there exists $j\in\mathbb{N}$ such that for every $i\in\mathbb{N}$ with $i\geq j$ the group $P_{i}(G)$ is uniformly powerful; see [7, Prop. 3.9 and Thm. 4.2]. Moreover, every minimal generating system $x_{1},\ldots,x_{d}$ of a uniformly powerful pro-$p$ group $U$ yields a product decomposition of $U=\overline{\langle x_{1}\rangle}\cdots\overline{\langle x_{d}\rangle}$ into $d$ procyclic subgroups, which in turn gives a system of $p$-adic coordinates $U\to\mathbb{Z}_{p}^{\,d}$, $x_{1}^{\,a_{1}}\cdots x_{d}^{\,a_{d}}\mapsto(a_{1},\ldots,a_{d})$ of ‘the second kind’; see [7, Thm. 4.9]. The dimension of $G$, denoted by $\dim(G)$, coincides with the rank of any uniformly powerful open subgroup of $G$.
Theorem 1.2.
Let $G$ be a $p$-adic analytic pro-$p$ group of dimension $d$. Then there exists $j\in\mathbb{N}$ such that
  1. (i)
    the group $U=P_{j}(G)\trianglelefteq_{\mathrm{o}}G$ is uniformly powerful and
  2. (ii)
    there is a minimal generating system $x_{1},\ldots,x_{d}$ for $U$ and there are rational numbers $\xi_{1},\ldots,\xi_{d}$ with $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$ and $c\in\mathbb{N}_{0}$ such that the following holds: for all $i\in\mathbb{N}$ with $i\geq\max\{j+2cd,6cd\}$, the following products of procyclic subgroups
    $\check{U}_{i}$ $=\overline{\langle x_{1}^{\,p^{\lfloor(i-j)\xi_{1}\rfloor+c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{\,p^{\lfloor(i-j)\xi_{d}\rfloor+c}}\rangle}$
    $\hat{U}_{i}$ $=\overline{\langle x_{1}^{\,p^{\lfloor(i-j)\xi_{1}\rfloor-c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{\,p^{\lfloor(i-j)\xi_{d}\rfloor-c}}\rangle}$
    form open normal subgroups $\check{U}_{i},\hat{U}_{i}\trianglelefteq_{\mathrm{o}}G$ such that
    $\check{U}_{i}\subseteq P_{i}(G)\subseteq\hat{U}_{i}$
    and the factor group $\hat{U}_{i}/\check{U}_{i}$ is homocyclic abelian of exponent $p^{2c}$.
Corollary 1.3.
In the set-up of Theorem 1.2 and with the additional notation $\sigma=\sum_{k=1}^{d}\xi_{k}\in[1,d]$, the index of $P_{i}(G)$ in $G$, for $i\geq\max\{j+2cd,6cd\}$, satisfies
$\Bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\big{(}\log_{p}\lvert G:P_{j}(G)\rvert+\lfloor(i-j)\sigma\rfloor\big{)}\Bigr{\rvert}\leq d(c+1).$
Corollary 1.4.
Let $G$ be a $p$-adic analytic pro-$p$ group of dimension $d$. Then any two approximations of the lower $p$-series as described in Theorem 1.2 involve the same rational growth parameters $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$.
The last corollary underlines that we have uncovered a new set of invariants which capture certain structural properties of $p$-adic analytic pro-$p$ groups.
One of our motivations for pinning down the lower $p$-series of $p$-adic analytic pro-$p$ groups was to settle some basic, but annoyingly outstanding questions about the Hausdorff spectra of such groups. The study and application of Hausdorff dimension in profinite groups is an ongoing affair; for instance, see [22] for an overview and [21, 9, 11, 13, 14, 15] for some recent developments. Let $G$ be an infinite finitely generated pro-$p$ group and let $\mathcal{S}\colon G_{i},\,i\in\mathbb{N}_{0}$, be a filtration series of $G$, that is, a descending chain $G=G_{0}\geq G_{1}\geq\ldots$ of open normal subgroups $G_{i}\trianglelefteq_{\mathrm{o}}G$ such that $\bigcap_{i\in\mathbb{N}_{0}}G_{i}=1$. In accordance with [1, 5], the Hausdorff dimension of a closed subgroup $H\leq_{\mathrm{c}}G$, with respect to $\mathcal{S}$, admits an algebraic interpretation as a logarithmic density:
[…]
and the Hausdorff spectrum of $G$, with respect to $\mathcal{S}$,
$\operatorname{hspec}^{\mathcal{S}}(G)=\{\operatorname{hdim}_{G}^{\mathcal{S}}(H)\mid H\leq_{\mathrm{c}}G\}\subseteq[0,1]$
provides a simple ‘snapshot’ of the typically rather complicated array of all subgroups in $G$. We say that $H$ has strong Hausdorff dimension in $G$, with respect to $\mathcal{S}$, if
$\operatorname{hdim}^{\mathcal{S}}_{G}(H)=\lim_{i\to\infty}\frac{\log_{p}\lvert HG_{i}:G_{i}\rvert}{\log_{p}\lvert G:G_{i}\rvert}$
is given by a proper limit.
Even for well-behaved groups, such as $p$-adic analytic pro-$p$ groups $G$, not only the Hausdorff dimension of individual subgroups, but also qualitative features of the Hausdorff spectrum of $G$ are known to be sensitive to the choice of $\mathcal{S}$; see [22, Thm. 1.3]. In our setting, where $G$ is a finitely generated pro-$p$ group, the group-theoretically relevant choices of $\mathcal{S}$ are primarily: the $p$-power series $\mathcal{P}$, the iterated $p$-power series $\mathcal{P}^{*}$, the Frattini series $\mathcal{F}$, the dimension subgroup (or Zassenhaus) series $\mathcal{D}$ and, last but not least, the lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. It was shown in [22, Prop. 1.5] that, if $G$ is a $p$-adic analytic pro-$p$ group, then the Hausdorff dimensions of $H\leq_{\mathrm{c}}G$ with respect to the series $\mathcal{P},\mathcal{F},\mathcal{D}$ coincide, and [5, Thm. 1.1] yields, for $\mathcal{S}$ any one of these series $\mathcal{P},\mathcal{F},\mathcal{D}$, that the spectrum
$\operatorname{hspec}^{\mathcal{S}}(G)=\Big{\{}\nicefrac{{\displaystyle\dim(H)}}{{\displaystyle\dim(G)}}\mid H\leq_{\mathrm{c}}G\Big{\}}$
is discrete and that all closed subgroups have strong Hausdorff dimension in $G$ with respect to $\mathcal{S}$. These results extend easily to Hausdorff dimensions with respect to the series $\mathcal{P}^{*}$.
It was seen in [22] that, contrary to indications in [5], the Hausdorff spectra of $p$-adic analytic pro-$p$ groups $G$ with respect to the lower $p$-series $\mathcal{L}$ follow a different pattern. Concretely, [22, Example 4.1] provides a family of $p$-adic analytic pro-$p$ groups $G$ such that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert/\dim(G)$ is unbounded as $\dim(G)\to\infty$. It remained open whether $\operatorname{hspec}^{\mathcal{L}}(G)$ could even be infinite for some $p$-adic analytic groups $G$. Theorem 1.1 allows us to establish that $\operatorname{hspec}^{\mathcal{L}}(G)$ is finite for every $p$-adic analytic pro-$p$ group $G$ and to put in place further restrictions.
Theorem 1.5.
Let $G$ be an infinite $p$-adic analytic pro-$p$ group and let $d=\dim(G)$. Then there exist $\xi_{1},\ldots,\xi_{d}\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$ such that
$\operatorname{hspec}^{\mathcal{L}}(G)\subseteq\left\{\frac{\varepsilon_{1}\xi_{1}+\ldots+\varepsilon_{d}\xi_{d}}{\xi_{1}+\ldots+\xi_{d}}\mid\varepsilon_{1},\ldots,\varepsilon_{d}\in\{0,1\}\right\}.$
In particular, this shows that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert\leq 2^{d}$. The parameters $\xi_{1},\ldots,\xi_{d}$ are the same as those appearing in Theorem 1.2, and all closed subgroups have strong Hausdorff dimension in $G$ with respect to $\mathcal{L}$.
It is a natural problem to improve, if possible, upon the upper bound for $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert$ as a function of $d=\dim(G)$. In Example 3.2 we manufacture a family of $p$-adic analytic pro-$p$ groups $G_{m}$, $m\in\mathbb{N}$, such that
$\frac{\log_{p}\bigl{\lvert}\operatorname{hspec}^{\mathcal{L}}(G_{m})\bigr{\rvert}}{\sqrt{\dim(G_{m})}}\to\infty\qquad\text{as $m\to\infty$.}$
Theorem 1.5 triggers more refined questions about Hausdorff spectra with respect to the lower $p$-series. Under additional assumptions on $G$, we obtain the following stronger results, which partly extend even to other filtration series.
Theorem 1.6.
Let $G$ be an infinite $p$-adic analytic pro-$p$ group.
If $G$ is just infinite, then every closed subgroup $H\leq_{\mathrm{c}}G$ has strong Hausdorff dimension $\operatorname{hdim}_{G}^{\mathcal{S}}(H)=\dim(H)/\dim(G)$ for every filtration series $\mathcal{S}$.
If $G$ is nilpotent, then every closed subgroup $H\leq_{\mathrm{c}}G$ has strong Hausdorff dimension $\operatorname{hdim}_{G}^{\mathcal{S}}(H)=\dim(H)/\dim(G)$ for $\mathcal{S}=\mathcal{L}$.
In the cases discussed above, the Hausdorff spectrum of $G$ with respect to the relevant filtration series is equal to $\Big{\{}\nicefrac{{\displaystyle\dim(H)}}{{\displaystyle\dim(G)}}\mid H\leq_{\mathrm{c}}G\Big{\}}$.
It remains an outstanding problem in the subject to understand, more generally, what renders the Hausdorff spectrum $\operatorname{hspec}^{\mathcal{S}}(G)$ of a finitely generated pro-$p$ group $G$ finite. In particular, it is open whether finiteness of the spectrum with respect to one of the standard filtration series $\mathcal{P},\mathcal{P}^{*},\mathcal{F},\mathcal{D},\mathcal{L}$ implies that $G$ is $p$-adic analytic; compare with [22, Prob. 1.4]. It is known that for every $p$-adic analytic pro-$p$ group $G$ whose abelianisation has torsion-free rank at least $2$, there is a filtration $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(G)$ is infinite; see [22, Thm. 1.3]. This leads to the question [22, Prob. 1.6]: does there exist, for every $p$-adic analytic pro-$p$ group $G$, a uniform bound $b(G)$ for $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert$, as $\mathcal{S}$ runs through all filtration series of $G$ with $\lvert\operatorname{hspec}^{\mathcal{S}}(G)\rvert<\infty$. We show that the answer is typically negative, in the strongest possible way.
Theorem 1.7.
Let $G$ be a countably based pro-$p$ group that has an open subgroup mapping surjectively onto $\mathbb{Z}_{p}\oplus\mathbb{Z}_{p}$. Then for every finite subset $X\subseteq[0,1]$ with $\{0,1\}\subseteq X$ there exists a filtration series $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(G)=X$.
Irrespective of whether or not we expect the finiteness of the Hausdorff spectrum to characterise $p$-adic analytic pro-$p$ groups, it is of interest to produce and study potential counterexamples to such a proposition. For the filtration series $\mathcal{F}$ and $\mathcal{D}$, it was shown in [9] that free pro-$p$ groups $F$ can be cancelled from the candidate list; in fact, non-abelian finite direct products $G$ of free pro-$p$ groups have full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{F}}(G)=\operatorname{hspec}^{\mathcal{D}}(G)=[0,1]$. Using Lie-theoretic tools, we produce an analogous result for the lower $p$-series $\mathcal{L}$.
Theorem 1.8.
Let $G=F_{1}\times\ldots\times F_{r}$ be a finite direct product of finitely generated free pro-$p$ groups $F_{j}$, at least one of which is non-abelian. Then $G$ has full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)=[0,1]$ with respect to the lower $p$-series $\mathcal{L}$.
It remains a challenge to compute the Hausdorff spectra with respect to the lower $p$-series for pro-$p$ groups of positive rank gradient, such as non-soluble Demushkin pro-$p$ groups. This would further complement the results in [9].
Notation. Generally, any new notation is explained when needed, sometimes implicitly. For convenience we collect basic conventions here.
[…]
The group-theoretic notation is mostly standard and in line, for instance, with its use in [7]. Tacitly, subgroups of profinite groups are generally understood to be closed subgroups. In some places, we emphasise that we are taking the topological closure of a set $X$ by writing $\overline{X}$.
Organisation. In Section 2, we study stratifications of the lower $p$-series and establish Theorems 1.1 and 1.2, as well as their corollaries. In Section 3 we turn to Hausdorff spectra with respect to the lower $p$-series: using the results from the previous section, we prove Theorems 1.5 and 1.6, for $p$-adic analytic pro-$p$ groups. In Section 4 we establish Theorem 1.7, yielding arbitrarily large finite Hausdorff spectra, and Theorem 1.8 about full Hausdorff spectra for finite direct products of free pro-$p$ groups.
2. Stratifications of the lower $p$-series
Let $G$ be a profinite group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. We are interested in situations, where the lower $p$-series $\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, of $L$, as defined in (1.1), satisfies $\bigcap_{i\in\mathbb{N}_{0}}\lambda_{i}(L)=\{0\}$ and hence gives a filtration series. This is the case if and only if $G$ acts unipotently on every principal congruence quotient $L/p^{j}L$, $j\in\mathbb{N}$, which in turn is true if and only if the action of $G$ on the $\mathbb{Z}_{p}$-lattice $L\cong\mathbb{Z}_{p}^{\,d}$ factors through a pro-$p$ subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$. Since every closed subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$ is topologically finitely generated, we may restrict without any true loss of generality to the situation that $G$ itself is a finitely generated pro-$p$ group. A theorem of Serre implies that every abstract action of the group $G$ on the group $L$ is continuous; see [7, Sec. 1]. Thus we arrive at the standard set-up which we use henceforth: $L$ is a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action, for a finitely generated pro-$p$ group $G$.
We start with a basic observation about growth rates of stratifications.
Lemma 2.1.
Let $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ be a filtration series of a $\mathbb{Z}_{p}$-lattice $L$ which admits a stratification. Then any two stratifications of $\mathcal{S}$ have the same growth rate.
Proof.
Let $d=\dim_{\mathbb{Z}_{p}}L$ and, for $i\in\mathbb{N}$, let $p^{m_{i,1}}$, …, $p^{m_{i,d}}$ with $0\leq m_{i,1}\leq\ldots\leq m_{i,d}$ be the elementary divisors of the finite $\mathbb{Z}_{p}$-module $L/L_{i}$ so that
$L/L_{i}\cong\mathbb{Z}_{p}/p^{m_{i,1}}\mathbb{Z}_{p}\,\oplus\ldots\oplus\,\mathbb{Z}_{p}/p^{m_{i,d}}\mathbb{Z}_{p}.$
Let $(x_{1},\ldots,x_{d})$ be the frame for a stratification of $\mathcal{S}$, with growth rate $(\xi_{1},\ldots,\xi_{d})$, say. Then there is a constant $c\in\mathbb{N}$ such that $\mathcal{S}$ is $c$-equivalent to the filtration series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Recall that $\xi_{1}\leq\ldots\leq\xi_{d}$ and choose $j\in\mathbb{N}$ such that $\lfloor j\xi_{e}\rfloor+c<\lfloor j\xi_{e+1}\rfloor-c$ for all $e\in\{1,\ldots,d-1\}$ with $\xi_{e}<\xi_{e+1}$. This implies that for all $i\in\mathbb{N}$ with $i\geq j$,
$\lvert m_{i,k}-\lfloor i\xi_{k}\rfloor\rvert\leq c\qquad\text{for $1\leq k\leq d$.}$
Thus each $\xi_{k}=\lim_{i\to\infty}m_{i,k}/i$ is uniquely determined by $\mathcal{S}$. ∎
Next we recall and adapt the following definitions from [19]: the lower and upper level of an open $\mathbb{Z}_{p}G$-submodule $M$ in $L$ are
$\ell_{L}(M)=\min\{k\in\mathbb{N}_{0}\mid p^{k}L\subseteq M\}\;\text{and}\;u_{L}(M)=\max\{k\in\mathbb{N}_{0}\mid M\subseteq p^{k}L\},$
and the rigidity of $L$ is
$r(L)=\sup\{\ell_{L}(M)-u_{L}(M)\mid\text{$M$ is an open $\mathbb{Z}_{p}G$-submodule of $L$}\}.$
The following auxiliary result is a variation of [19, Prop. 4.3] and can be proved in the same way, using compactness.
Proposition 2.2.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Then $r(L)$ is finite if and only if $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}G$-module.
In conjunction with Lemma 2.1, the next proposition already establishes a special case of Theorem 1.1.
Proposition 2.3.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Suppose that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}G$-module. Then the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification with constant growth rate $(\xi,\ldots,\xi)$, where $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$ for $d=\dim_{\mathbb{Z}_{p}}L$.
Proof.
Denote by $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, the terms of the lower $p$-series of $L$, and for $i\in\mathbb{N}_{0}$ put $u_{i}=u_{L}(L_{i})$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity $r=r(L)$. Consequently, the lower $p$-series of $L$ is equivalent to the filtration series $\mathcal{S}\colon L\geq p^{u_{1}}L\geq p^{u_{2}}L\geq\ldots$, and it suffices to show that $\mathcal{S}$ is equivalent to a series of the form $p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$, for suitable $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$.
[…]
$\varphi_{i}\colon p^{u_{i}}L\to p^{u_{i}}L/p^{u_{i}+r}L\cong R$
defined by the dictum $p^{u_{i}}x_{k}\mapsto z_{k}$ for $1\leq k\leq d$. For each $i\in\mathbb{N}$, we have $\ell_{L}(L_{i})\leq u_{i}+r$, hence $p^{u_{i}+r}L\leq L_{i}\leq p^{u_{i}}L$, and we consider $R_{i}=L_{i}\varphi_{i}$.
Since $R$ is finite, there exist $j,m\in\mathbb{N}$ such that $R_{j+m}=R_{j}$, and therefore $L_{j+m}=p^{n}L_{j}$ for $n=u_{j+m}-u_{j}\geq 1$. Furthermore $p^{m}L_{j}\subseteq L_{j+m}$ yields $n\leq m$. We see that $L_{i+m}=L_{j+m}.\mathfrak{a}_{\mathbb{Z}_{p}G}^{\,i-j}=p^{n}L_{j}.\mathfrak{a}_{\mathbb{Z}_{p}G}^{\,i-j}=p^{n}L_{i}$ for all $i\in\mathbb{N}$ with $i\geq j$. In particular, this shows that $u_{i+m}=u_{i}+n$ for $i\geq j$. Put $\xi=n/m\in(0,1]\cap\mathbb{Q}$ and
$c=1+\max\big{\{}\lvert u_{i}-i\xi\rvert\bigm{|}0\leq i\leq j+m-1\big{\}}.$
Then
$\lvert u_{i}-\lfloor i\xi\rfloor\rvert\leq\lvert u_{i}-i\xi\rvert+1\leq c\qquad\text{for all $i\in\mathbb{N}_{0}$,}$
and the filtration series $\mathcal{S}$ is $c$-equivalent to $\mathcal{S}^{*}\colon p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$. Finally, we notice that $L=L_{0}\supsetneqq L_{1}\supsetneqq\ldots$ yields the trivial lower bound $\lvert L:L_{i}\rvert\geq p^{i}$ and hence
$\xi=\lim_{i\to\infty}\frac{d\lfloor i\xi\rfloor}{di}=\lim_{i\to\infty}\frac{\log_{p}\lvert L:p^{\lfloor i\xi\rfloor}L\rvert}{di}=\lim_{i\to\infty}\frac{\log_{p}\lvert L:L_{i}\rvert}{di}\geq\nicefrac{{1}}{{d}}.\qed$
The next proposition provides a key tool for proving Theorem 1.1; with its corollary it yields, in particular, the last two assertions in the theorem.
Proposition 2.4.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Suppose that the lower $p$-series of $L$ admits a stratification, with the frame $(x_{1},\ldots,x_{d})$ and growth rate $(\xi_{1},\ldots,\xi_{d})$.
Suppose that $e\in\{1,\ldots,d-1\}$ is such that $\xi_{e}<\xi_{e+1}$. Then the $\mathbb{Z}_{p}$-lattice $M=\bigoplus_{k=1}^{e}\mathbb{Z}_{p}x_{k}$ is $G$-invariant. Moreover the lower $p$-series of $M$ has a stratification with frame $(x_{1},\ldots,x_{e})$ and growth rate $(\xi_{1},\ldots,\xi_{e})$.
Proof.
The $\mathbb{Z}_{p}$-lattice $L$ decomposes as the direct sum of its sublattices $M$ and $N=\bigoplus_{k=e+1}^{d}\mathbb{Z}_{p}x_{k}$. Choose $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Then
$L_{i}\cap N\subseteq p^{\lfloor i\xi_{e+1}\rfloor-c_{1}}N\qquad\text{for all $i\in\mathbb{N}_{0}$,}$ (2.1)
where we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret negative exponents occurring for small values of $i$.
For a contradiction, assume that $M$ is not $G$-invariant, i.e. not a $\mathbb{Z}_{p}G$-submodule. Then there are elements $\alpha_{1},\ldots,\alpha_{e}\in\mathbb{Z}_{p}G$ and $c_{2}\in\mathbb{N}$ such that $y=\sum_{k=1}^{e}x_{k}\alpha_{k}\in N\smallsetminus p^{c_{2}}N$. For every $i\in\mathbb{N}$, we have $p^{\lfloor i\xi_{e}\rfloor+c_{1}}M\subseteq L_{i}$ and hence $p^{\lfloor i\xi_{e}\rfloor+c_{1}}y\in(L_{i}\cap N)\smallsetminus p^{\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}}N$. Thus (2.1) yields $\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}>\lfloor i\xi_{e+1}\rfloor-c_{1}$. Since $\xi_{e}<\xi_{e+1}$ this yields the desired contradiction for sufficiently large $i\in\mathbb{N}$.
It remains to prove that the lower $p$-series $M_{i}=\lambda_{i}(M)$, $i\in\mathbb{N}_{0}$, of $M$ has a stratification with the frame $(x_{1},\ldots,x_{e})$ and growth rate $(\xi_{1},\ldots,\xi_{e})$. As before we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret any negative powers of $p$, whose occurrence does not cause any serious problems. We put
$b=c_{1}+b_{1}\qquad\text{with}\qquad b_{1}=\left\lceil\frac{3c_{1}+3}{\xi_{e+1}-\xi_{e}}\right\rceil.$
It is clear that, for all $i\in\mathbb{N}_{0}$,
$M_{i}\subseteq L_{i}\cap M\subseteq\left(\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}\,x_{k}\right)\cap M\subseteq p^{-b}\,\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k},$
and it suffices to show that, for all $i\in\mathbb{N}_{0}$ with $i\geq b_{1}$,
$p^{b}\,\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}\subseteq M_{i}.$
Let $i\in\mathbb{N}_{0}$ with $i\geq b_{1}$. From $p\in\mathfrak{a}_{G}$ it follows that $p^{b_{1}}M_{i-b_{1}}\subseteq M_{i-b_{1}}.\mathfrak{a}_{G}^{\,b_{1}}=M_{i}$, and it suffices to establish that $p^{c_{1}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}\subseteq M_{i-b_{1}}$. There is no harm in working modulo $p\big{(}p^{c_{1}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}\big{)}$ and hence it is enough to show that
$\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}}\mathbb{Z}_{p}\,x_{k}\subseteq M_{i-b_{1}}+\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}+1}\mathbb{Z}_{p}\,x_{k}.$
Recall that $L=M\oplus N$ as a $\mathbb{Z}_{p}$-lattice and let $\tau\colon L\to M$ denote the projection along $N$. Recall further that $\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}}\mathbb{Z}_{p}\,x_{k}\subseteq L_{i}$. Hence it is enough to show that
$L_{i}\tau\subseteq M_{i-b_{1}}+\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}+1}\mathbb{Z}_{p}\,x_{k}.$ (2.2)
From $L_{b_{1}}\subseteq M+p^{\lfloor b_{1}\xi_{e+1}\rfloor-c_{1}}N$ we conclude that
$L_{i}\tau=\big{(}L_{b_{1}}.\mathfrak{a}_{G}^{\,i-b_{1}}\big{)}\tau\subseteq\big{(}\underbrace{M.\mathfrak{a}_{G}^{\,i-b_{1}}}_{=M_{i-b_{1}}}+p^{\lfloor b_{1}\xi_{e+1}\rfloor-c_{1}}\underbrace{N.\mathfrak{a}_{G}^{\,i-b_{1}}}_{\subseteq L_{i-b_{1}}}\big{)}\tau,$
and hence $L_{i-b_{1}}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor(i-b_{1})\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}x_{k}$ gives
$L_{i}\tau$ $\subseteq M_{i-b_{1}}+\big{(}p^{\lfloor b_{1}\xi_{e+1}\rfloor-c_{1}}L_{i-b_{1}}\big{)}\tau$
$\subseteq M_{i-b_{1}}+\bigoplus_{k=1}^{e}p^{\lfloor b_{1}\xi_{e+1}\rfloor+\lfloor(i-b_{1})\xi_{k}\rfloor-2c_{1}}\mathbb{Z}_{p}\,x_{k}.$
Now our choice of $b_{1}$ guarantees that, for each $k\in\{1,\ldots,e\}$,
$\lfloor b_{1}\xi_{e+1}\rfloor+\lfloor(i-b_{1})\xi_{k}\rfloor-2c_{1}\geq b_{1}\xi_{e+1}+(i-b_{1})\xi_{k}-2c_{1}-2\\ =i\xi_{k}+c_{1}+b_{1}(\xi_{e+1}-\xi_{k})-(3c_{1}+2)\geq\lfloor i\xi_{k}\rfloor+c_{1}+1,$
and this yields (2.2). ∎
Corollary 2.5.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Then every stratification of the lower $p$-series of $L$ is given by $G$-invariant sublattices.
Proof.
Suppose that the lower $p$-series of $L$ admits a stratification, with the frame $(x_{1},\ldots,x_{d})$ and growth rate $(\xi_{1},\ldots,\xi_{d})$. There are $\vartheta_{1}<\ldots<\vartheta_{r}$, for some $r\in\{1,\ldots,d\}$, such that $\{\xi_{k}\mid 1\leq k\leq d\}=\{\vartheta_{s}\mid 1\leq s\leq r\}$. For $1\leq s\leq r$ let $d_{s}$ denote the multiplicity of $\vartheta_{s}$ in $(\xi_{1},\ldots,\xi_{d})$, and for $0\leq s\leq r$ put $D(s)=d_{1}+\ldots+d_{s}$. Then the $\mathbb{Z}_{p}$-lattice $L$ decomposes as
$L=N_{1}\oplus\ldots\oplus N_{r},\quad\text{where}\quad N_{s}=\bigoplus_{k=D(s-1)+1}^{D(s)}\mathbb{Z}_{p}x_{k}\quad\text{for $1\leq s\leq r$,}$
and our stratification of the lower $p$-series of $L$ is realised by the series
$\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}N_{s},\quad i\in\mathbb{N}_{0}.$ (2.3)
For each $s\in\{1,\ldots,r\}$, Proposition 2.4 yields $N_{s}.\mathfrak{a}_{G}\subseteq N_{1}\oplus\ldots\oplus N_{s}$, because the $\mathbb{Z}_{p}$-lattice $N_{1}\oplus\ldots\oplus N_{s}$ is $G$-invariant. Recalling $\vartheta_{1}<\ldots<\vartheta_{r}$ we deduce that for every $i\in\mathbb{N}_{0}$,
$\Big{(}\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}N_{s}\Big{)}.\mathfrak{a}_{G}=\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}\big{(}N_{s}.\mathfrak{a}_{G}\big{)}\\ \subseteq\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}\big{(}N_{1}\oplus\ldots\oplus N_{s}\big{)}=\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}N_{s}.$
Hence all terms of the series (2.3) are $G$-invariant. ∎
Lemma 2.6.
Let $G$ be a finitely generated pro-$p$ group, and let $L,\tilde{L}$ be $\mathbb{Z}_{p}$-lattices equipped with $G$-actions such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L\cong\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}\tilde{L}$ as $\mathbb{Q}_{p}G$-modules. Put $d=\dim_{\mathbb{Z}_{p}}L=\dim_{\mathbb{Z}_{p}}\tilde{L}$. If the lower $p$-series of $L$ admits a stratification with growth rate $(\xi_{1},\ldots,\xi_{d})$ then so does the lower $p$-series of $\tilde{L}$.
Proof.
Without loss of generality, $L$ and $\tilde{L}$ are $G$-invariant full $\mathbb{Z}_{p}$-lattices in one and the same $\mathbb{Q}_{p}G$-module $V=\mathbb{Q}_{p}L=\mathbb{Q}_{p}\tilde{L}$ and $d=\dim_{\mathbb{Q}_{p}}V$. Then there exists $c_{1}\in\mathbb{N}$ such that $\tilde{L}\subseteq p^{-c_{1}}L$. Since $p^{-c_{1}}L$ and $L$ are isomorphic as $\mathbb{Z}_{p}G$-modules, we may assume without loss of generality that $\tilde{L}\subseteq L$.
Furthermore, there is $c_{2}\in\mathbb{N}$ such that $p^{c_{2}}L\subseteq\tilde{L}$. This implies that $p^{c_{2}}\lambda_{i}(L)\subseteq\lambda_{i}(\tilde{L})\subseteq\lambda_{i}(L)$ for all $i\in\mathbb{N}$. Let $(x_{1},\ldots,x_{d})$ be the frame for a stratification of the lower $p$-series of $L$, with growth rate $(\xi_{1},\ldots,\xi_{d})$. Then there is $c_{3}\in\mathbb{N}$ such that, for $i\in\mathbb{N}$,
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor+c_{3}}\mathbb{Z}_{p}x_{k}\subseteq\lambda_{i}(L)\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{3}}\mathbb{Z}_{p}x_{k}.$
Below we produce a $\mathbb{Z}_{p}$-basis $y_{1},\ldots,y_{d}$ for $\tilde{L}$ and $c_{4}\in\mathbb{N}$ such that
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor+c_{4}}\mathbb{Z}_{p}x_{k}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}.$ (2.4)
With $c=c_{2}+c_{3}+c_{4}$, this yields for all $i\in\mathbb{N}_{0}$,
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor+c}\mathbb{Z}_{p}y_{k}\subseteq\lambda_{i}(\tilde{L})\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c}\mathbb{Z}_{p}y_{k}.$
Thus the lower $p$-series of $\tilde{L}$ admits a stratification with frame $(y_{1},\ldots,y_{d})$ and growth rate $(\xi_{1},\ldots,\xi_{d})$.
It remains to establish (2.4) for suitable $y_{1},\ldots,y_{d}$ and $c_{4}$. For each $e\in\{1,\ldots,d\}$ we choose
$y_{e}=a_{e,1}x_{1}+\ldots+a_{e,e-1}x_{e-1}+p^{b_{e}}x_{e}\in\tilde{L},$
where $a_{e,1},\ldots,a_{e,e-1}\in\mathbb{Z}_{p}$ denote suitable coefficients and $b_{e}\in\{0,1,\ldots,c_{2}\}$, in such a way that $\bigoplus_{k=1}^{e}\mathbb{Z}_{p}y_{k}=\tilde{L}\cap\bigoplus_{k=1}^{e}\mathbb{Z}_{p}x_{k}$. We observe that $y_{1},\ldots,y_{d}$ do form a $\mathbb{Z}_{p}$-basis for $\tilde{L}$ and that $b_{1}+\ldots+b_{d}=\log_{p}\lvert L:\tilde{L}\rvert$. Furthermore, we see that for each $i\in\mathbb{N}$ and each $e\in\{1,\ldots,d\}$,
$p^{\lfloor i\xi_{e}\rfloor}y_{e}=\left(\sum_{k=1}^{e-1}\big{(}p^{\lfloor i\xi_{e}\rfloor-\lfloor i\xi_{k}\rfloor}a_{e,k}\big{)}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}\big{)}\right)+p^{b_{e}}\big{(}p^{\lfloor i\xi_{e}\rfloor}x_{e}\big{)}.$
This implies that
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\leq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$
with
$\log_{p}\left\lvert\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}:\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\right\rvert=b_{1}+\ldots+b_{d}=\log_{p}\lvert L:\tilde{L}\rvert.$
Hence (2.4) holds for $c_{4}=\log_{p}\lvert L:\tilde{L}\rvert$. ∎
We are now in a position to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a $G$-action, for a finitely generated pro-$p$ group $G$. In view of Lemma 2.1, Proposition 2.4 and Corollary 2.5, it remains to prove that the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification with rational growth rates in $[\nicefrac{{1}}{{d}},1]$, where $d=\dim_{\mathbb{Z}_{p}}L$.
We argue by induction on $d$. Of course, the claim holds trivially if $L=\{0\}$, and we now suppose that $d\geq 1$. Let $N$ be a $\mathbb{Z}_{p}G$-submodule of $L$ such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Choose $n\in\mathbb{N}$ such that $p^{n+1}L\cap N\subseteq pN$. Then $\tilde{L}=p^{n}L+N$ is a $G$-invariant open $\mathbb{Z}_{p}$-sublattice of $L$, and $N$ is a direct summand of the $\mathbb{Z}_{p}$-lattice $\tilde{L}$. By Lemma 2.6, there is no harm in replacing $L$ with $\tilde{L}$ and we may thus suppose that $L$ itself decomposes as $L=M\oplus N$ for a $\mathbb{Z}_{p}$-sublattice $M$.
Suppose that we can arrange for $M$ to be $G$-invariant. Then $L$ is a direct sum $M\oplus N$ of $\mathbb{Z}_{p}G$-modules and its lower $p$-series decomposes accordingly: $\lambda_{i}(L)=\lambda_{i}(M)\oplus\lambda_{i}(N)$ for $i\in\mathbb{N}_{0}$. If both $\dim_{\mathbb{Z}_{p}}M$ and $\dim_{\mathbb{Z}_{p}}N$ are less than $d$, our claim follows by induction, while in the remaining case, $L=N$ and $M=\{0\}$, the proof can be finished by appealing to Proposition 2.3.
Now suppose that neither $M$ nor any other $\mathbb{Z}_{p}$-sublattice $M^{\prime}$ of $L$, which complements $N$ to yield an open sublattice $M^{\prime}\oplus N$ of $L$, is $G$-invariant. Let us fix some additional notation. We write
$L_{i}=\lambda_{i}(L),\quad i\in\mathbb{N}_{0},$
for the lower $p$-series. Put $e=\dim_{\mathbb{Z}_{p}}M$. By Proposition 2.3 there is $\eta\in[\nicefrac{{1}}{{(d-e)}},1]\cap\mathbb{Q}$ such that the lower $p$-series of $N$ is equivalent to the series $p^{\lfloor i\eta\rfloor}N$, $i\in\mathbb{N}_{0}$. By induction on the dimension, the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L/N$, which is $(L_{i}+N)/N$, $i\in\mathbb{N}$, admits a stratification, based on a frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$, with growth rates $\xi_{1}\leq\ldots\leq\xi_{e}$ in $[\nicefrac{{1}}{{e}},1]\cap\mathbb{Q}$. Set $\xi_{0}=0$ for notational convenience.
We choose representatives $x_{1},\ldots,x_{e}\in M$ for the cosets $\mathbf{x}_{1},\ldots,\mathbf{x}_{e}\in L/N$, and pick $\tilde{e}\in\{0,1,\ldots,e-1\}$ as large as possible subject to the conditions:
[…]
In fact, we arrange that among all possible choices – for the complement $M$ for $N$ in $L$ (initially up to finite index), for the stratification based on the frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$ and for the representatives $x_{1},\ldots,x_{e}$ in $L$ – we make one that renders $\tilde{e}$ as large as possible. In addition we choose $\hat{e}\in\{\tilde{e}+1,\ldots,e\}$ maximal with $\xi_{\hat{e}}=\xi_{\tilde{e}+1}$, and for short we put
$\vartheta=\xi_{\tilde{e}+1}=\xi_{\tilde{e}+2}=\ldots=\xi_{\hat{e}}\qquad\text{and}\qquad T=\bigoplus_{k=\tilde{e}+1}^{\hat{e}}\mathbb{Z}_{p}x_{k}.$
Our motivation here is that, by Proposition 2.4, the $\mathbb{Z}_{p}$-lattice $\tilde{M}\oplus T\oplus N$ is $G$-invariant, even though $\tilde{M}\oplus T$ is not.
We proceed to establish three auxiliary claims. As on previous occasions we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret terms involving negative powers of $p$ and lying outside $L$.
Claim 1. There are elements $w_{i,k}\in N$, for $i\in\mathbb{N}_{0}$ and $\tilde{e}<k\leq e$, and $n_{i}\in\mathbb{N}$ with $n_{i}\leq\lfloor i\eta\rfloor$, for $i\in\mathbb{N}_{0}$, such that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0}.$ (2.5)
Indeed, our construction already incorporates a stratification of the lower $p$-series modulo $N$ and accordingly there exists $c_{1}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{1}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq L_{i}+N\quad\text{and}\quad p^{c_{1}}L_{i}\subseteq\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;N.$
Moreover, the $G$-invariant $\mathbb{Z}_{p}$-lattice $\tilde{M}=\bigoplus_{k=1}^{\tilde{e}}\mathbb{Z}_{p}x_{k}$ embeds isomorphically as $(\tilde{M}+N)/N$ into $L/N$. By Proposition 2.4, applied to $L/N$, there exists $c_{2}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{2}}\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq\lambda_{i}(\tilde{M})\subseteq L_{i}.$
We put $c_{3}=\max\{c_{1},c_{2}\}$ and observe that, a fortiori, for all $i\in\mathbb{N}_{0}$,
$p^{c_{3}}p^{\lfloor i\xi_{k}\rfloor}x_{k}\in L_{i}\qquad\text{for $1\leq k\leq\tilde{e}$.}$
Choose $j\in\mathbb{N}$ such that $L_{j}\subseteq p^{c_{3}}L$. Then for all $i\in\mathbb{N}_{0}$ with $i\geq j$,
$p^{c_{3}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq(L_{i}+N)\cap p^{c_{3}}L=L_{i}+(N\cap p^{c_{3}}L)=L_{i}+p^{c_{3}}N$
allows us to select $w_{i,\tilde{e}+1},\ldots,w_{i,e}\in N$ such that
$p^{c_{3}}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\in L_{i}\qquad\text{for $\tilde{e}<k\leq e$.}$
For $i\in\mathbb{N}_{0}$ with $i<j$ we set $w_{i,\tilde{e}+1}=\ldots=w_{i,e}=0$; small values of $i$ play an insignificant role and this simple choice has no relevance beyond the fact that the series in (2.5) is properly defined. To streamline the notation, we also set
$w_{i,1}=\ldots=w_{i,\tilde{e}}=0\qquad\text{for all $i\in\mathbb{N}_{0}$.}$ (2.6)
For $i\in\mathbb{N}_{0}$ with $i\geq j$ we deduce directly from our definitions that
$p^{c_{3}}\Big{(}\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;(L_{i}\cap N)\Big{)}\subseteq L_{i},$
and next we observe that
$p^{c_{1}}L_{i}\subseteq\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;N=\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;N$
implies
$p^{c_{1}+c_{3}}L_{i}=p^{c_{3}}\big{(}p^{c_{1}}L_{i}\big{)}\subseteq\Big{(}\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\underbrace{p^{c_{3}}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}}_{\in L_{i}}\Big{)}\;\oplus\;(L_{i}\cap N).$
We have established that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the filtration series
$\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;(L_{i}\cap N),\quad i\in\mathbb{N}_{0}.$
By Proposition 2.2, the filtration series $L_{i}\cap N$, $i\in\mathbb{N}_{0}$, of $N$ is equivalent to a filtration series of the form $p^{n_{i}}N$, $i\in\mathbb{N}_{0}$, and our choice of $\eta$, in connection with the lower $p$-series of $N$, yields $c_{4}\in\mathbb{N}$ such that $p^{\lfloor i\eta\rfloor+c_{4}}N\subseteq\lambda_{i}(N)\subseteq L_{i}$ for $i\in\mathbb{N}_{0}$. This allows us to arrange that $n_{i}\in\mathbb{N}$ satisfies $n_{i}\leq\lfloor i\eta\rfloor$ for all $i\in\mathbb{N}_{0}$. Recalling the notational ‘nullnummer’ (2.6), we deduce that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the filtration series in (2.5), as claimed. $\diamond$
Claim 2. There exists $b\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$L_{i}\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N.$ (2.7)
Clearly, the $\mathbb{Z}_{p}$-lattice $L$ decomposes as a direct sum
$L=\tilde{M}\oplus K\oplus N$
of its sublattices $\tilde{M}$, $K=\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}x_{k}$ and $N$. Furthermore, the lattices $\tilde{M}$ and $N$ are $G$-invariant, whereas $K$ is not.
Recall that $\mathfrak{a}_{G}\trianglelefteq\mathbb{Z}_{p}G$ denotes the $p$-augmentation ideal and yields the description $L_{i}=L.\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the lower $p$-series of $L$, and similarly for $N$. According to our set-up, there exists $c_{1}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{1}}\big{(}K.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{c_{1}}L_{i}\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor}K\oplus N\qquad\text{and}\qquad p^{c_{1}}\big{(}N.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{\lfloor i\eta\rfloor}N.$
By decomposing elements into their $\tilde{M}$-, $K$- and $N$-coordinates and tracing their ‘trajectory’ as we descend along the lower $p$-series, we can make the inclusion recorded above more precise. We choose to work modulo $\tilde{M}$ and can thus focus on the $K$- and $N$-coordinates. Of significance are the $K$-coordinates, because they can give rise to non-trivial $K$- and $N$-coordinates at the next step. We introduce a parameter $j$, which indicates that we follow the $K$-coordinates for $j$ steps, then switch to $N$-coordinates and continue to follow the latter for the remaining $i-j$ steps. Concretely, we deduce that for all $i\in\mathbb{N}_{0}$,
$K.\mathfrak{a}_{G}^{\,i}$ $\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor-c_{1}}N.\mathfrak{a}_{G}^{\,i-1-j}$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor+\lfloor(i-1-j)\eta\rfloor-2c_{1}}N$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus p^{\min\{\lfloor(i-1)\vartheta\rfloor,\lfloor(i-1)\eta\rfloor\}-2c_{1}-2}N.$
This yields, for $i\in\mathbb{N}_{0}$,
$L_{i}=\tilde{M}.\mathfrak{a}_{G}^{\,i}+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\subseteq M+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\\ \subseteq M+K.\mathfrak{a}_{G}^{\,i}+p^{\lfloor i\eta\rfloor-c_{1}}N\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-2c_{1}-3}N.$
Thus (2.7) holds for $b=2c_{1}+3$. $\diamond$
Claim 3. Let $\mathcal{V}=\big{(}p^{-b}N\big{)}^{\hat{e}-\tilde{e}}$, with $b$ as in Claim 2. Then there exists $m\in\mathbb{N}_{0}$ such that for every choice of $\mathbf{v}=(v_{\tilde{e}+1},\ldots,v_{\hat{e}})\in\mathcal{V}$, we have
$p^{m}N\subseteq\tilde{M}+\sum_{k=\tilde{e}+1}^{\hat{e}}(x_{k}+v_{k}).\mathbb{Z}_{p}G.$ (2.8)
For each $\mathbf{v}=(v_{\tilde{e}+1},\ldots,v_{\hat{e}})\in\mathcal{V}$, we think of $\tilde{T}_{\mathbf{v}}=\sum_{k=\tilde{e}+1}^{\hat{e}}\mathbb{Z}_{p}(x_{k}+v_{k})$ as a deformation of $T$ inside $M\oplus p^{-b}N\subseteq\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$. Recall that, by our choice of $\tilde{e}$, the $\mathbb{Z}_{p}$-lattice $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\subseteq L$ is not $G$-invariant, whereas $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\oplus N=\tilde{M}\oplus p^{b}T\oplus N$ is $G$-invariant by Proposition 2.4. This implies that
$\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N\neq\{0\}.$
Recall that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Proposition 2.2 shows that the non-zero $G$-invariant $\mathbb{Z}_{p}$-lattice $\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N$ is open in $N$ of lower level $l_{\mathbf{v}}\in\mathbb{N}_{0}$, say.
Clearly, the function $\mathcal{V}\to\mathbb{N}_{0}$, $\mathbf{v}\mapsto l_{\mathbf{v}}$ is locally constant and, since the domain $\mathcal{V}$ is compact, the maximum $m=\max\{l_{\mathbf{v}}\mid\mathbf{v}\in\mathcal{V}\}$ exists and has the property described in (2.8). $\diamond$
Now we just need to put the pieces together. Combining the assertions of Claim 1 and Claim 2, we see that there exists $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the filtration series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0},$
where
$w_{i,k}\in p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N\quad\text{and}\quad\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b\leq n_{i}\leq\lfloor i\eta\rfloor$
for all relevant indices $i,k$.
First consider the case $\eta\leq\vartheta$. In this situation
$w_{i,k}\in p^{\lfloor i\eta\rfloor-b}N\quad\text{and}\quad\lfloor i\eta\rfloor-b\leq n_{i}\leq\lfloor i\eta\rfloor$
for all relevant indices $i,k$ and it follows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+c_{1})$-equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}p^{\lfloor i\xi_{k}\rfloor}x_{k}\;\oplus\;p^{\lfloor i\eta\rfloor}N,\quad i\in\mathbb{N}_{0},$
We choose any $\mathbb{Z}_{p}$-basis $y_{1},\ldots,y_{d-e}$ for $N$. Then $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to
$\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=1}^{d-e}p^{\lfloor i\eta\rfloor}\mathbb{Z}_{p}y_{k}$
and, if $k\in\{1,\ldots,\tilde{e}\}$ is minimal subject to $\eta\leq\xi_{k}$, this yields a stratification with frame
$(x_{1},x_{2},\ldots,x_{k-1},y_{1},y_{2},\ldots,y_{d-e},x_{k},x_{k+1},\ldots,x_{e})$
and growth rate
$(\xi_{1},\xi_{2},\ldots,\xi_{k-1},\underbrace{\eta,\eta,\ldots,\eta}_{\text{$d-e$ entries}},\xi_{k},\xi_{k+1},\ldots,\xi_{e})\in\big{(}[\nicefrac{{1}}{{\max\{e,d-e\}}},1]\cap\mathbb{Q}\big{)}^{d}.$
Finally, suppose that $\vartheta<\eta$. In this situation
$w_{i,k}=p^{\lfloor i\vartheta\rfloor}v_{i,k}\quad\text{with}\quad v_{i,k}\in p^{-b}N\qquad\text{and}\qquad\lfloor i\vartheta\rfloor-b\leq n_{i}\leq\lfloor i\eta\rfloor.$
Using Claim 3, we deduce that for $i\in\mathbb{N}_{0}$,
$L_{i}\supseteq p^{c_{1}}\Big{(}p^{\lfloor i\vartheta\rfloor}\tilde{M}+\bigoplus_{k=\tilde{e}+1}^{\hat{e}}p^{\lfloor i\vartheta\rfloor}\mathbb{Z}_{p}(x_{k}+v_{i,k})\Big{)}.\mathbb{Z}_{p}G\\ \supseteq p^{c_{1}+\lfloor i\vartheta\rfloor}\Big{(}\tilde{M}+\sum_{k=\tilde{e}+1}^{\hat{e}}(x_{k}+v_{i,k}).\mathbb{Z}_{p}G\Big{)}\supseteq p^{c_{1}+m+\lfloor i\vartheta\rfloor}N$
and hence
$p^{n_{i}}N\supseteq p^{c_{1}}L_{i}\cap N\supseteq p^{2c_{1}+m+\lfloor i\vartheta\rfloor}N.$
It follows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+2c_{1}+m)$-equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;p^{\lfloor i\vartheta\rfloor}N,\quad i\in\mathbb{N}_{0}.$
We choose any $\mathbb{Z}_{p}$-basis $y_{1},\ldots,y_{d-e}$ for $N$. Then $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to
$\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=1}^{d-e}p^{\lfloor i\vartheta\rfloor}\mathbb{Z}_{p}y_{k}$
and this yields a stratification with frame
$(x_{1},x_{2},\ldots,x_{\tilde{e}},y_{1},y_{2},\ldots,y_{d-e},x_{\tilde{e}+1},x_{\tilde{e}+2},\ldots,x_{e})$
and growth rate
$(\xi_{1},\xi_{2},\ldots,\xi_{\tilde{e}},\underbrace{\vartheta,\vartheta,\ldots,\vartheta}_{\text{$d-e$ entries}},\xi_{\tilde{e}+1},\xi_{\tilde{e}+2},\ldots,\xi_{e})\in\big{(}[\nicefrac{{1}}{{\max\{e,d-e\}}},1]\cap\mathbb{Q}\big{)}^{d}.\qed$
Next we use $p$-adic Lie theory to derive Theorem 1.2.
Proof of Theorem 1.2.
Being $p$-adic analytic, the pro-$p$ group $G$ has finite rank. By [7, Prop. 3.9], there exists $j\in\mathbb{N}$ such that $U=P_{j}(G)$ is uniformly powerful of rank $\mathrm{rk}(U)=\dim(G)=d$ and such that $P_{i}(G)$ is powerfully embedded in $U$ for all $i\in\mathbb{N}$ with $i\geq j$. We recall that there is an explicit isomorphism of categories translating between uniformly powerful pro-$p$ groups and powerful $\mathbb{Z}_{p}$-Lie lattices; see [7, Sec. 4 and 9] and [20]. This means that the underlying set of the pro-$p$ group $U$ can be equipped in a canonical way with the structure of a $\mathbb{Z}_{p}$-Lie lattice $L$ carrying the same topology and incorporating essentially all information about $U$. In particular, the construction is such that exponentiation in $U$ corresponds to scalar multiplication in $L$ and that the conjugation action of $G$ on $U$ translates to a continuous $\mathbb{Z}_{p}$-linear action of $G$ on $L$. In order to distinguish between the two parallel points of view, it is convenient to write $\underline{x}$ for $x\in U$ when it features as an element of the Lie lattice $L$ rather than as a group element. Using this notation, the earlier statements can be phrased as follows: the (identity) map $U\to L$, $x\mapsto\underline{x}$ is a homeomorphism; we have $\underline{x^{a}}=a\underline{x}$ for all $x\in U$, $a\in\mathbb{Z}_{p}$; and conjugation, encapsulated in the canonical homomorphism $G\to\operatorname{Inn}(G)$, induces a continuous homomorphism
$G\to\operatorname{Aut}_{\mathbb{Z}_{p}}(L),\quad g\mapsto\alpha_{g},\qquad\text{where $\underline{x}\alpha_{g}=\underline{x^{g}}$ for all $x\in U$, $g\in G$.}$
Furthermore, for every $i\in\mathbb{N}$ with $i\geq j$, the powerfully embedded subgroup $P_{i}(G)\trianglelefteq_{\mathrm{o}}U$ gives, on the Lie lattice side, a powerfully embedded Lie sublattice $L_{i^{*}}\trianglelefteq_{\mathrm{o}}L$, where we write $i^{*}=i-j$ to ease notationally a recurrent constant shift by $j$ in the index; see [20, Thm. 1.4].
Clearly, $j^{*}=0$ gives $L_{j^{*}}=L=\lambda_{j^{*}}(L)$, and we contend that, in fact, $L_{i^{*}}=\lambda_{i^{*}}(L)$ for all $i\geq j$. Let $i\in\mathbb{N}$ with $i>j$. Since $P_{i-1}(G)$ is powerful, its Frattini subgroup equals $P_{i-1}(G)^{p}$ and corresponds to the Lie sublattice $pL_{i^{*}-1}$. Multiplicative cosets of $P_{i-1}(G)^{p}$ in $P_{i-1}(G)$ are the same as additive cosets of $pL_{i^{*}-1}$ in $L_{i^{*}-1}$, and multiplication in $P_{i-1}(G)/P_{i-1}(G)^{p}$ is the same operation as addition $L_{i^{*}-1}/pL_{i^{*}-1}$; see [7, Cor. 4.15]. Since
$\nicefrac{{\displaystyle P_{i-1}(G)}}{{\displaystyle P_{i}(G)}}\cong\nicefrac{{\left(\displaystyle\nicefrac{{\displaystyle P_{i-1}(G)}}{{\displaystyle P_{i-1}(G)^{p}}}\right)}}{{\left(\displaystyle\nicefrac{{\displaystyle P_{i}(G)}}{{\displaystyle P_{i-1}(G)^{p}}}\right)}}$
is the largest $G$-central quotient of the multiplicative group $P_{i-1}(G)/P_{i-1}(G)^{p}$, the corresponding term
$\nicefrac{{\displaystyle L_{i^{*}-1}}}{{\displaystyle L_{i^{*}}}}\cong\nicefrac{{\left(\displaystyle\nicefrac{{\displaystyle L_{i^{*}-1}}}{{\displaystyle pL_{i^{*}-1}}}\right)}}{{\left(\displaystyle\nicefrac{{\displaystyle L_{i^{*}}}}{{\displaystyle pL_{i^{*}-1}}}\right)}}$
is the largest $G$-central quotient of the additive group $L_{i^{*}-1}/pL_{i^{*}-1}$. By induction, $L_{i^{*}-1}$ equals $\lambda_{i^{*}-1}(L)$ and thus its largest $G$-central quotient of exponent $p$ is $L_{i^{*}-1}/\lambda_{i^{*}}(L)$. This implies that $L_{i^{*}}$ is indeed equal to $\lambda_{i^{*}}(L)$.
Theorem 1.1 supplies us with a stratification of the lower $p$-series $L_{i^{*}}$, $i\in\mathbb{N}$ with $i\geq j$, based on a frame $(\underline{x_{1}},\ldots,\underline{x_{d}})$ and with a growth rate $(\xi_{1},\ldots,\xi_{d})$, say, where $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$ are rational numbers. This means in particular that $\underline{x_{1}},\ldots,\underline{x_{d}}$ form a $\mathbb{Z}_{p}$-basis for $L$ and that there exists $c\in\mathbb{N}_{0}$ such that, for all $i\geq j$,
$\check{M}_{i^{*}}\subseteq L_{i^{*}}\subseteq\hat{M}_{i^{*}}$
for
$\check{M}_{i^{*}}=\bigoplus_{k=1}^{d}p^{\lfloor i^{*}\xi_{k}\rfloor+c}\mathbb{Z}_{p}\underline{x_{k}}\quad\text{and}\quad\hat{M}_{i^{*}}=\bigoplus_{k=1}^{d}p^{\lfloor i^{*}\xi_{k}\rfloor-c}\mathbb{Z}_{p}\underline{x_{k}},$
where once more we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to resolve unambiguously multiplication by negative powers of $p$. By Corollary 2.5, the lattices $\check{M}_{i^{*}}$ and $\hat{M}_{i^{*}}$ are $G$-invariant for all $i\geq j$.
We consider $i\in\mathbb{N}$ with $i\geq\max\{j+2cd,6cd\}$. We see that
$[P_{i-2cd}(G),P_{i-2cd}(G)]\subseteq P_{2i-4cd}(G)\subseteq P_{i+2cd}(G).$
In conjunction with [8, Cor. 3.5], the description of the Lie bracket in terms of the group multiplication, as described in [7, Sec. 4] and recorded in [20, Equ. (2.1)] yields the corresponding result on the Lie lattice side,
$[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}}\subseteq L_{i^{*}+2cd},$
where we write $[\cdot,\cdot]_{\mathrm{Lie}}$ for the Lie bracket to distinguish it from group commutators. Similar to the simpler situation discussed above, the finite abelian group section $P_{i-2cd}(G)/P_{i+2cd}(G)$ of $G$ corresponds to a finite abelian Lie ring section $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ of the $\mathbb{Z}_{p}$-Lie lattice $L$: the multiplicative cosets are the same as the additive cosets, multiplication and addition are the same and the Lie bracket is trivial; compare [10, Thm. 4.5].
Indeed, we already observed that the Lie bracket on $L_{i^{*}-2cd}$ is trivial modulo $L_{i^{*}+2cd}$, and the Lie ring $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ is abelian. In conjunction with [20, Lem. B.2], the description of group multiplication via the Hausdorff series $\Phi$, namely
$\underline{xy}=\Phi(\underline{x},\underline{y})=\underline{x}+\underline{y}+\tfrac{1}{2}[\underline{x},\underline{y}]_{\mathrm{Lie}}+\sum_{n=3}^{\infty}u_{n}(\underline{x},\underline{y})\qquad\text{for $x,y\in U$}$
as described in [20], yields that multiplicative cosets of $P_{i+2cd}(G)$ in $P_{i-2cd}(G)$ and additive cosets of $L_{i^{*}+2cd}$ in $L_{i^{*}-2cd}$ agree with one another and that multiplication modulo $P_{i+2cd}(G)$ is the same operation as addition modulo $L_{i^{*}+2cd}$. Here the details are more tricky to work out and we indicate how to proceed when $p\geq 3$; the case $p=2$ is very similar, but would need notational refinements as in [7, 20] which we want to forgo. If $x,y\in P_{i-2cd}(G)$, then on the Lie lattice side we obtain
$u_{2}(\underline{x},\underline{y})=\tfrac{1}{2}[\underline{x},\underline{y}]_{\mathrm{Lie}}\in[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}}\subseteq L_{i^{*}+2cd}$
and, for $n\geq 3$,
$u_{n}(\underline{x},\underline{y})\in p^{-\lfloor(n-1)/(p-1)\rfloor}\big{[}\!\!\cdot\!\cdot\!\cdot\!\!\big{[}[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}},\underbrace{L_{i^{*}-2cd}]_{\mathrm{Lie}},\ldots,L_{i^{*}-2cd}}_{\text{$n-2$ terms}}\big{]}_{\mathrm{Lie}}\\ \subseteq p^{-\lfloor(n-1)/(p-1)\rfloor}p^{n-2}L_{i^{*}+2cd}\subseteq L_{i^{*}+2cd}.$
This shows that all $x,y\in P_{i-2cd}(G)$ satisfy
$\underline{xy}=\Phi(\underline{x},\underline{y})\equiv\underline{x}+\underline{y}\quad\text{modulo $L_{i^{*}+2cd}$.}$ (2.9)
In particular, the elements $x$ and $y$ are multiplicative inverses of one another modulo $P_{i+2cd}(G)$ if and only if $\underline{x}$ and $\underline{y}$ are additive inverses of one another modulo $L_{i^{*}+2cd}$. By considering inverses of inverses, we conclude that multiplicative cosets of $P_{i+2cd}(G)$ are the same as additive cosets of $L_{i^{*}+2cd}$. Now (2.9) shows that multiplication modulo $P_{i+2cd}(G)$ is the same operation as addition modulo $L_{i^{*}+2cd}$: we have established the isomorphism of abelian groups
$\nicefrac{{\displaystyle P_{i-2cd}(G)}}{{\displaystyle P_{i+2cd}(G)}}\cong\nicefrac{{\displaystyle L_{i^{*}-2cd}}}{{\displaystyle L_{i^{*}+2cd}}}\qquad\text{via}\quad xP_{i+2cd}(G)\mapsto\underline{x}+L_{i^{*}+2cd}.$
Since $G$ acts unipotently on the finite $p$-group $P_{i}(G)/P_{i}(G)^{p^{2c}}$ of order $p^{2cd}$, we conclude that $P_{i+2cd}(G)\subseteq P_{i}(G)^{p^{2c}}$ and likewise $P_{i}(G)\subseteq P_{i-2cd}(G)^{p^{2c}}$. On the Lie lattice side these inclusions translate into
$L_{i^{*}+2cd}\subseteq p^{2c}L_{i^{*}}\subseteq\check{M}_{i^{*}}\subseteq L_{i^{*}}\subseteq\hat{M}_{i^{*}}\subseteq p^{-2c}L_{i^{*}}\subseteq L_{i^{*}-2cd}.$
Recalling that $\check{M}_{i^{*}}$ and $\hat{M}_{i^{*}}$ are $G$-invariant and satisfy $p^{2c}\hat{M}_{i^{*}}=\check{M}_{i^{*}}$, we conclude that they yield open normal subgroups $\check{U}_{i},\hat{U}_{i}\trianglelefteq_{\mathrm{o}}G$ such that
$\check{U}_{i}\subseteq P_{i}(G)\subseteq\hat{U}_{i}$
and such that the abelian factor group $\hat{U}_{i}/\check{U}_{i}\cong\hat{M}_{i^{*}}/\check{M}_{i^{*}}$ is homocyclic of exponent $p^{2c}$.
It remains to factorise $\check{U}_{i},\hat{U}_{i}$ into products of procyclic subgroups. We consider $\check{U}_{i}$; the proof for $\hat{U}_{i}$ works in the same way. Since $\check{U}_{i}=\check{M}_{i^{*}}$ as sets, we certainly have $x_{k}^{\,p^{\lfloor i^{*}\xi_{k}\rfloor+c}}\in\check{U}_{i}$ for $1\leq k\leq d$. This shows that
$\check{V}_{i}=\overline{\langle x_{1}^{p^{\lfloor i^{*}\xi_{1}\rfloor+c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rangle}\subseteq_{\mathrm{c}}U$
is a subset of $\check{U}_{i}$. Furthermore, [7, Prop. 3.7 and Sec. 4.2] show that $\check{V}_{i}$ is the disjoint union of $p^{\sum_{k=1}^{d}(\lfloor i^{*}\xi_{d}\rfloor-\lfloor i^{*}\xi_{k}\rfloor)}$ cosets of the open normal subgroup
$\overline{\langle x_{1}^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rangle}=U^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\trianglelefteq_{\mathrm{o}}U$
of index $\lvert U:U^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rvert=p^{(\lfloor i^{*}\xi_{d}\rfloor+c)d}$. This yields
$\mu_{U}(\check{V}_{i})=p^{\sum_{k=1}^{d}(\lfloor i^{*}\xi_{d}\rfloor-\lfloor i^{*}\xi_{k}\rfloor)}\big{\lvert}U:U^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\big{\rvert}^{-1}=p^{-cd-\sum_{k=1}^{d}\lfloor i^{*}\xi_{k}\rfloor},$
where $\mu_{U}$ denotes the normalised Haar measure on the group $U$. We recall that the normalised Haar measures $\mu_{U}$ and $\mu_{L}$ on the compact space $U=L$ are the same; see [20, Prop. A.2] and [6, Lem. 1.9.4]. This gives
$\mu_{U}(\check{U}_{i})=\mu_{L}(\check{M}_{i^{*}})=\lvert L:\check{M}_{i^{*}}\rvert^{-1}=p^{-cd-\sum_{k=1}^{d}\lfloor i^{*}\xi_{k}\rfloor}=\mu_{U}(\check{V}_{i}),$
and hence $\check{U}_{i}=\check{V}_{i}$ as wanted. ∎
It does not take much extra effort to prove Corollaries 1.3 and 1.4.
Proof of Corollary 1.3.
We continue to use the notation established in the proof of Theorem 1.2. Let $i\in\mathbb{N}$ with $i\geq\max\{j+2cd,6cd\}$. In the special case $c=0$ everything is easier and the following argument, which implicitly uses $c\geq 1$, can be skipped. Recall that $A_{i}=\hat{U}_{i}/\check{U}_{i}$ is a homocyclic group of rank $d$ and exponent $p^{2c}$; it has a unique homocyclic subgroup of rank $d$ and exponent $p^{c}$, namely $B_{i}=A_{i}^{\,p^{c}}$, whose pre-image in $\hat{U}_{i}$ is
$G_{i}=\overline{\langle x_{1}^{\,p^{\lfloor(i-j)\xi_{1}\rfloor}},\ldots,x_{d}^{\,p^{\lfloor(i-j)\xi_{d}\rfloor}}\rangle}\trianglelefteq_{\mathrm{o}}U.$
The group $P_{i}(G)$ contains $\check{U}_{i}$ and maps, modulo $\check{U}_{i}$, onto a subgroup $C_{i}$ of $A_{i}$. Thus $C_{i}\cong\prod_{k=1}^{d}C_{p^{m_{i,k}}}$ for suitable parameters $0\leq m_{i,1}\leq\ldots\leq m_{i,d}\leq 2c$, and
$\bigl{\lvert}\log_{p}\lvert A_{i}:C_{i}\rvert-\log_{p}\lvert A_{i}:B_{i}\rvert\bigr{\rvert}\leq\sum_{k=1}^{d}\lvert m_{i,k}-c\rvert\leq cd.$
This yields
$\bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\log_{p}\lvert G:G_{i}\rvert\bigr{\rvert}\leq cd,$
and clearly
$\log_{p}\lvert G:G_{i}\rvert=\log_{p}\lvert G:P_{j}(G)\rvert+\sum_{k=1}^{d}\lfloor(i-j)\xi_{k}\rfloor.$
Thus it suffices to observe that, with the notation $\sigma=\sum_{k=1}^{d}\xi_{k}$,
$\lfloor(i-j)\sigma\rfloor-d\leq\sum_{k=1}^{d}\lfloor(i-j)\xi_{k}\rfloor\leq\lfloor(i-j)\sigma\rfloor.\qed$
Proof of Corollary 1.4.
The uniqueness of the parameters $\xi_{1},\ldots,\xi_{d}$ follows by looking at the elementary divisors of increasing sections, similar to the proof of Lemma 2.1. Indeed, suppose that $j$, $U$, $\xi_{1},\ldots,\xi_{d}$, $c$ and $\check{U}_{i}$ satisfy the assertions of Theorem 1.2. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$. By [7, Prop. 3.9], there is $j^{\prime}\in\mathbb{N}$ with $j^{\prime}\geq j$ such that, for all $i\in\mathbb{N}$ with $i\geq j^{\prime}$, the subgroups $P_{i}(G),\check{U}_{i}\trianglelefteq_{\mathrm{o}}G$ are uniformly powerful and thus give rise to full $\mathbb{Z}_{p}$-sublattices $A_{i},\check{M}_{i}\leq_{\mathrm{o}}\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$, by the Lie correspondence.
For $i_{1},i_{2}\in\mathbb{N}$ with $j^{\prime}\leq i_{1}\leq i_{2}$ the finite abelian $p$-group $A_{i_{1}}/A_{i_{2}}$ decomposes as
$A_{i_{1}}/A_{i_{2}}\cong\mathbb{Z}_{p}/p^{m_{i_{1},i_{2},1}}\mathbb{Z}_{p}\,\oplus\ldots\oplus\,\mathbb{Z}_{p}/p^{m_{i_{1},i_{2},d}}\mathbb{Z}_{p}$
with parameters $0\leq m_{i_{1},i_{2},1}\leq\ldots\leq m_{i_{1},i_{2},d}$ reflecting the elementary divisors. As in the proof of Lemma 2.1, we conclude that
$\xi_{k}=\lim_{\begin{subarray}{c}j^{\prime}\leq i_{1}\leq i_{2}\\ \text{s.t.}\ i_{2}-i_{1}\to\infty\end{subarray}}\frac{m_{i_{1},i_{2},k}}{i_{2}-i_{1}}\qquad\text{for $1\leq k\leq d$.}\qed$
3. Hausdorff spectra with respect to the lower $p$-series
Our next task is to use the stratification results from Section 2 to pin down the Hausdorff dimensions of closed subgroups in $p$-adic analytic pro-$p$ groups with respect to the lower $p$-series.
Theorem 3.1.
Let $L$ be a non-zero $\mathbb{Z}_{p}$-lattice of dimension $d=\dim_{\mathbb{Z}_{p}}(L)$. Let $\mathcal{S}$ be a filtration series of $L$ which admits a stratification with growth rate $(\xi_{1},\ldots,\xi_{d})$. Then the Hausdorff spectrum of the additive group $L$, with respect to $\mathcal{S}$, is
$\operatorname{hspec}^{\mathcal{S}}(L)=\left\{\frac{\varepsilon_{1}\xi_{1}+\ldots+\varepsilon_{d}\xi_{d}}{\xi_{1}+\ldots+\xi_{d}}\mid\varepsilon_{1},\ldots,\varepsilon_{d}\in\{0,1\}\right\}.$
In particular, this shows that $\lvert\operatorname{hspec}^{\mathcal{S}}(L)\rvert\leq 2^{d}$.
Furthermore, every subgroup $H\leq_{\mathrm{c}}L$ has strong Hausdorff dimension with respect to $\mathcal{S}$.
Proof.
We write $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and choose a frame $(x_{1},\ldots,x_{d})$ for a stratification of $\mathcal{S}$. We also recall that $0<\xi_{1}\leq\xi_{2}\leq\ldots\leq\xi_{d}$. By [22, Lem. 2.2], any two equivalent filtration series define the same Hausdorff dimension on closed subgroups of $L$. Hence we may suppose that
$L_{i}=\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}.$
Let $H\leq_{\mathrm{c}}L$ be a $\mathbb{Z}_{p}$-sublattice of dimension $\dim_{\mathbb{Z}_{p}}(H)=r$, say. Since $\mathbb{Z}_{p}$ is a principal ideal domain, $H$ has a $\mathbb{Z}_{p}$-basis $(y_{1},\ldots,y_{r})$ such that the parameters
$j(s)=\max\big{(}\{0\}\cup\{k\mid 1\leq k\leq d\text{ and }a_{s,k}\neq 0\}\big{)},\quad 1\leq s\leq r,$
where we write $y_{s}=\sum_{k=1}^{d}a_{s,k}x_{k}$ with suitable coefficients $a_{s,k}\in\mathbb{Z}_{p}$, satisfy $j(1)<\ldots<j(r)$. We contend that $H$ has strong Hausdorff dimension
$\operatorname{hdim}_{L}^{\mathcal{S}}(H)=\frac{\xi_{j(1)}+\ldots+\xi_{j(r)}}{\xi_{1}+\ldots+\xi_{d}}.$
If $r=0$, the lattice $H$ equals $\{0\}$ and indeed $\operatorname{hdim}_{L}^{\mathcal{S}}(H)=0$. Now suppose that $r\geq 1$ and proceed by induction. We write $H=H_{1}\oplus H_{2}$, where $H_{1}=\bigoplus_{s=1}^{r-1}\mathbb{Z}_{p}y_{s}$ and $H_{2}=\mathbb{Z}_{p}y_{r}$. By induction, we have
$\lim_{i\to\infty}\frac{\log_{p}\lvert H_{1}+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}=\operatorname{hdim}_{L}^{\mathcal{S}}(H_{1})=\frac{\xi_{j(1)}+\ldots+\xi_{j(r-1)}}{\xi_{1}+\ldots+\xi_{d}}.$ (3.1)
Now consider $H_{2}=\mathbb{Z}_{p}y_{r}$ and recall that $y_{r}=\sum_{k=1}^{j(r)}a_{r,k}x_{k}$ with $a_{r,j(r)}\neq 0$. Let $v_{p}\colon\mathbb{Z}_{p}\to\mathbb{N}_{0}\cup\{\infty\}$ denote the $p$-adic valuation map, and put
$m=\min\{v_{p}(a_{r,k})\mid 1\leq k\leq j(r)\text{ with }\xi_{k}=\xi_{j(r)}\}\in\mathbb{N}_{0}.$
Then we see that, for all sufficiently large $i\in\mathbb{N}$,
$\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert=\lfloor i\xi_{j(r)}\rfloor-m$
and thus
$\lim_{i\to\infty}\frac{\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}=\lim_{i\to\infty}\frac{\lfloor i\xi_{j(r)}\rfloor-m}{\lfloor i\xi_{1}\rfloor+\ldots+\lfloor i\xi_{d}\rfloor}=\frac{\xi_{j(r)}}{\xi_{1}+\ldots+\xi_{d}}.$ (3.2)
For each $i\in\mathbb{N}_{0}$, we see that
$\log_{p}\lvert H+L_{i}:L_{i}\rvert+\log_{p}\lvert(H_{1}+L_{i})\cap(H_{2}+L_{i}):L_{i}\rvert\\ =\log_{p}\lvert H_{1}+L_{i}:L_{i}\rvert+\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert,$
and combining this observation with (3.1) and (3.2) we arrive at
[…]
It suffices to prove that the upper limit on the right-hand side equals 00; in particular, this also shows that it is a proper limit. Since $H_{2}=\mathbb{Z}_{p}y_{r}$, we find that for each $i\in\mathbb{N}$ there exists $n_{i}\in\mathbb{N}_{0}$ such that
$(H_{1}+L_{i})\cap(H_{2}+L_{i})=p^{n_{i}}\mathbb{Z}_{p}\,y_{r}+L_{i}.$
Moreover,
$p^{n_{i}}y_{r}\in H_{1}+L_{i}\subseteq\bigoplus_{k=1}^{j(r)-1}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=j(r)}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$
shows that $n_{i}+v_{p}(a_{r,j(r)})\geq\lfloor i\xi_{j(r)}\rfloor$ for all $i\in\mathbb{N}_{0}$. This implies that indeed
[…]
We are now in position to prove Theorem 1.5.
Proof of Theorem 1.5.
As in the proof of Theorem 1.2 we find $j\in\mathbb{N}$ such that $U=P_{j}(G)\leq_{\mathrm{o}}G$ is uniformly powerful and forms the starting point for the analysis carried out there. In particular, there are matching filtration series
$\mathcal{S}_{U}\colon U_{i^{*}}=P_{i}(G),\;i\in\mathbb{N}_{\geq j},\quad\text{and}\quad\mathcal{S}_{L}\colon L_{i^{*}}=\lambda_{i^{*}}(L),\;i\in\mathbb{N}_{\geq j},$
of the group $U$ and of the powerful $\mathbb{Z}_{p}$-Lie lattice $L$ associated to $U$, where we write $i^{*}=i-j$ as before.
Hausdorff dimensions of subgroups in $G$ with respect to the lower $p$-series $\mathcal{L}$ are closely linked to Hausdorff dimensions of subgroups in $U$ with respect to $\mathcal{S}_{U}$. Indeed, for every $H\leq_{\mathrm{c}}G$ and every open subgroup $W\leq_{\mathrm{o}}H$ with $W\subseteq U$,
$\operatorname{hdim}_{G}^{\mathcal{L}}(H)=\operatorname{hdim}_{U}^{\mathcal{S}_{U}}(W),$
and $H$ has strong Hausdorff dimension in $G$ if and only if $W$ does in $U$.
Moreover, we claim that, if $W$ is chosen so that it is powerful (using [7, Prop. 3.9]) with corresponding Lie sublattice $M\leq_{\mathrm{o}}L$, then
$\operatorname{hdim}_{U}^{\mathcal{S}_{U}}(W)=\operatorname{hdim}_{L}^{\mathcal{S}_{L}}(M),$ (3.3)
and again one of them is strong if and only if the other one is. The proof then concludes by the application of Theorems 1.1 and 3.1.
To see (3.3), let $i\in\mathbb{N}$ with $i\geq j$ and choose $m\in\mathbb{N}$ such that $U^{p^{m}}\leq U_{i^{*}}$, and accordingly $p^{m}L\leq L_{i^{*}}$. By [7, Cor. 4.15], multiplicative cosets of $U^{p^{m}}$ in $U$ correspond to additive cosets of $p^{m}L$ in $L$. Consequently, $U_{i^{*}}$ is the union of as many $U^{p^{m}}$-cosets as $L_{i^{*}}$ is the union of $p^{m}L$-cosets, hence
$\lvert U_{i^{*}}:U^{p^{m}}\rvert=\lvert L_{i^{*}}:p^{m}L\rvert.$
Since $W$ is powerful and $U_{i^{*}}$ is powerfully embedded in $U$, the group $WU_{i^{*}}$ is powerful and corresponds to the Lie lattice $M+L_{i^{*}}$. Again we obtain
$\lvert WU_{i^{*}}:U^{p^{m}}\rvert=\lvert M+L_{i^{*}}:p^{m}L\rvert.$
In combination this gives
$\frac{\log_{p}\lvert WU_{i^{*}}:U_{i^{*}}\rvert}{\log_{p}\lvert U:U_{i^{*}}\rvert}=\frac{\log_{p}\lvert WU_{i^{*}}:U^{p^{m}}\rvert-\log_{p}\lvert U_{i^{*}}:U^{p^{m}}\rvert}{\log_{p}\lvert U:U^{p^{m}}\rvert-\log_{p}\lvert U_{i^{*}}:U^{p^{m}}\rvert}\\ =\frac{\log_{p}\lvert M+L_{i^{*}}:p^{m}L\rvert-\log_{p}\lvert L_{i^{*}}:p^{m}L\rvert}{\log_{p}\lvert L:p^{m}L\rvert-\log_{p}\lvert L_{i^{*}}:p^{m}L\rvert}=\frac{\log_{p}\lvert M+L_{i^{*}}:L_{i^{*}}\rvert}{\log_{p}\lvert L:L_{i^{*}}\rvert},$
which shows that the Hausdorff dimension of $W$ in $U$ and the one of $M$ in $L$ are given by the lower limit of the same sequence. ∎
As indicated in the introduction, it is natural to look for sharp bounds for the size of the Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)$ of a $p$-adic analytic group $G$, with respect to its lower $p$-series, in terms of the dimension $\dim(G)$.
Example 3.2.
For $m\in\mathbb{N}$, let $q_{1},q_{2},\ldots,q_{m}$ denote the first $m$ consecutive primes, starting at $q_{1}=2$. We consider the metabelian pro-$p$ groups
$G_{m}=C\ltimes(A_{1}\times\ldots\times A_{m}),$
where $C=\langle y\rangle\cong\mathbb{Z}_{p}$ is infinite procyclic and, for each $j\in\{1,\ldots,m\}$, the group $A_{j}=\langle x_{j,1},\ldots,x_{j,q_{j}}\rangle\cong\mathbb{Z}_{p}^{\,q_{j}}$ is torsion-free abelian of rank $q_{j}$ and normal in $G_{m}$, the action of $C$ on $A_{j}$ being determined by
$[x_{j,k},y]=\begin{cases}x_{j,k+1}&\text{if $1\leq k\leq q_{j}-1$,}\\ x_{j,1}^{\,p}&\text{if $k=q_{j}$.}\end{cases}$
Clearly, the groups $G_{m}$ are $p$-adic analytic. Below we determine asymptotically the size of their Hausdorff spectra, with respect to the lower $p$-series. In particular, the estimate yields
$\frac{\log_{p}\bigl{\lvert}\operatorname{hspec}^{\mathcal{L}}(G_{m})\bigr{\rvert}}{\sqrt{\dim(G_{m})}}\to\infty\qquad\text{as $m\to\infty$.}$ (3.4)
We fix $m\in\mathbb{N}$ and set $G=G_{m}$. It is straightforward to check that $G$ has dimension
$\dim(G)=1+\sum_{j=1}^{m}q_{j}.$
Moreover, its lower $p$-series admits a stratification with growth rate
$(1,\,\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}},\,\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\,\ldots,\,\underbrace{\nicefrac{{1}}{{q_{m}}},\ldots,\nicefrac{{1}}{{q_{m}}}}_{\text{$q_{m}$ entries}}).$
Theorem 1.5 readily provides an upper bound for the size of the Hausdorff spectrum, namely
$\bigl{\lvert}\operatorname{hspec}^{\mathcal{L}}(G)\bigr{\rvert}\leq 2\prod_{j=1}^{m}(q_{j}+1).$
We arrive at a lower bound as follows. For $H\leq_{\mathrm{c}}G$ of the form $H=\prod_{j=1}^{m}B_{j}$ with $B_{j}\leq A_{j}$ and $\dim(B_{j})=d_{j}$, say, a direct analysis, similar to the one in the proof of Theorem 3.1, yields
$\operatorname{hdim}^{\mathcal{L}}_{G}(H)=\frac{\nicefrac{{d_{1}}}{{q_{1}}}+\nicefrac{{d_{2}}}{{q_{2}}}+\ldots+\nicefrac{{d_{m}}}{{q_{m}}}}{m+1}.$
We observe that, upon restricting to $0\leq d_{j}<q_{j}$ for $j\in\{1,\ldots,m\}$, the formula yields
$\prod_{j=1}^{m}q_{j}$
distinct values in the Hausdorff spectrum of $G$ with respect to $\mathcal{L}$.
The Prime Number Theorem provides the asymptotic equivalences
$\log\Big{(}2\prod_{j=1}^{m}(q_{j}+1)\Big{)}\sim\log\Big{(}\prod_{j=1}^{m}q_{j}\Big{)}\sim q_{m}\quad\text{and}\quad\sum_{j=1}^{m}q_{j}\sim\frac{q_{m}^{\,2}}{2\log q_{m}},$
as $m$ tends to infinity; compare with [25, Cor. 1] (and [2] for a more careful approximation). These estimates describe the size of the Hausdorff spectrum asymptotically and imply (3.4). ∎
Next we turn to the special situation, where the $p$-adic analytic group is just infinite. This class of groups is of considerable interest and has been studied in some detail; for instance, see [18]. We start with two somewhat more general auxiliary propositions.
Proposition 3.3.
Let $G$ be an infinite $p$-adic analytic pro-$p$ group, and let $\mathcal{S}$ be a filtration series of $G$. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to a uniformly powerful open normal subgroup $U\trianglelefteq_{\mathrm{o}}G$, equipped with the induced $G$-action, and suppose that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}G$-module.
Then every $H\leq_{\mathrm{c}}G$ has strong Hausdorff dimension
$\operatorname{hdim}_{G}^{\mathcal{S}}(H)=\frac{\dim(H)}{\dim(G)}.$
Proof.
Let $H\leq_{\mathrm{c}}G$. Without loss of generality, we may suppose that every open normal subgroup $N\trianglelefteq_{\mathrm{o}}G$ with $N\subseteq U$ is powerfully embedded in $U$; see [7, Prop. 3.9]. Then $\mathcal{S}|_{U}\colon U_{i}=G_{i}\cap U,\,i\in\mathbb{N}_{0}$ is a filtration series of $U$, consisting of uniformly powerful subgroups, and
$\operatorname{hdim}^{\mathcal{S}}_{G}(H)=\operatorname{hdim}^{\mathcal{S}}_{G}(H\cap U)=\operatorname{hdim}^{\mathcal{S}|_{U}}_{U}(H\cap U);$ (3.5)
compare with [21, Lem. 2.1].
The series $\mathcal{S}|_{U}$ translates over to the Lie lattice side to a filtration series $\mathcal{S}|_{L}\colon L_{i},\,i\in\mathbb{N}_{0}$, of $L$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity. Consequently, $\mathcal{S}|_{L}$ is equivalent to a filtration series of the form $\mathcal{S}|_{L}^{*}\colon L\geq p^{n_{1}}L\geq p^{n_{2}}L\geq\ldots$, for a suitable non-decreasing sequence of non-negative integers $n_{1},n_{2},\ldots$ tending to infinity.
Pick a uniformly powerful open subgroup $W\leq_{\mathrm{o}}H\cap U$, and let $M\leq L$ denote the corresponding Lie sublattice. Arguing as in the proof of Theorem 1.5, we deduce that
$\operatorname{hdim}^{\mathcal{S}|_{U}}_{U}(H\cap U)=\operatorname{hdim}^{\mathcal{S}|_{U}}_{U}(W)=\operatorname{hdim}^{\mathcal{S}|_{L}}_{L}(M)=\operatorname{hdim}^{\mathcal{S}|_{L}^{*}}_{L}(M).$ (3.6)
A direct inspection shows that the additive subgroup $M$ of $L$ has strong Hausdorff dimension
$\operatorname{hdim}^{\mathcal{S}|_{L}^{*}}_{L}(M)=\frac{\dim(M)}{\dim(L)}=\frac{\dim(H)}{\dim(G)}.$
Tracing our way back through (3.5) and (3.6) concludes the proof. ∎
Lemma 3.4.
Let $U$ be a uniformly powerful pro-$p$ group, and let $L$ be the associated $\mathbb{Z}_{p}$-Lie lattice. Then the following are equivalent:
  1. (a)
    $U$ is insoluble and just infinite.
  2. (b)
    $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}$-Lie algebra.
  3. (c)
    $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}U$-module.
Proof.
It is known that (a) implies (b); for instance, see [18, Prop. III.6].
Next we show that (b) implies (c). Suppose that $\mathfrak{L}=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}$-Lie algebra. Let $\mathfrak{M}$ denote a non-zero $\mathbb{Q}_{p}U$-submodule of $\mathfrak{L}$. In order to show that $\mathfrak{M}=\mathfrak{L}$ it suffices to prove that $M=\mathfrak{M}\cap L$ is a Lie ideal in $L$.
Let $\underline{x}\in M$ and $\underline{y}\in L$. Since $M\subseteq_{\mathrm{c}}L$, it suffices to prove that
$[\underline{x},\underline{y}]_{\mathrm{Lie}}\in M+p^{k-1}L\qquad\text{for all $k\in\mathbb{N}$ with $k\geq 2$.}$ (3.7)
Let $k\in\mathbb{N}$ with $k\geq 2$. From [24, IV (3.2.7)] we see that
$\underline{x}.y^{p^{k}}-\underline{x}=\sum_{n=1}^{\infty}\tfrac{1}{n!}\,[\underline{x},_{n}p^{k}\underline{y}]_{\mathrm{Lie}}=p^{k}\,[\underline{x},\underline{y}]_{\mathrm{Lie}}+\sum_{n=2}^{\infty}\tfrac{p^{nk}}{n!}\,[\underline{x},_{n}\underline{y}]_{\mathrm{Lie}}.$
We recall that $(n!)^{-1}\in p^{-\lfloor(n-1)/(p-1)\rfloor}\mathbb{Z}_{p}$ and observe that for $n\geq 2$,
$nk-\lfloor(n-1)/(p-1)\rfloor\geq k+(k-1)(n-1)\geq k+(k-1).$
This gives
$[\underline{x},\underline{y}]_{\mathrm{Lie}}\equiv_{p^{k-1}L}p^{-k}\big{(}\underline{x}.y^{p^{k}}-\underline{x}\big{)}\in\mathfrak{M},$
and (3.7) holds.
Finally, we argue that (c) implies (a). Suppose that $\mathfrak{L}=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}U$-module. Let $N\trianglelefteq_{\mathrm{c}}U$ be a non-trivial normal subgroup. Then $N$ contains a uniformly powerful characteristic open subgroup $K$, which supplies a corresponding non-trivial Lie sublattice $M\leq_{\mathrm{c}}L$. Since $K$ is normal in $U$, the lattice $M$ is $U$-invariant. This implies that the $\mathbb{Q}_{p}U$-submodule $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}M$ equals $\mathfrak{L}$ and gives
$\dim(N)=\dim(K)=\dim_{\mathbb{Z}_{p}}(M)=\dim_{\mathbb{Z}_{p}}(L)=\dim(U).$
Hence $N$ is open in $U$. ∎
Now we prove Theorem 1.6.
Proof of Theorem 1.6.
We treat the two situations, $G$ is just infinite and $G$ is nilpotent, one after the other.
(1) Suppose that $G$ is a just infinite $p$-adic analytic pro-$p$ group.
Case 1: $G$ is soluble. In this situation $G$ is virtually abelian and, in fact, an irreducible $p$-adic space group: there is an abelian open normal subgroup $A\trianglelefteq_{\mathrm{o}}G$ such that $A\cong\mathbb{Z}_{p}^{d}$ is a $\mathbb{Z}_{p}$-lattice and the finite $p$-group $\Gamma=G/A$ acts faithfully on $A$ and irreducibly on $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}A$. Thus the claims follow from Proposition 3.3.
Case 2: $G$ is insoluble. Let $U$ be a uniformly powerful open normal subgroup of $G$, and let $L$ be the Lie lattice associated to $U$. The $\mathbb{Q}_{p}$-Lie algebra $\mathfrak{L}=\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is known to be semisimple of homogeneous type: $\mathfrak{L}=\bigoplus_{k=1}^{q}\mathfrak{L}_{k}$, for some $p$-power $q$, with simple components $\mathfrak{L}_{1}\cong\ldots\cong\mathfrak{L}_{q}$; see [18, Prop. III.6]. Furthermore, the proof of [18, Prop. III.6] shows that $G$ permutes transitively the $q$ components $\mathfrak{L}_{k}$ of $\mathfrak{L}$. Using Lemma 3.4, we deduce that $\mathfrak{L}$ is a simple $\mathbb{Q}_{p}G$-module and the result follows from Proposition 3.3.
(2) Suppose that the $p$-adic analytic pro-$p$ group $G$ is nilpotent of class $c$. As in the proof of Theorem 1.2, let $j\in\mathbb{N}$ be such that $U=P_{j}(G)$ is uniformly powerful and $P_{i}(G)$ is powerfully embedded in $U$ for all $i\in\mathbb{N}$ with $i\geq j$. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$, equipped with the induced $G$-action, and let $L_{i^{*}}=\lambda_{i^{*}}(L)$, $i\in\mathbb{N}$ with $i\geq j$, denote the Lie sublattices corresponding to $P_{i}(G)$ which form with the lower $p$-series of $L$. As before, we write $i^{*}=i-j$ to streamline the notation.
Intersecting the terms $\gamma_{s}(G)$ of the lower central series of $G$ with $U$, we arrive at a descending series
$U=V_{1}\geq V_{2}\geq\ldots\geq V_{c+1}=1$
of closed normal subgroups of $G$ such that $[V_{s},G]\subseteq V_{s+1}$ for $1\leq s\leq c$. For each $s\in\{1,\ldots,c+1\}$, let
$W_{s}=\{x\in U\mid\exists m\in\mathbb{N}:x^{p^{m}}\in V_{s}\}\leq_{\mathrm{c}}U$
denote the isolator of $V_{s}$ in $U$, a powerful subgroup that is normal in $G$; compare with [7, Scholium to Thm. 9.10]. Raising elements to some power commutes with applying the $G$-action; hence it is straightforward to check that $[W_{s},G]\subseteq W_{s+1}$ for $1\leq s\leq c$. On the Lie lattice side, we arrive at a corresponding descending series
$L=M_{1}\geq M_{2}\geq\ldots\geq M_{c+1}=\{0\}$
of Lie sublattices such that $G$ acts trivially on each section $M_{s}/M_{s+1}$, $1\leq s\leq c$. Writing $\mathfrak{b}=\sum_{g\in G}(g-1)\mathbb{Z}_{p}G$ we deduce that, for $i\in\mathbb{N}$ with $i\geq j$,
$L_{i^{*}}=\lambda_{i^{*}}(L)=L.\mathfrak{a}_{G}^{\,i}=L.\Big{(}\mathfrak{b}+p\mathbb{Z}_{p}\Big{)}^{i}\subseteq\sum_{s=0}^{c}p^{i-s}L.\mathfrak{b}^{s}$
and hence
$p^{i^{*}}L\subseteq L_{i^{*}}\subseteq p^{i^{*}-c}L,$
where negative powers of $p$ are interpreted by embedding $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$. In other words, the filtration series $L_{i^{*}}$, $i\in\mathbb{N}$ with $i\geq j$, and $p^{i^{*}}L$, $i\in\mathbb{N}$ with $i\geq j$, are $c$-equivalent. Arguing as in the proof of Proposition 3.3, we obtain that every $H\leq_{\mathrm{c}}G$ has strong Hausdorff dimension
$\operatorname{hdim}_{G}^{\mathcal{L}}(H)=\frac{\dim(H)}{\dim(G)}.\qed$
4. Large finite Hausdorff spectra and free pro-$p$ groups
In this section we establish Theorem 1.7, yielding arbitrarily large finite Hausdorff spectra for one and the same pro-$p$ group, when we are free to vary the underlying filtration series. Subsequently, we study the Hausdorff spectra $\operatorname{hspec}^{\mathcal{L}}(G)$ of finite direct products $G$ of free pro-$p$ groups, with respect to the lower $p$-series $\mathcal{L}$, and prove Theorem 1.8.
By [22, Prop. 2.1], Theorem 1.7 is a consequence of the next result.
Proposition 4.1.
Let $L\cong\mathbb{Z}_{p}\oplus\mathbb{Z}_{p}$, and let $X\subseteq[0,1]$ be finite with $\{0,1\}\subseteq X$. Then there exists a filtration series $\mathcal{S}$ such that $\operatorname{hspec}^{\mathcal{S}}(L)=X$.
Proof.
We write $L=\mathbb{Z}_{p}x\oplus\mathbb{Z}_{p}y$ and $X=\{\xi_{1},\ldots,\xi_{n}\}$, where $n=\lvert X\rvert\geq 2$ and $10=\xi_{1}<\ldots<\xi_{n}=$ 1 are real numbers. For $k\in\{1,\ldots,n\}$ and $j\in\mathbb{N}_{0}$ we set
$t_{k,j}=\big{\lfloor}\tfrac{1}{k}\big{(}2^{2^{k+jn}}-2^{2^{k+jn-1}}\big{)}(1-\xi_{k})\big{\rfloor}k\in k\mathbb{N};$
thus $t_{k,j}$ is the maximal positive multiple of $k$ subject to the condition $t_{k,j}\leq\big{(}2^{2^{k+jn}}-2^{2^{k+jn-1}}\big{)}(1-\xi_{k})$. We put $L_{0}=L$, and for $k\in\{1,\ldots,n\}$ and $j\in\mathbb{N}_{0}$, we set
$L_{k+jn}=p^{2^{2^{k+jn-1}}}\mathbb{Z}_{p}\tilde{x}_{k,j}\oplus p^{2^{2^{k+jn}}}\mathbb{Z}_{p}y,\qquad\text{where}\quad\tilde{x}_{k,j}=x+\frac{1-p^{k+t_{k,j}}}{1-p^{k}}y.$
We observe that
$p^{2^{2^{k+jn}}}L\subseteq L_{k+jn}\subseteq p^{2^{2^{k+jn-1}}}L\quad\text{and}\quad\log_{p}\lvert L:L_{k+jn}\rvert=2^{2^{k+jn-1}}+2^{2^{k+jn}}.$
In particular, $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ forms a filtration series of $L$.
We decompose $\mathbb{N}$ into its residue classes modulo $n$,
$\mathbb{N}=I_{1}\cup\ldots\cup I_{n},\qquad\text{where}\;\;I_{k}=k+n\mathbb{N}_{0}\;\;\text{for}\;\;1\leq k\leq n,$
and set
$z_{k}=x+(1-p^{k})^{-1}\,y,\qquad\text{for $1\leq k\leq n$}.$
We observe that each of the $1$-dimensional subgroups $\mathbb{Z}_{p}z_{k}$, $1\leq k\leq n$, is isolated in $L$, i.e. equal to its isolator in $L$, and that any two of them intersect trivially: $\mathbb{Z}_{p}z_{k}\cap\mathbb{Z}_{p}z_{l}=\{0\}$ for $1\leq k<l\leq n$.
The Hausdorff spectrum of $L$ with respect to $\mathcal{S}$ trivially contains 00 and $1$. Any further points in the spectrum arise from $1$-dimensional subgroups. Let $H\leq L$ be a $1$-dimensional subgroup. We show below that, for each $k\in\{1,\ldots,n\}$,
$\lim_{i\in I_{k}}\frac{\log_{p}\lvert H+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}=\begin{cases}\xi_{k}&\text{if $H\leq\mathbb{Z}_{p}z_{k}$,}\\ 1&\text{otherwise.}\end{cases}$ (4.1)
This implies that
[…]
which concludes the proof.
It remains to establish (4.1). Let $k\in\{1,\ldots,n\}$, and first suppose that $H\leq\mathbb{Z}_{p}z_{k}$. Then $\operatorname{hdim}_{L}^{\mathcal{S}}(H)=\operatorname{hdim}_{L}^{\mathcal{S}}(\mathbb{Z}_{p}z_{k})$, and it suffices to show that $\operatorname{hdim}_{L}^{\mathcal{S}}(\mathbb{Z}_{p}z_{k})=\xi_{k}$. From the definition of $t_{k,j}$ we see that, for $j\in\mathbb{N}$,
$p^{2^{2^{j+kn-1}}}z_{k}\equiv_{L_{k+jn}}p^{2^{2^{j+kn-1}}}(z_{k}-\tilde{x}_{k,j})=p^{2^{2^{j+kn-1}}}p^{k+t_{k,j}}(1-p^{k})^{-1}\,y.$ (4.2)
This yields
$\begin{split}\lim_{i\in I_{k}}\frac{\log_{p}\lvert\mathbb{Z}_{p}z_{k}+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}&=\lim_{j\to\infty}\frac{\log_{p}\lvert\mathbb{Z}_{p}z_{k}+L_{k+jn}:L_{k+jn}\rvert}{\log_{p}\lvert L:L_{k+jn}\rvert}\\ &=\lim_{j\to\infty}\frac{2^{2^{k+jn-1}}+\big{(}2^{2^{k+jn}}-2^{2^{k+jn-1}}-k-t_{k,j}\big{)}}{2^{2^{k+jn-1}}+2^{2^{k+jn}}}\\ &=\lim_{j\to\infty}\frac{2^{2^{k+jn-1}}+\big{(}2^{2^{k+jn}}-2^{2^{k+jn-1}}\big{)}\big{(}1-(1-\xi_{k})\big{)}}{2^{2^{k+jn-1}}+2^{2^{k+jn}}}\\ &=\xi_{k}.\end{split}$
Now suppose that $H\not\leq\mathbb{Z}_{p}z_{k}$. We write $H=\mathbb{Z}_{p}w$, where $w=p^{m}x+by$ with $m\in\mathbb{N}_{0}$ and $b\neq p^{m}(1-p^{k})^{-1}$. Let $l\in\mathbb{N}_{0}$ such that $a=b-p^{m}(1-p^{k})^{-1}\in p^{l}\mathbb{Z}_{p}\smallsetminus p^{l+1}\mathbb{Z}_{p}$. We observe from (4.2) that for every $j\in\mathbb{N}_{0}$ with $k+t_{k,j}>l$, we have
$p^{2^{2^{k+jn-1}}}z_{k}\equiv_{L_{k+jn}}p^{2^{2^{k+jn-1}}}a_{j}^{\prime}\,y\quad\text{with}\quad a_{j}^{\prime}=p^{k+t_{k,j}}(1-p^{k})^{-1}\in p^{l+1}\mathbb{Z}_{p}$
and hence
$p^{2^{2^{k+jn-1}}-m}w=p^{2^{2^{k+jn-1}}}p^{-m}(p^{m}z_{k}+ay)\equiv_{L_{k+jn}}p^{2^{2^{k+jn-1}}}(a_{j}^{\prime}+p^{-m}a)y,$
where we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret negative powers of $p$. This yields
$\lim_{i\in I_{k}}\frac{\log_{p}\lvert H+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}=\lim_{j\to\infty}\frac{\log_{p}\lvert\mathbb{Z}_{p}w+L_{k+jn}:L_{k+jn}\rvert}{\log_{p}\lvert L:L_{k+jn}\rvert}\\ =\lim_{j\to\infty}\frac{\big{(}2^{2^{k+jn-1}}-m\big{)}+\big{(}2^{2^{k+jn}}-2^{2^{k+jn-1}}+m-l\big{)}}{2^{2^{k+jn-1}}+2^{2^{k+jn}}}=1.\qed$
Finally we turn to the Hausdorff spectra of non-abelian direct products of free pro-$p$ groups. We carefully adapt Lie-theoretic methods used by Garaialde Ocaña, Garrido and Klopsch [9, Sec. 4] to deal with other filtration series, in particular the Zassenhaus series. In parts, we follow closely a line of reasoning from a preprint version (arXiv:1901.03101v2) of [9] that adjusts well to our present setting. We slightly adjust our notation so that it matches better with the relevant parts of [9].
It is convenient to establish first the following central case of Theorem 1.8.
Theorem 4.2.
Let $F$ be a finitely generated non-abelian free pro-$p$ group. Then $F$ has full Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(F)=[0,1]$ with respect to the lower $p$-series $\mathcal{L}$.
The proof of the theorem requires some preparation. We consider the free pro-$p$ group $F$ on finitely many free generators $x_{1},\ldots,x_{d}$, for a given $d\geq 2$. As described in [24, Sec. II.1] and summarised in [23, Sec. 5], the lower $p$-series $P_{n}(F)$, $n\in\mathbb{N}$, gives rise to a mixed $\mathbb{F}_{p}$-Lie algebra $\boldsymbol{\Lambda}=(\Lambda,\mathsf{P})$. In concrete terms, this is the $\mathbb{N}$-graded $\mathbb{F}_{p}$-Lie algebra
$\Lambda=\bigoplus_{n=1}^{\infty}\Lambda_{n}$
whose homogeneous components are the finite elementary abelian $p$-groups $\Lambda_{n}=P_{n}(F)/P_{n+1}(F)$, for $n\in\mathbb{N}$, and whose Lie bracket on homogeneous elements is induced by the group commutator map $P_{m}(F)\times P_{n}(F)\to P_{m+n}(F)$, $(g,h)\mapsto[g,h]$; additionally, $\Lambda$ comes equipped with a family of $\mathbb{F}_{p}$-linear operators $\mathsf{P}_{n}\colon\Lambda_{n}\to\Lambda_{n+1}$, for $n\in\mathbb{N}$, induced by the $p$-power map $P_{n}(F)\to P_{n+1}(F)$, $g\mapsto g^{p}$ and thus satisfying a certain set of axioms that match the more general concept of a mixed Lie algebra.
For $p>2$, the structure of $\boldsymbol{\Lambda}$ is rather easy to describe. Indeed, the images $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}\in\Lambda_{1}$ of $x_{1},\ldots,x_{d}\in P_{1}(F)$ generate a free $\mathbb{F}_{p}$-Lie subalgebra $L=\bigoplus_{n=1}^{\infty}L_{n}$ of $\Lambda$, furnished with the grading $L_{n}=L\cap\Lambda_{n}$ which reflects the degree of Lie monomials. The mixed Lie algebra $\boldsymbol{\Lambda}$ can be understood as a free Lie ring on $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ over the polynomial ring $\mathbb{F}_{p}[\pi]$ in the sense that we may write
$\Lambda_{n}=\bigoplus_{m=1}^{n}\pi^{n-m}L_{m},\qquad\text{for $n\in\mathbb{N}$,}$ (4.3)
and interpret the operators $\mathsf{P}_{n}\colon\Lambda_{n}\to\Lambda_{n+1}$ as multiplication by the transcendental element $\pi$. For $p=2$, there is a small wrinkle for $\mathsf{P}_{1}\colon\Lambda_{1}\to\Lambda_{2}$, but the free $\mathbb{F}_{p}$-Lie algebra $L$ still appears as a subalgebra of $\Lambda$ and the mixed Lie subalgebra $\boldsymbol{\Lambda}^{\circ}$ based on $\Lambda^{\circ}=[\Lambda,\Lambda]=\bigoplus_{n=2}^{\infty}\Lambda_{n}$ admits exactly the same description as for $p>2$. It is convenient to postpone taking this extra hurdle and for the time being we assume for simplicity that $p>2$; at the end of the proof of Theorem 4.2 we explain how our reasoning easily adapts to the remaining case $p=2$.
For all practical purposes, $\boldsymbol{\Lambda}$ and its $\mathbb{F}_{p}$-Lie subalgebra $L$ can be realised as subalgebras of the free associative algebra $A=\mathbb{F}_{p}[\pi]\langle\!\langle\mathsf{x}_{1},\ldots,\mathsf{x}_{d}\rangle\!\rangle$ over the polynomial ring $\mathbb{F}_{p}[\pi]$, turned into an $\mathbb{F}_{p}$-Lie algebra via $[\mathsf{u},\mathsf{v}]=\mathsf{u}\mathsf{v}-\mathsf{v}\mathsf{u}$ and equipped with a ‘uniform’ $\mathsf{P}$-operator $\mathsf{u}\mapsto\pi\mathsf{u}$. Denoting by $A_{n}$ the $\mathbb{F}_{p}$-subspace spanned by elements of the form $\pi^{n-m}\mathsf{u}$, where $\mathsf{u}$ runs through all products of length $m\in\{0,1,\ldots,n\}$ in $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$, we see that $A$ decomposes as $A=\bigoplus_{n=0}^{\infty}A_{n}$ and becomes a graded algebra. From this perspective, the homogeneous components of $\Lambda$ arise simply as the intersections $\Lambda_{n}=\Lambda\cap A_{n}$ with homogeneous components of $A$.
It is well known that each homogeneous component $L_{n}=L\cap\Lambda_{n}$ of the free $\mathbb{F}_{p}$-Lie subalgebra $L$ on $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ is spanned, as an $\mathbb{F}_{p}$-vector space, by linearly independent basic Lie commutators in $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ of weight $n$; compare [12, Sec. 11] or [3, Sec. 2]. We recall and extend the pertinent notions.
Definition 4.3.
The weight of a Lie commutator in $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ is defined inductively: each generator $\mathsf{x}_{i}$ has weight $\operatorname{wt}(\mathsf{x}_{i})=1$, and, if $\mathsf{c}_{1},\mathsf{c}_{2}$ are Lie commutators in $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$, then $[\mathsf{c}_{1},\mathsf{c}_{2}]$ is a Lie commutator of weight $\operatorname{wt}([\mathsf{c}_{1},\mathsf{c}_{2}])=\operatorname{wt}(\mathsf{c}_{1})+\operatorname{wt}(\mathsf{c}_{2})$. The basic commutators are also defined inductively. The basic commutators of weight $1$ are the generators $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ in their natural order $\mathsf{x}_{1}<\ldots<\mathsf{x}_{d}$. For each $n\in\mathbb{N}$ with $n\geq 2$, after defining basic commutators of weight less than $n$, the basic commutators of weight $n$ are those Lie commutators $[\mathsf{u},\mathsf{v}]$ such that:
  1. (i)
    $\mathsf{u},\mathsf{v}$ are basic commutators with $\operatorname{wt}(\mathsf{u})+\operatorname{wt}(\mathsf{v})=n$,
  2. (ii)
    $\mathsf{u}>\mathsf{v}$; and if $\mathsf{u}=[\mathsf{y},\mathsf{z}]$ with basic commutators $\mathsf{y},\mathsf{z}$, then $\mathsf{v}\geq\mathsf{z}$.
Lastly, the total order is extended, subject to the condition that $\mathsf{u}<\mathsf{v}$ whenever $\operatorname{wt}(\mathsf{u})<\operatorname{wt}(\mathsf{v})$ and arbitrarily among the new basic commutators of weight $n$.
We extend the terminology to elements of $\Lambda$, the free $\mathbb{F}_{p}[\pi]$-Lie ring on $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$, as follows. For $n\in\mathbb{N}$, a generalised basic commutator $\tilde{\mathsf{c}}$ of weight $\operatorname{wt}(\tilde{\mathsf{c}})=n$ in $\Lambda$ is a homogeneous element of the form $\tilde{\mathsf{c}}=\pi^{n-m}\mathsf{c}\in\Lambda_{n}$, where $m\in\{1,2,\ldots,n\}$ and $\mathsf{c}$ is a basic commutator of weight $m$. We refer to $\mathsf{c}$ as the core part of the generalised basic commutator $\tilde{\mathsf{c}}$.
For $m\in\mathbb{N}$, we consider the principal Lie ideals
$I_{m}=\bigoplus_{n\geq m}\Lambda_{n}\trianglelefteq\Lambda\quad\text{and}\quad J_{m}=I_{m}\cap L=\bigoplus_{n\geq m}L_{n}\trianglelefteq L.$
For $\mathbb{F}_{p}$-Lie subalgebras $H$ of $\Lambda$ and $M$ of $L$ we define the relative densities
[…]
we say that a subalgebra has strong density if its density is given by a proper limit.
Witt’s formula describes the dimensions $\dim_{\mathbb{F}_{p}}(L_{n})$ of the homogenous components of the free Lie algebra $L$ and yields the following asymptotic estimates for the dimensions of principal quotients of $\Lambda$ and $L$; compare [17, Lem. 4.3].
Lemma 4.4.
In the set-up described above, including $d\geq 2$,
$\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})$ $=\frac{d^{n+2}}{n(d-1)^{2}}\big{(}1+o(1)\big{)}$ as $n\to\infty$.
and
$\dim_{\mathbb{F}_{p}}(L/J_{n+1})$ $=\frac{d^{n+1}}{n(d-1)}\big{(}1+o(1)\big{)}$ as $n\to\infty$.
From [4, Thm. 1] we obtain the following useful consequence.
Proposition 4.5.
Let $L$ be a non-abelian free $\mathbb{F}_{p}$-Lie algebra on finitely many generators. Then every finitely generated proper graded Lie subalgebra $M\lneqq L$ has strong density $\operatorname{den}_{L}(M)=0$.
In order to make use of this insight, we need to transfer the conclusion to the $\mathbb{F}_{p}[\pi]$-Lie subalgebra $H$ of $\Lambda$ generated by an $\mathbb{F}_{p}$-Lie subalgebra $M$ of $L$.
Lemma 4.6.
Let $L$ be the non-abelian free $\mathbb{F}_{p}$-Lie algebra on $d$ generators, regarded as an $\mathbb{F}_{p}$-subalgebra of the free $\mathbb{F}_{p}[\pi]$-Lie algebra $\Lambda$ on the same $d$ generators, as in the set-up described above. Let $M$ be a graded $\mathbb{F}_{p}$-Lie subalgebra of $L$ that has strong density $\operatorname{den}_{L}(M)=0$ in $L$. Then the $\mathbb{F}_{p}[\pi]$-Lie algebra $H$ of $\Lambda$ generated by $M$ has strong density $\operatorname{den}_{\Lambda}(H)=0$ in $\Lambda$.
Proof.
It is convenient to write, for $n\in\mathbb{N}$,
$\delta_{n}=\frac{\dim_{\mathbb{F}_{p}}((M+J_{n+1})/J_{n+1})}{\dim_{\mathbb{F}_{p}}(L/J_{n+1})}\quad\text{and}\quad\Delta_{n}=\frac{\dim_{\mathbb{F}_{p}}((H+I_{n+1})/I_{n+1})}{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})}.$
We are to deduce from $\lim_{n\to\infty}\delta_{n}=0$ that $\lim_{n\to\infty}\Delta_{n}=0$.
Let $\varepsilon\in(0,1)$, and choose $c\in\mathbb{N}$ such that $d^{-c}<\varepsilon/4$. By Lemma 4.4 there exists $n_{0}\in\mathbb{N}$ with $n_{0}>c$ such that for all $n\geq n_{0}$,
$\frac{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n-c+1})}{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})}\leq(1+\varepsilon)d^{-c}\leq(1+\varepsilon)\varepsilon/4<\varepsilon/2$ (4.4)
and, because $\delta_{n}\to 0$ as $n\to\infty$, we can also arrange that
$\bar{\delta}_{c}(n)=\max\{\delta_{m}\mid n-c+1\leq m\leq n\}\leq\varepsilon/2\qquad\text{for $n\geq n_{0}$.}$ (4.5)
We write $M=\bigoplus_{n=1}^{\infty}M_{n}$, where $M_{n}=M\cap L_{n}=M\cap\Lambda_{n}$ for $n\in\mathbb{N}$. Clearly, this gives $H=\bigoplus_{n=1}^{\infty}H_{n}$, with $H_{n}=\bigoplus_{m=1}^{n}\pi^{n-m}M_{m}$ for $n\in\mathbb{N}$. We deduce that for $n\in\mathbb{N}$,
$H+I_{n+1}$ $=\left(\bigoplus_{k=0}^{n-1}\pi^{k}\bigoplus_{m=1}^{n-k}M_{m}\right)\oplus I_{n+1}$
and likewise
$\Lambda$ $=\left(\bigoplus_{k=0}^{n-1}\pi^{k}\bigoplus_{m=1}^{n-k}L_{m}\right)\oplus I_{n+1}.$
This yields
$\dim_{\mathbb{F}_{p}}\big{(}(\pi^{c}\Lambda+I_{n+1})/I_{n+1}\big{)}=\sum_{k=c}^{n-1}\sum_{m=1}^{n-k}\dim_{\mathbb{F}_{p}}(L_{m})=\dim_{\mathbb{F}_{p}}(\Lambda/I_{n-c+1})$
and
$\frac{\dim_{\mathbb{F}_{p}}\big{(}(H+\pi^{c}\Lambda+I_{n+1})/(\pi^{c}\Lambda+I_{n+1})\big{)}}{\dim_{\mathbb{F}_{p}}\big{(}\Lambda/(\pi^{c}\Lambda+I_{n+1})\big{)}}=\frac{\sum_{k=0}^{c-1}\sum_{m=1}^{n-k}\dim_{\mathbb{F}_{p}}(M_{m})}{\sum_{k=0}^{c-1}\sum_{m=1}^{n-k}\dim_{\mathbb{F}_{p}}(L_{m})}\\ \leq\max_{0\leq k\leq c-1}\frac{\sum_{m=1}^{n-k}\dim_{\mathbb{F}_{p}}(M_{m})}{\sum_{m=1}^{n-k}\dim_{\mathbb{F}_{p}}(L_{m})}=\bar{\delta}_{c}(n).$
Using (4.4) and (4.5) we conclude that for $n\geq n_{0}$,
$\Delta_{n}\leq\frac{\dim_{\mathbb{F}_{p}}\big{(}(H+\pi^{c}\Lambda+I_{n+1})/(\pi^{c}\Lambda+I_{n+1})\big{)}+\dim_{\mathbb{F}_{p}}\big{(}(\pi^{c}\Lambda+I_{n+1})/I_{n+1}\big{)}}{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})}\\ \leq\left(1-\frac{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n-c+1})}{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})}\right)\bar{\delta}_{c}(n)+\frac{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n-c+1})}{\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})}<\varepsilon/2+\varepsilon/2=\varepsilon.$
Thus $\Delta_{n}\to 0$ as $n\to\infty$. ∎
The next result is a straightforward adaptation of [9, Thm 4.10].
Theorem 4.7.
Let $\Lambda$ be a non-abelian free $\mathbb{F}_{p}[\pi]$-Lie algebra on finitely many generators. Then there exists, for each $\alpha\in[0,1]$, an $\mathbb{F}_{p}[\pi]$-Lie subalgebra $H\leq\Lambda$ that is contained in $\Lambda^{\circ}=[\Lambda,\Lambda]$, that can be generated by generalised basic commutators and has density $\operatorname{den}_{\Lambda}(H)=\alpha$.
Proof.
As before, let $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$, for $d\geq 2$, be free generators of $\Lambda$, and recall that the $\mathbb{F}_{p}$-subalgebra generated by $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ is a free $\mathbb{F}_{p}$-Lie algebra $L=\bigoplus_{n=1}^{\infty}L_{n}$, furnished with the standard grading. We write $L^{\circ}=[L,L]=\bigoplus_{n=2}^{\infty}L_{n}$ and observe that
$\Lambda^{\circ}=\bigoplus_{n=2}^{\infty}\Lambda_{n}^{\circ},\qquad\text{with}\quad\Lambda_{n}^{\circ}=\bigoplus_{m=2}^{n}\pi^{n-m}L_{m},$
is the $\mathbb{F}_{p}[\pi]$-Lie subalgebra generated by $L^{\circ}$. Furthermore we have $\dim_{\mathbb{F}_{p}}(\Lambda_{n})=d+\dim_{\mathbb{F}_{p}}(\Lambda_{n}^{\circ})$ for each $n\in\mathbb{N}$. This implies that the sequence
$l^{\circ}(n)=\dim_{\mathbb{F}_{p}}\big{(}(\Lambda^{\circ}+I_{n+1})/I_{n+1}\big{)}=\sum\nolimits_{m=2}^{n}\dim_{\mathbb{F}_{p}}(\Lambda_{m}^{\circ}),\quad n\in\mathbb{N},$
is strictly increasing, grows exponentially in $n$ and satisfies
$\nicefrac{{\displaystyle l^{\circ}(n)\,}}{{\displaystyle\,\dim_{\mathbb{F}_{p}}(\Lambda/I_{n+1})}}\to 1\qquad\text{as $n\to\infty$;}$ (4.6)
see Lemma 4.4. We conclude that the subalgebra $\Lambda^{\circ}$ has strong density $1$ in $\Lambda$. Clearly, the trivial subalgebra $\{0\}$ has strong density 00 in $\Lambda$.
Suppose now that $\alpha\in(0,1)$. In order to produce an $\mathbb{F}_{p}[\pi]$-Lie subalgebra $H\leq\Lambda^{\circ}$ that can be generated by generalised basic commutators and has density $\operatorname{den}_{\Lambda}(H)=\alpha$ it suffices, by (4.6), to build an $\mathbb{F}_{p}[\pi]$-Lie subalgebra $H=\bigoplus_{m=1}^{\infty}H_{m}\leq\Lambda^{\circ}$, where $H_{m}=H\cap\Lambda_{m}^{\circ}=H\cap\Lambda_{m}$ for $m\in\mathbb{N}$, which is generated by generalised basic commutators and such that
  1. (i)
    $\alpha-\nicefrac{{1}}{{l^{\circ}(n)}}\leq\tfrac{1}{l^{\circ}(n)}\sum_{m=1}^{n}\dim_{\mathbb{F}_{p}}(H_{m})$ for all $n\in\mathbb{N}$ with $n\geq 2$; and
  2. (ii)
    $\frac{1}{l^{\circ}(n)}\sum_{m=1}^{n}\dim_{\mathbb{F}_{p}}(H_{m})\leq\alpha$ for infinitely many $n\in\mathbb{N}$.
We construct such a Lie subalgebra $H$ inductively as the union $H=\bigcup_{k=2}^{\infty}H(k)$ of an ascending chain of $\mathbb{F}_{p}[\pi]$-Lie subalgebras $H(2)\subseteq H(3)\subseteq\ldots$, where each term $H(k)$ is generated by a finite set $\widetilde{Y}_{k}$ of generalised basic commutators of weight at most $k$ and $\widetilde{Y}_{2}\subseteq\widetilde{Y}_{3}\subseteq\ldots$.
Let $k\in\mathbb{N}$ with $k\geq 2$. For $k=2$, observe that $l^{\circ}(2)=\dim_{\mathbb{F}_{p}}(\Lambda_{2}^{\circ})$ and pick $a\in\{0,1,\ldots,l^{\circ}(2)-1\}$ such that $\alpha-\nicefrac{{1}}{{l^{\circ}(2)}}\leq\nicefrac{{a}}{{l^{\circ}(2)}}\leq\alpha$. Let $H(2)$ denote the $\mathbb{F}_{p}[\pi]$-Lie subalgebra generated by a subset $\widetilde{Y}_{2}\subseteq\Lambda_{2}^{\circ}=L_{2}$, consisting of an arbitrary choice of $a$ basic commutators of weight $2$.
Now suppose that $k\geq 3$. Suppose further that we have already constructed an $\mathbb{F}_{p}[\pi]$-Lie subalgebra $H(k-1)\leq\Lambda^{\circ}$ which is generated by a finite set $\widetilde{Y}_{k-1}$ of generalised basic commutators of weight at most $k-1$, and that the inequalities in (i) hold for $H(k-1)$ in place of $H$ and $11\leq n\leq k-$ 1. For $n\geq k$, consider
$\beta_{n}=\frac{\sum_{i=1}^{n}\dim_{\mathbb{F}_{p}}(H(k-1)_{i})}{l^{\circ}(n)},$
where $H(k-1)_{i}=H(k-1)\cap\Lambda_{i}$ denotes the $i$th homogeneous component of the graded subalgebra $H(k-1)$.
First suppose that $\beta_{k}\leq\alpha$. From
$\frac{\sum_{i=1}^{k-1}\dim_{\mathbb{F}_{p}}(H(k-1)_{i})+\dim_{\mathbb{F}_{p}}(\Lambda_{k}^{\circ})}{l^{\circ}(k)}\geq\left(\alpha-\tfrac{1}{l^{\circ}(k-1)}\right)\tfrac{l^{\circ}(k-1)}{l^{\circ}(k)}+\tfrac{l^{\circ}(k)-l^{\circ}(k-1)}{l^{\circ}(k)}\\ \geq\alpha-\tfrac{1}{l^{\circ}(k)},$
we deduce that there is a finite set $\widetilde{Z}_{k}\subseteq\Lambda_{k}^{\circ}\smallsetminus H(k-1)_{k}$, consisting of generalised basic commutators of weight $k$, such that
$\alpha-\tfrac{1}{l^{\circ}(k)}\leq\frac{\sum_{i=1}^{k-1}\dim_{\mathbb{F}_{p}}\big{(}H(k-1)_{i}\big{)}+\dim_{\mathbb{F}_{p}}\big{(}H(k-1)_{k}\oplus\operatorname{\mathbb{F}_{p}-span}\widetilde{Z}_{k}\big{)}}{l^{\circ}(k)}\leq\alpha.$
Let $H(k)$ be the $\mathbb{F}_{p}[\pi]$-Lie subalgebra generated by $\widetilde{Y}_{k}=\widetilde{Y}_{k-1}\cup\widetilde{Z}_{k}$. Observe that the inequality in (i), respectively (ii), holds for $H(k)$ in place of $H$ and $1\leq n\leq k$, respectively $n=k$.
Now suppose that $\beta_{k}>\alpha$. Let $Y_{k-1}\subseteq L^{\circ}$ be the finite set consisting of the core parts of the elements of $\widetilde{Y}_{k-1}$, and let $M$ denote the $\mathbb{F}_{p}$-Lie subalgebra of $L$ generated by $Y_{k-1}$. By Proposition 4.5, the subalgebra $M$, has strong density $\operatorname{den}_{L}(M)=0$ in $L$. Clearly, $H(k-1)$ is contained in the $\mathbb{F}_{p}[\pi]$-Lie algebra generated by $Y_{k-1}$. By Lemma 4.6, the latter has strong density 00 in $\Lambda$, hence so does $H(k-1)$ in $\Lambda^{\circ}$. In particular, we find a minimal $k_{0}\geq k+1$ such that $\beta_{k_{0}}\leq\alpha$. Putting $H(k_{0}-1)=H(k_{0}-2)=\ldots=H(k)=H(k-1)$ and $\widetilde{Y}_{k_{0}-1}=\widetilde{Y}_{k_{0}-2}=\ldots=\widetilde{Y}_{k}=\widetilde{Y}_{k-1}$, we return to the previous case for $k_{0}>k$ in place of $k$. ∎
Our proof of Theorem 4.2 uses, in addition to Theorem 4.7, a partial correspondence between selected Lie subalgebras of $\Lambda$ and subgroups of the free pro-$p$ group $F$, which gives rise to the mixed Lie algebra $\boldsymbol{\Lambda}$ in the first place. The correspondence relies on a direct application of Philip Hall’s collection process, which we presently recall, following almost verbatim the account provided in Section 4.3 of a preprint version (arXiv:1901.03101v2) of [9].
We describe the collection process in a free group $\Gamma$ on finitely many generators $x_{1},\ldots,x_{d}$; by taking homomorphic images it can subsequently be applied to arbitrary $d$-generated groups with a chosen set of $d$ generators. The weight of group commutators in $x_{1},\ldots,x_{d}$ and basic group commutators in $x_{1},\ldots,x_{d}$, including a total ordering, are defined in complete analogy to Definition 4.3. At the heart of the Hall collection process for groups is the basic identity
$uv=vu[u,v],\qquad\text{valid for all $u,v\in\Gamma$.}$
A finite product
$w=c_{1}\cdots c_{k}c_{k+1}\cdots c_{m}$ (4.7)
of basic commutators $c_{1},\ldots,c_{m}$ is ordered if $c_{1}\leq\ldots\leq c_{m}$. The collected part of $w$ is the longest prefix $c_{1}\cdots c_{k}$ such that $c_{1}\leq\ldots\leq c_{k}$ and $c_{k}\leq c_{l}$ for all $l\in\{k+1,\ldots,m\}$. The rest of the product is called the uncollected part.
The collection process produces ordered product expressions for $w$ modulo uncollected terms of higher weight, as follows. Suppose that the leftmost occurrence of the smallest basic commutator in the uncollected part of the expression (4.7) for $w$ is $c_{l}$, in position $l$. Rewriting $c_{l-1}c_{l}$, we obtain a new expression for $w$, with $m+1$ factors:
$w=c_{1}\cdots c_{k}c_{k+1}\cdots c_{l}c_{l-1}[c_{l-1},c_{l}]c_{l+1}\cdots c_{m},$
where the new factor $[c_{l-1},c_{l}]$ is a basic commutator of higher weight than $c_{l}$; in particular, $[c_{l-1},c_{l}]>c_{l}$. If $c_{l}$, now occupying position $l-1$ in the product, is still in the uncollected part, we repeat the above step until $c_{l}$ becomes the rightmost commutator in the collected part. For any given $r\in\mathbb{N}$, carrying out finitely many iterations of the above procedure leads to an ordered product expression for $w$, modulo $\gamma_{r+1}(\Gamma)$, the $(r+1)$th term of the lower central series of $\Gamma$.
In order to apply the collection process to arbitrary elements of $\Gamma$, i.e. to products in $x_{1},\ldots,x_{d}$ and their inverses, one would need to accommodate also for inverses $c_{k}^{\,-1}$ of basic commutators as factors in our expressions. That is, one would have to consider collecting $u$ or $u^{-1}$ in expressions $v^{-1}u$, $vu^{-1}$ or $v^{-1}u^{-1}$. This can be done, as explained in [12, §11], but the group words we have to deal with are of the simple form (4.7).
The collection process can be used to show that, for every element $w\in\Gamma$ and every $r\in\mathbb{N}$, there is a unique approximation
$w\equiv_{\gamma_{r+1}(\Gamma)}b_{1}^{\,j_{1}}\cdots b_{t}^{\,j_{t}}$
of $w$ as a product of powers of the basic commutators $b_{1},\ldots,b_{t}$ of weight at most $r$, appearing in order; see [12, Thm. 11.2.4]. The existence of the decomposition follows from the collection process; the uniqueness requires further considerations.
Proof of Theorem 4.2.
We may regard $F=\widehat{\Gamma}_{p}$ as the pro-$p$ completion of the free group $\Gamma$ on $x_{1},\ldots,x_{d}$, with $d\geq 2$. The trivial subgroup $\{1\}$ and $F$ have Hausdorff dimensions 00 and $1$. Now suppose that $\alpha\in(0,1)$. Using Theorem 4.7, we find an $\mathbb{F}_{p}[\pi]$-Lie subalgebra $H\leq\Lambda$ that is contained in $\Lambda^{\circ}=[\Lambda,\Lambda]$, generated by an infinite set $\widetilde{Y}$ of generalised basic Lie commutators and satisfies $\operatorname{den}_{L}(H)=\alpha$. There is no harm in assuming that the set $\widetilde{Y}$ is minimal subject to generating $H$. In particular, this implies that no two elements of $\widetilde{Y}$ have the same core part. From $\widetilde{Y}\subseteq H\subseteq\Lambda^{\circ}$ we deduce that the core parts of all elements of $\widetilde{Y}$ lie in $L^{\circ}$.
There is a naive one-to-one correspondence between basic Lie commutators in $\mathsf{x}_{1},\ldots,\mathsf{x}_{d}$ and basic group commutators in $x_{1},\ldots,x_{d}$, which works by replacing occurrences of $\mathsf{x}_{i}$ by $x_{i}$ and exchanging Lie brackets with group commutators. If $\tilde{\mathsf{c}}=\pi^{k}\mathsf{c}\in\Lambda$ is a generalised basic Lie commutator, with $k\in\mathbb{N}_{0}$ and core part $\mathsf{c}$, and if $c\in\Gamma$ is the basic group commutator derived from $\mathsf{c}$, we call $c^{p^{k}}$ the generalised basic group commutator derived from $\tilde{\mathsf{c}}$, with core part $c$. Let $\widetilde{Y}_{\mathrm{grp}}$ denote the set of generalised basic group commutators derived from the elements of $\widetilde{Y}$. Consider the subgroup $\Delta=\langle\widetilde{Y}_{\mathrm{grp}}\rangle\leq\Gamma$; from $\widetilde{Y}\subseteq\Lambda^{\circ}$ we see that $\Delta\subseteq[\Gamma,\Gamma]$. Let $\overline{\Delta}$ denote the topological closure of $\Delta$ in $F$.
The canonical embedding $\Gamma\hookrightarrow\widehat{\Gamma}_{p}=F$ induces canonical isomorphisms $P_{n}(\Gamma)/P_{n+1}(\Gamma)\cong P_{n}(F)/P_{n+1}(F)=\Lambda_{n}$ for $n\in\mathbb{N}$. We change perspective and regard $\boldsymbol{\Lambda}=(\Lambda,\mathsf{P})$ as the mixed Lie algebra associated to the discrete group $\Gamma$, with respect to its lower $p$-series. Let
$\varphi\colon\Gamma\to\Lambda=\bigoplus_{n=1}^{\infty}\Lambda_{n},\quad\text{with components $\Lambda_{n}=P_{n}(\Gamma)/P_{n+1}(\Gamma)$,}$
denote the canonical map that sends $1$ to $01\varphi=$ 0 and each non-identity element $w\in P_{n}(\Gamma)\smallsetminus P_{n+1}(\Gamma)$ to $w\varphi=wP_{n+1}(\Gamma)\in\Lambda_{n}$, for $n\in\mathbb{N}$. It is easily seen that $\Delta\varphi\subseteq[\Gamma,\Gamma]\varphi=\Lambda^{\circ}$ is an $\mathbb{F}_{p}[\pi]$-Lie subalgebra and that
$\operatorname{hdim}_{F}^{\mathcal{L}}(\overline{\Delta})=\operatorname{den}_{\Lambda}(\Delta\varphi).$
Hence it suffices to show that $\Delta\varphi=H$; this duly implies $\alpha\in\operatorname{hspec}^{\mathcal{L}}(F)$.
We observe that, for every basic group commutator $c\in[\Gamma,\Gamma]$ and for $k\in\mathbb{N}_{0}$, the Lie element $(c^{p^{k}})\varphi$ is equal to the generalised basic Lie commutator $\pi^{k}\mathsf{c}\in\Lambda_{n}$, where $\mathsf{c}=c^{\mathrm{Lie}}$ is the basic Lie commutator corresponding to $c$ and $n=k+\operatorname{wt}(\mathsf{c})$. In particular, the construction ensures that $\widetilde{Y}_{\mathrm{grp}}\varphi=\widetilde{Y}$. Thus $\Delta\varphi$ contains $\widetilde{Y}$, and $H\subseteq\Delta\varphi$. It remains to show that, conversely, $w\varphi\in H$ for every non-trivial element $w\in\Delta$.
Let $w\in\Delta\smallsetminus\{1\}$, and choose $r\in\mathbb{N}$ such that $w\not\in P_{r+1}(\Gamma)$. The element $w$ is a finite product of generalised basic group commutators from the generating set $\widetilde{Y}_{\mathrm{grp}}$ of $\Delta$ and their inverses. However, working modulo the finite-index subgroup $P_{r+1}(\Gamma)$, we can avoid using inverses and assume that $w=\tilde{c}_{1}\cdots\tilde{c}_{m}$ with $m\in\mathbb{N}$ and $\tilde{c}_{1},\ldots,\tilde{c}_{m}\in\widetilde{Y}_{\mathrm{grp}}$, similar to (4.7).
We now apply the collection process, treating the generalised basic commutators $\tilde{c}_{i}=c_{i}^{\,p^{k(i)}}$ as if they were equal to their core parts $c_{i}$ carrying the exponents $p^{k(i)}$ simply as decorations; we recall that, by the minimal choice of $\widetilde{Y}$, each exponent $p^{k(i)}$ is uniquely determined by the core part $c_{i}$. As $P_{r+1}(\Gamma)\supseteq\gamma_{r+1}(\Gamma)$, the collection process shows that, modulo $P_{r+1}(\Gamma)$, there is a finite product decomposition
$w\equiv_{P_{r+1}(\Gamma)}\tilde{b}_{1}^{\,p^{e(1)}j_{1}}\,\tilde{b}_{2}^{\,p^{e(2)}j_{2}}\,\cdots\,\tilde{b}_{t}^{\,p^{e(t)}j_{t}},$ (4.8)
where (i) $\tilde{b}_{1},\ldots,\tilde{b}_{t}$ are commutator expressions in $\tilde{c}_{1},\ldots,\tilde{c}_{m}$ which upon replacing $\tilde{c}_{1},\ldots,\tilde{c}_{m}$ by their core parts $c_{1},\ldots,c_{m}$ yield basic group commutators $b_{1},\ldots,b_{t}$ that appear in order, and (ii) $e(1),\ldots,e(t),j_{1},\ldots,j_{t}\in\mathbb{N}_{0}$ with $p\nmid j_{1}\cdots j_{t}$.
For $1\leq s\leq t$, we define the weight of the expression $\tilde{b}_{s}$ to be
$\operatorname{wt}(\tilde{b}_{s})=\operatorname{wt}(b_{s})+A_{s},\qquad\text{with}\quad A_{s}=\sum_{i=1}^{m}a_{s,i}\,k(i),$
where $\operatorname{wt}(b_{s})$ is the weight of the corresponding ‘core’ commutator and $a_{s,i}$ denotes the number of occurrences of $c_{i}$ in $b_{s}$. Let $\mathsf{b}_{s}=b_{s}^{\mathrm{Lie}}$ denote the basic Lie commutator corresponding to $b_{s}$. Induction on $\operatorname{wt}(b_{s})$, which measures the ‘length’ of the commutator expression $\tilde{b}_{s}$, shows that
$\tilde{b}_{s}\in P_{\operatorname{wt}(\tilde{b}_{s})}(\Gamma)\smallsetminus P_{\operatorname{wt}(\tilde{b}_{s})+1}(\Gamma)\qquad\text{and}\qquad\tilde{b}_{s}\varphi=\pi^{A_{s}}\mathsf{b}_{s}.$ (4.9)
Let
$n=\min\big{\{}e(s)+\operatorname{wt}(\tilde{b}_{s})\mid 1\leq s\leq t\big{\}}$
and set
$S=\big{\{}s\mid 1\leq s\leq t\text{ and }e(s)+\operatorname{wt}(\tilde{b}_{s})=n\big{\}}.$
Then $w\in P_{n}(\Gamma)\smallsetminus P_{n+1}(\Gamma)$ and $w\varphi\in\Lambda_{n}$ is obtained from (4.8) by ignoring those factors $\tilde{b}_{s}^{\,p^{e(s)}j_{s}}$ for which $e(s)+\operatorname{wt}(\tilde{b}_{s})>n$. Indeed, (4.8) and (4.9) show that $w\varphi$ can be written as a non-trivial linear combination
$w\varphi=\sum_{s\in S}j_{s}\,\pi^{e(s)}\tilde{\mathsf{b}}_{s},\qquad\text{with$\tilde{\mathsf{b}}_{s}=\pi^{A_{s}}\mathsf{b}_{s}$ for $s\in S$},$
of the $\mathbb{F}_{p}$-linearly independent generalised basic Lie commutators $\pi^{e(s)+A_{s}}\mathsf{b}_{s}=\pi^{e(s)}\tilde{\mathsf{b}}_{s}$, $s\in S$, of weight $n$.
Each $\tilde{\mathsf{b}}_{s}$ is a Lie commutator of the generalised basic Lie commutators $\tilde{\mathsf{c}}_{1}=\pi^{k(1)}\mathsf{c}_{1},\ldots,\tilde{\mathsf{c}}_{m}=\pi^{k(m)}\mathsf{c}_{m}\in\widetilde{Y}$, in the same way as $\tilde{b}_{s}$ is a group commutator of $\tilde{c}_{1},\ldots,\tilde{c}_{m}$. As observed before, we have $\tilde{\mathsf{c}}_{i}=\tilde{c}_{i}\varphi\in\widetilde{Y}$ for $1\leq i\leq m$. Thus, $w\varphi$ lies indeed in $H$.
Up to this point we assumed, for convenience, that $p>2$. This meant that we could regard the mixed Lie algebra $\boldsymbol{\Lambda}$ as a free $\mathbb{F}_{p}[\pi]$-Lie algebra which simplified our discussion. As indicated earlier, for $p=2$, the mixed Lie algebra $\boldsymbol{\Lambda}$ has a slightly more complicated structure; but the mixed subalgebra $\boldsymbol{\Lambda}^{\circ}$ based on $\Lambda^{\circ}=[\Lambda,\Lambda]$ retains its properties. Moreover, it remains to settle the case that $\dim_{\mathbb{F}_{p}}(\Lambda_{n})=d+\dim_{\mathbb{F}_{p}}(\Lambda_{n}^{\circ})$ for each $n\in\mathbb{N}$. By going through our proofs up to this point, it is easily seen that these ingredients are all we really need. Thus we arrive at the same conclusions, also for $p=2$, notwithstanding the fact that $\boldsymbol{\Lambda}$ has a somewhat different structure. ∎
It remains to generalise our result to finite direct products of free pro-$p$ groups. We use, without any substantial change, the line of reasoning laid out in the proof of Theorem 1.3 in a preprint version (arXiv:1901.03101v2) of [9].
Proof of Theorem 1.8.
We consider the finite direct product $G=F_{1}\times\ldots\times F_{r}$ of finitely generated free pro-$p$ groups $F_{j}$, where $d=\max\{d(F_{j})\mid 1\leq j\leq r\}\geq 2$. Observe that the lower $p$-series $\mathcal{L}\colon P_{n}(G)$, $n\in\mathbb{N}$, of $G$ decomposes as the product of the lower $p$-series $\mathcal{L}_{j}\colon P_{n}(F_{j})$, $n\in\mathbb{N}$, of the direct factors $F_{j}$, for $1\leq j\leq r$. Setting $t=\#\{j\mid 1\leq j\leq r,\,d(F_{j})=d\}$, we may assume that $d(F_{j})=d$ for $1\leq j\leq t$ and $d(F_{j})<d$ for $t+1\leq j\leq r$. Lemma 4.4 shows that, for $t+1\leq j\leq r$,
$\lim_{n\to\infty}\frac{\log_{p}\lvert F_{j}:P_{n}(F_{j})\rvert}{\log_{p}\lvert F_{1}:P_{n}(F_{1})\rvert}=0.$ (4.10)
Now let $\alpha\in[0,1]$ and choose $k\in\{1,\ldots,t\}$ such that $\nicefrac{{(k-1)}}{{t}}\leq\alpha\leq\nicefrac{{k}}{{t}}$. By Theorem 4.2, we find a subgroup $H_{1}\leq_{c}F_{1}$ with $\operatorname{hdim}_{F_{1}}^{\mathcal{L}_{1}}(H_{1})=t\alpha-(k-1)$. Then
$H=H_{1}\times F_{2}\times\ldots\times F_{k}\times 1\times\ldots\times 1\leq_{\mathrm{c}}G$
has Hausdorff dimension
[…]
$=\frac{\operatorname{hdim}_{F_{1}}^{\mathcal{L}_{1}}(H_{1})+(k-1)}{t}$
$=\alpha.\qed$