### v2

# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

Abstract.
Let $G$ be a $p$-adic analytic pro-$p$ group of dimension $d$, with lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. We produce an approximate series which descends regularly in strata and whose terms deviate from $P_{i}(G)$ in a uniformly bounded way. This brings to light a new set of rational invariants $\xi_{1},\ldots,\xi_{d}\in[\nicefrac{{1}}{{d}},1]$, canonically associated to $G$, such that
$\sup_{i\in\mathbb{N}}\;\Bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\Big{(}\sum_{k=1}^{d}\xi_{k}\Big{)}(i-1)\Bigr{\rvert}<\infty$
and such that the Hausdorff dimensions $\operatorname{hdim}_{G}^{\mathcal{L}}(H)$ of closed subgroups $H\leq_{\mathrm{c}}G$ with respect to $\mathcal{L}$ are confined to the set
$\Big{\{}\Big{(}\sum_{k=1}^{d}\varepsilon_{k}\xi_{k}\Big{)}/\Big{(}\sum_{k=1}^{d}\xi_{k}\Big{)}\mid\varepsilon_{1},\ldots,\varepsilon_{d}\in\{0,1\}\Big{\}}.$
In particular, the Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)$ is discrete and consists of at most $2^{d}$ rational numbers.
Key words and phrases:
Lower $p$-series, $p$-adic analytic groups, Hausdorff dimension, Hausdorff spectrum
2020 Mathematics Subject Classification:
Primary 20E18; Secondary 22E20, 28A78
Funding acknowledgements. The first author was supported by the Spanish Government, grant MTM2017-86802-P, partly with FEDER funds, and by the Basque Government, grant IT974-16; he received funding from the European Union’s Horizon 2021 research and innovation programme under the Marie Sklodowska-Curie grant agreement, project 101067088. The second author received support from the Knut and Alice Wallenberg Foundation. The third author acknowledges support from EPSRC, grant EP/T005068/1 and from the Folke Lannér’s Fund. The research was partially conducted in the framework of the DFG-funded research training group GRK 2240: Algebro-Geometric Methods in Algebra, Arithmetic and Topology.
1. Introduction
Throughout $p$ denotes a prime, and $\mathbb{Z}_{p}$ is the ring of $p$-adic integers. We are interested in the lower $p$-series of $p$-adic analytic pro-$p$ groups and questions about Hausdorff dimension with respect to the lower $p$-series.
Let $G$ be a finitely generated pro-$p$ group. The lower $p$-series of $G$ – sometimes called the lower $p$-central series – is defined recursively as follows:
$P_{1}(G)=G,$ $\quad\text{and}\quad P_{i}(G)=\overline{P_{i-1}(G)^{p}\,[P_{i-1}(G),G]}\quad\text{for $i\in\mathbb{N}$ with $i\geq 2$,}$
where we adopt the notation from [5, Sec. 1.2]. It is the fastest descending central series of $G$ with elementary abelian sections and it consists of open normal subgroups forming a base of neighbourhoods for the element $1$. The lower $p$-series plays a significant role in the study of pro-$p$ groups. For instance, $\bigoplus_{i=1}^{\infty}P_{i}(G)/P_{i+1}(G)$ naturally carries the structure of an $\mathbb{F}_{p}$-Lie algebra and can frequently be used to analyse the structure of $G$ or its finite quotients; compare with [13, Chap. VIII], [19, Chap. II] and [18, Sec. 5].
In order to understand the lower $p$-series of $p$-adic analytic pro-$p$ groups, it is useful to consider an analogous series for $\mathbb{Z}_{p}$-lattices furnished with a group action. Let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a (continuous) right $G$-action, in other words, let $L$ be a right $\mathbb{Z}_{p}G$-module that is finitely generated and torsion-free as a $\mathbb{Z}_{p}$-module. The two-sided ideal
$\mathfrak{a}_{G}=\sum_{g\in G}(g-1)\mathbb{Z}_{p}G+p\mathbb{Z}_{p}G\,\trianglelefteq\,\mathbb{Z}_{p}G$
is the kernel of the natural epimorphism from $\mathbb{Z}_{p}G$ onto the finite field $\mathbb{F}_{p}$ that sends each group element to $1$; it is convenient to refer to $\mathfrak{a}_{G}$, somewhat informally, as the $p$-augmentation ideal of the group ring $\mathbb{Z}_{p}G$. The powers $\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the $p$-augmentation ideal intersect in $\{0\}$ and induce a topology on $\mathbb{Z}_{p}G$ which gives rise to the completed group ring $\mathbb{Z}_{p}[\![G]\!]$; see [5, Sec. 7.1]. The descending series of open $\mathbb{Z}_{p}G$-submodules
$\lambda_{i}(L)=L.\mathfrak{a}_{G}^{\,i}\quad\text{for $i\in\mathbb{N}_{0}$,}$ (1.1)
is the fastest descending $G$-central series of $L$ with elementary abelian sections. Accordingly we call the series (1.1) the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$, taking note of the notational shift in the index in comparison to the lower $p$-series of a group. As $G$ is a pro-$p$ group, it acts unipotently on every principal congruence quotient $L/p^{j}L$, $j\in\mathbb{N}$, and this implies that $\bigcap_{i\in\mathbb{N}_{0}}\lambda_{i}(L)=\{0\}$.
In general, a filtration series of $L$, regarded as a $\mathbb{Z}_{p}$-lattice, is a descending chain $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ of open $\mathbb{Z}_{p}$-sublattices $L_{i}\leq_{\mathrm{o}}L$ such that $\bigcap_{i\in\mathbb{N}_{0}}L_{i}=\{0\}$. For $c\in\mathbb{N}_{0}$, we say that two filtration series $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and $\mathcal{S}^{*}\colon L=L_{0}^{*}\geq L_{1}^{*}\geq\ldots$ are $c$-equivalent if for all $i\in\mathbb{N}$,
$p^{c}L_{i}\subseteq L_{i}^{*}\quad\text{and}\quad p^{c}L_{i}^{*}\subseteq L_{i}.$
We say that $\mathcal{S}$ and $\mathcal{S}^{*}$ are equivalent if they are $c$-equivalent for some $c\in\mathbb{N}_{0}$. This is but one of several ways to establish the concept or a closely related one. For instance, if $\log_{p}\lvert L_{i}+L_{i}^{*}:L_{i}\cap L_{i}^{*}\rvert\leq c$ for all $i\in\mathbb{N}$, then $\mathcal{S}$ and $\mathcal{S}^{*}$ are $c$-equivalent; conversely, if $\mathcal{S}$ and $\mathcal{S}^{*}$ are $c$-equivalent, then $\log_{p}\lvert L_{i}+L_{i}^{*}:L_{i}\cap L_{i}^{*}\rvert\leq 2c$ for all $i\in\mathbb{N}$. Thus $\mathcal{S}$ and $\mathcal{S}^{*}$ are equivalent if and only if $\lvert L_{i}+L_{i}^{*}:L_{i}\cap L_{i}^{*}\rvert$ is uniformly bounded as $i$ ranges over all positive integers.
Our aim is to show that the lower $p$-series of $L$ is equivalent to a filtration series that is descending very regularly, in a sense that we make precise now. Let $d=\dim_{\mathbb{Z}_{p}}L$ denote the dimension of the $\mathbb{Z}_{p}$-lattice $L$, and let $x_{1},\ldots,x_{d}$ be a $\mathbb{Z}_{p}$-basis for $L$. We say that a filtration series $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ of $L$ has a (vertical) stratification given by the frame $(x_{1},\ldots,x_{d})$ if there are positive real parameters $\xi_{1},\ldots,\xi_{d}$ with $\xi_{1}\leq\ldots\leq\xi_{d}$ such that $\mathcal{S}$ is equivalent to the filtration series
$\mathcal{S}^{*}\colon L_{i}^{*}=\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k};$
in this situation, $(\xi_{1},\ldots,\xi_{d})$ is uniquely determined by the frame and we refer to this $d$-tuple as the growth rate of the stratification. Informally, we call $\xi_{1},\ldots,\xi_{d}$ the ‘growth rates’ of the stratification. We say that the stratification $\mathcal{S}^{*}$ is given by $G$-invariant sublattices if all terms $L_{i}^{*}$ are $G$-invariant.
Theorem 1.1.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a $G$-action. Then the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification, and any two stratifications have the same growth rate.
Moreover, if $d=\dim_{\mathbb{Z}_{p}}L$ and if $(x_{1},\ldots,x_{d})$ is a frame for a stratification of the lower $p$-series of $L$ with growth rate $(\xi_{1},\ldots,\xi_{d})$, then
  1. (i)
    $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$ are rational numbers,
  2. (ii)
    for every $e\in\{1,\ldots,d-1\}$ with $\xi_{e}<\xi_{e+1}$ the sublattice $\bigoplus_{k=1}^{e}\mathbb{Z}_{p}x_{k}$ is $G$-invariant, and
  3. (iii)
    the entire stratification is given by $G$-invariant sublattices.
The above result yields an analogous description of the lower $p$-series of a $p$-adic analytic pro-$p$ group. We recall that, for every $p$-adic analytic pro-$p$ group $G$, there exists $j\in\mathbb{N}$ such that for every $i\in\mathbb{N}$ with $i\geq j$ the group $P_{i}(G)$ is uniformly powerful; see [5, Prop. 3.9 and Thm. 4.2]. Moreover, every minimal generating system $x_{1},\ldots,x_{d}$ of a uniformly powerful pro-$p$ group $U$ yields a product decomposition of $U=\overline{\langle x_{1}\rangle}\cdots\overline{\langle x_{d}\rangle}$ into $d$ procyclic subgroups, which in turn gives a system of $p$-adic coordinates $U\to\mathbb{Z}_{p}^{\,d}$, $x_{1}^{\,a_{1}}\cdots x_{d}^{\,a_{d}}\mapsto(a_{1},\ldots,a_{d})$ of ‘the second kind’; see [5, Thm. 4.9]. The dimension of $G$, denoted by $\dim(G)$, coincides with the rank of any uniformly powerful open subgroup of $G$.
Theorem 1.2.
Let $G$ be a $p$-adic analytic pro-$p$ group of dimension $d$. Then there exists $j\in\mathbb{N}$ such that
  1. (i)
    the group $U=P_{j}(G)\trianglelefteq_{\mathrm{o}}G$ is uniformly powerful and
  2. (ii)
    there is a minimal generating system $x_{1},\ldots,x_{d}$ for $U$ and there are rational numbers $\xi_{1},\ldots,\xi_{d}$ with $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$ and $c\in\mathbb{N}_{0}$ such that the following holds: for all $i\in\mathbb{N}$ with $i\geq\max\{j+2cd,6cd\}$, the following products of procyclic subgroups
    $\check{U}_{i}$ $=\overline{\langle x_{1}^{\,p^{\lfloor(i-j)\xi_{1}\rfloor+c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{\,p^{\lfloor(i-j)\xi_{d}\rfloor+c}}\rangle}$
    $\hat{U}_{i}$ $=\overline{\langle x_{1}^{\,p^{\lfloor(i-j)\xi_{1}\rfloor-c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{\,p^{\lfloor(i-j)\xi_{d}\rfloor-c}}\rangle}$
    form open normal subgroups $\check{U}_{i},\hat{U}_{i}\trianglelefteq_{\mathrm{o}}G$ such that
    $\check{U}_{i}\subseteq P_{i}(G)\subseteq\hat{U}_{i}$
    and the factor group $\hat{U}_{i}/\check{U}_{i}$ is homocyclic abelian of exponent $p^{2c}$.
Corollary 1.3.
In the set-up of Theorem 1.2 and with the additional notation $\sigma=\sum_{k=1}^{d}\xi_{k}\in[1,d]$, the index of $P_{i}(G)$ in $G$, for $i\geq\max\{j+2cd,6cd\}$, satisfies
$\Bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\big{(}\log_{p}\lvert G:P_{j}(G)\rvert+\lfloor(i-j)\sigma\rfloor\big{)}\Bigr{\rvert}\leq d(c+1).$
Corollary 1.4.
Let $G$ be a $p$-adic analytic pro-$p$ group of dimension $d$. Then any two approximations of the lower $p$-series as described in Theorem 1.2 involve the same rational growth parameters $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$.
The last corollary underlines that we have uncovered a new set of invariants which capture certain structural properties of $p$-adic analytic pro-$p$ groups.
One of our motivations for pinning down the lower $p$-series of $p$-adic analytic pro-$p$ groups was to settle some basic, but annoyingly outstanding questions about the Hausdorff spectra of such groups. The study and application of Hausdorff dimension in profinite groups is an ongoing affair; for instance, see [17] for an overview and [16, 7, 9, 10, 11, 12] for some recent developments. Let $G$ be an infinite finitely generated pro-$p$ group and let $\mathcal{S}\colon G_{i},\,i\in\mathbb{N}_{0}$, be a filtration series of $G$, that is, a descending chain $G=G_{0}\geq G_{1}\geq\ldots$ of open normal subgroups $G_{i}\trianglelefteq_{\mathrm{o}}G$ such that $\bigcap_{i\in\mathbb{N}_{0}}G_{i}=1$. In accordance with [1, 3], the Hausdorff dimension of a closed subgroup $H\leq_{\mathrm{c}}G$, with respect to $\mathcal{S}$, admits an algebraic interpretation as a logarithmic density:
[…]
and the Hausdorff spectrum of $G$, with respect to $\mathcal{S}$,
$\operatorname{hspec}^{\mathcal{S}}(G)=\{\operatorname{hdim}_{G}^{\mathcal{S}}(H)\mid H\leq_{\mathrm{c}}G\}\subseteq[0,1]$
provides a simple ‘snapshot’ of the typically rather complicated array of all subgroups in $G$. We say that $H$ has strong Hausdorff dimension in $G$, with respect to $\mathcal{S}$, if
$\operatorname{hdim}^{\mathcal{S}}_{G}(H)=\lim_{i\to\infty}\frac{\log_{p}\lvert HG_{i}:G_{i}\rvert}{\log_{p}\lvert G:G_{i}\rvert}$
is given by a proper limit.
Even for well-behaved groups, such as $p$-adic analytic pro-$p$ groups $G$, not only the Hausdorff dimension of individual subgroups, but also qualitative features of the Hausdorff spectrum of $G$ are known to be sensitive to the choice of $\mathcal{S}$; see [17, Thm. 1.3]. In our setting, where $G$ is a finitely generated pro-$p$ group, the group-theoretically relevant choices of $\mathcal{S}$ are primarily: the $p$-power series $\mathcal{P}$, the iterated $p$-power series $\mathcal{P}^{*}$, the Frattini series $\mathcal{F}$, the dimension subgroup (or Zassenhaus) series $\mathcal{D}$ and, last but not least, the lower $p$-series $\mathcal{L}\colon P_{i}(G),\,i\in\mathbb{N}$. It was shown in [17, Prop. 1.5] that, if $G$ is a $p$-adic analytic pro-$p$ group, then the Hausdorff dimensions of $H\leq_{\mathrm{c}}G$ with respect to the series $\mathcal{P},\mathcal{F},\mathcal{D}$ coincide. Furthermore, [3, Thm. 1.1] yields, for $\mathcal{S}$ any one of these series $\mathcal{P},\mathcal{F},\mathcal{D}$, that the spectrum
$\operatorname{hspec}^{\mathcal{S}}(G)=\Big{\{}\nicefrac{{\displaystyle\dim(H)}}{{\displaystyle\dim(G)}}\mid H\leq_{\mathrm{c}}G\Big{\}}$
is discrete and that all closed subgroups have strong Hausdorff dimension in $G$ with respect to $\mathcal{S}$. These results extend easily to Hausdorff dimensions with respect to the series $\mathcal{P}^{*}$.
It was seen in [17] that, contrary to indications in [3], the Hausdorff spectra of $p$-adic analytic pro-$p$ groups $G$ with respect to the lower $p$-series $\mathcal{L}$ follow a different pattern. Concretely, [17, Example 4.1] provides a family of $p$-adic analytic pro-$p$ groups $G$ such that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert/\dim(G)$ is unbounded as $\dim(G)\to\infty$. It remained open whether $\operatorname{hspec}^{\mathcal{L}}(G)$ could even be infinite for some $p$-adic analytic groups $G$. Theorem 1.1 allows us to establish that $\operatorname{hspec}^{\mathcal{L}}(G)$ is in fact finite for every $p$-adic analytic pro-$p$ group $G$ and to put in place further restrictions.
Theorem 1.5.
Let $G$ be an infinite $p$-adic analytic pro-$p$ group and let $d=\dim(G)$. Then there exist $\xi_{1},\ldots,\xi_{d}\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$ such that
$\operatorname{hspec}^{\mathcal{L}}(G)\subseteq\left\{\frac{\varepsilon_{1}\xi_{1}+\ldots+\varepsilon_{d}\xi_{d}}{\xi_{1}+\ldots+\xi_{d}}\mid\varepsilon_{1},\ldots,\varepsilon_{d}\in\{0,1\}\right\}.$
In particular, this shows that $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert\leq 2^{d}$. The parameters $\xi_{1},\ldots,\xi_{d}$ are the same as those appearing in Theorem 1.2, and all closed subgroups have strong Hausdorff dimension in $G$ with respect to $\mathcal{L}$.
It is a natural problem to improve, if possible, upon the upper bound for $\lvert\operatorname{hspec}^{\mathcal{L}}(G)\rvert$ as a function of $d=\dim(G)$. In Example 3.2 we manufacture a family of $p$-adic analytic pro-$p$ groups $G_{m}$, $m\in\mathbb{N}$, such that
$\frac{\log_{p}\bigl{\lvert}\operatorname{hspec}^{\mathcal{L}}(G_{m})\bigr{\rvert}}{\sqrt{\dim(G_{m})}}\to\infty\qquad\text{as $m\to\infty$.}$
Theorem 1.5 triggers more refined questions about Hausdorff spectra with respect to the lower $p$-series. We address some of these in a separate paper. For instance, one outstanding problem is to understand what renders the Hausdorff spectrum of a finitely generated pro-$p$ group finite. Using Lie-theoretic tools we show that non-abelian finite direct products of free pro-$p$ groups have full Hausdorff spectrum $[0,1]$ with respect to the lower $p$-series, in analogy to [7].
Notation. Generally, any new notation is explained when needed, sometimes implicitly. For convenience we collect basic conventions here.
[…]
The group-theoretic notation is mostly standard and in line, for instance, with its use in [5]. Tacitly, subgroups of profinite groups are generally understood to be closed subgroups. In some places, we emphasise that we are taking the topological closure of a set $X$ by writing $\overline{X}$.
Organisation. In Section 2, we study stratifications of the lower $p$-series and establish Theorems 1.1 and 1.2, as well as their corollaries. In Section 3 we turn to Hausdorff spectra with respect to the lower $p$-series: using the results from the previous section, we prove Theorem 1.5.
2. Stratifications of the lower $p$-series
Let $G$ be a profinite group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. We are interested in situations, where the lower $p$-series $\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, of $L$, as defined in (1.1), satisfies $\bigcap_{i\in\mathbb{N}_{0}}\lambda_{i}(L)=\{0\}$ and hence gives a filtration series. This is the case if and only if $G$ acts unipotently on every principal congruence quotient $L/p^{j}L$, $j\in\mathbb{N}$, which in turn is true if and only if the action of $G$ on the $\mathbb{Z}_{p}$-lattice $L\cong\mathbb{Z}_{p}^{\,d}$ factors through a pro-$p$ subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$. Since every closed subgroup of $\mathsf{GL}_{d}(\mathbb{Z}_{p})$ is topologically finitely generated, we may restrict without any true loss of generality to the situation that $G$ itself is a finitely generated pro-$p$ group. A theorem of Serre implies that every abstract action of the group $G$ on the group $L$ is continuous; see [5, Sec. 1]. Thus we arrive at the standard set-up which we use henceforth: $L$ is a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action, for a finitely generated pro-$p$ group $G$.
We start with a basic observation about growth rates of stratifications.
Lemma 2.1.
Let $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ be a filtration series of a $\mathbb{Z}_{p}$-lattice $L$ which admits a stratification. Then any two stratifications of $\mathcal{S}$ have the same growth rate.
Proof.
Let $d=\dim_{\mathbb{Z}_{p}}L$ and, for $i\in\mathbb{N}$, let $p^{m_{i,1}}$, …, $p^{m_{i,d}}$ with $0\leq m_{i,1}\leq\ldots\leq m_{i,d}$ be the elementary divisors of the finite $\mathbb{Z}_{p}$-module $L/L_{i}$ so that
$L/L_{i}\cong\mathbb{Z}_{p}/p^{m_{i,1}}\mathbb{Z}_{p}\,\oplus\ldots\oplus\,\mathbb{Z}_{p}/p^{m_{i,d}}\mathbb{Z}_{p}.$
Let $(x_{1},\ldots,x_{d})$ be the frame for a stratification of $\mathcal{S}$, with growth rate $(\xi_{1},\ldots,\xi_{d})$, say. Then there is a constant $c\in\mathbb{N}$ such that $\mathcal{S}$ is $c$-equivalent to the filtration series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Recall that $\xi_{1}\leq\ldots\leq\xi_{d}$ and choose $j\in\mathbb{N}$ such that $\lfloor j\xi_{e}\rfloor+c<\lfloor j\xi_{e+1}\rfloor-c$ for all $e\in\{1,\ldots,d-1\}$ with $\xi_{e}<\xi_{e+1}$. By definition of stratification, this implies that for all $i\in\mathbb{N}$ with $i\geq j$,
$\lvert m_{i,k}-\lfloor i\xi_{k}\rfloor\rvert\leq c\qquad\text{for $1\leq k\leq d$.}$
Thus each $\xi_{k}=\lim_{i\to\infty}m_{i,k}/i$ is uniquely determined by $\mathcal{S}$. ∎
Next we recall and adapt the following definitions from [14]: the lower and upper level of an open $\mathbb{Z}_{p}G$-submodule $M$ in $L$ are
$\ell_{L}(M)=\min\{k\in\mathbb{N}_{0}\mid p^{k}L\subseteq M\}\;\text{and}\;u_{L}(M)=\max\{k\in\mathbb{N}_{0}\mid M\subseteq p^{k}L\},$
and the rigidity of $L$ is
$r(L)=\sup\{\ell_{L}(M)-u_{L}(M)\mid\text{$M$ is an open $\mathbb{Z}_{p}G$-submodule of $L$}\}.$
For the trivial lattice $L=\{0\}$ the formalism yields $r(L)=-\infty$.
The following auxiliary result is a variation of [14, Prop. 4.3] and can be proved in the same way, using compactness.
Proposition 2.2.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Then $r(L)$ is finite if and only if $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}G$-module.
In conjunction with Lemma 2.1, the next proposition already establishes a special case of Theorem 1.1.
Proposition 2.3.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Suppose that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ is a simple $\mathbb{Q}_{p}G$-module. Then the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification with constant growth rate $(\xi,\ldots,\xi)$, where $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$ for $d=\dim_{\mathbb{Z}_{p}}L$.
Proof.
Denote by $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, the terms of the lower $p$-series of $L$, and for $i\in\mathbb{N}_{0}$ put $u_{i}=u_{L}(L_{i})$. By Proposition 2.2 the $\mathbb{Z}_{p}G$-module $L$ has finite rigidity $r=r(L)$. Consequently, the lower $p$-series of $L$ is $r$-equivalent to the filtration series $\mathcal{S}\colon L\geq p^{u_{1}}L\geq p^{u_{2}}L\geq\ldots$, and it suffices to show that $\mathcal{S}$ is equivalent to a series of the form $p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$, for suitable $\xi\in[\nicefrac{{1}}{{d}},1]\cap\mathbb{Q}$.
[…]
$\varphi_{i}\colon p^{u_{i}}L\to p^{u_{i}}L/p^{u_{i}+r}L\cong R$
defined by the dictum $p^{u_{i}}x_{k}\mapsto z_{k}$ for $1\leq k\leq d$. For each $i\in\mathbb{N}$, we have $\ell_{L}(L_{i})\leq u_{i}+r$, hence $p^{u_{i}+r}L\leq L_{i}\leq p^{u_{i}}L$ by the definition of $u_{i}$ and $r$, and we consider $R_{i}=L_{i}\varphi_{i}$.
Since $R$ is finite, there exist $j,m\in\mathbb{N}$ such that $R_{j+m}=R_{j}$, and therefore $L_{j+m}=p^{n}L_{j}$ for $n=u_{j+m}-u_{j}\geq 1$. Furthermore $p^{m}L_{j}\subseteq L_{j+m}$ yields $n\leq m$. We see that $L_{i+m}=L_{j+m}.\mathfrak{a}_{\mathbb{Z}_{p}G}^{\,i-j}=p^{n}L_{j}.\mathfrak{a}_{\mathbb{Z}_{p}G}^{\,i-j}=p^{n}L_{i}$ for all $i\in\mathbb{N}$ with $i\geq j$. In particular, this shows that $u_{i+m}=u_{i}+n$ for $i\geq j$. Put $\xi=n/m\in(0,1]\cap\mathbb{Q}$ and
$c=\bigl{\lceil}1+\max\big{\{}\lvert u_{i}-i\xi\rvert\bigm{|}0\leq i\leq j+m-1\big{\}}\bigr{\rceil}\in\mathbb{N}.$
Then
$\lvert u_{i}-\lfloor i\xi\rfloor\rvert\leq\lvert u_{i}-i\xi\rvert+1\leq c\qquad\text{for all $i\in\mathbb{N}_{0}$,}$
and the filtration series $\mathcal{S}$ is $c$-equivalent to $\mathcal{S}^{*}\colon p^{\lfloor i\xi\rfloor}L$, $i\in\mathbb{N}_{0}$. Furthermore, we notice that $L=L_{0}\supsetneqq L_{1}\supsetneqq\ldots$ yields the trivial lower bound $\lvert L:L_{i}\rvert\geq p^{i}$. Combining these two results, we arrive at
$\xi=\lim_{i\to\infty}\frac{d\lfloor i\xi\rfloor}{di}=\lim_{i\to\infty}\frac{\log_{p}\lvert L:p^{\lfloor i\xi\rfloor}L\rvert}{di}=\lim_{i\to\infty}\frac{\log_{p}\lvert L:L_{i}\rvert}{di}\geq\nicefrac{{1}}{{d}}.\qed$
The next proposition provides a key tool for proving Theorem 1.1; with its corollary it yields, in particular, the last two assertions in the theorem.
Proposition 2.4.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Suppose that the lower $p$-series of $L$ admits a stratification, with the frame $(x_{1},\ldots,x_{d})$ and growth rate $(\xi_{1},\ldots,\xi_{d})$.
Suppose that $e\in\{1,\ldots,d-1\}$ is such that $\xi_{e}<\xi_{e+1}$. Then the $\mathbb{Z}_{p}$-lattice $M=\bigoplus_{k=1}^{e}\mathbb{Z}_{p}x_{k}$ is $G$-invariant. Moreover the lower $p$-series of $M$ has a stratification with frame $(x_{1},\ldots,x_{e})$ and growth rate $(\xi_{1},\ldots,\xi_{e})$.
Proof.
The $\mathbb{Z}_{p}$-lattice $L$ decomposes as the direct sum of its sublattices $M$ and $N=\bigoplus_{k=e+1}^{d}\mathbb{Z}_{p}x_{k}$. Choose $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}=\lambda_{i}(L)$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the series $\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$, $i\in\mathbb{N}_{0}$. Then $\xi_{e+1}=\min\{\xi_{e+1},\ldots,\xi_{d}\}$ implies
$L_{i}\cap N\subseteq\left(\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}x_{k}\right)\cap N\subseteq p^{\lfloor i\xi_{e+1}\rfloor-c_{1}}N\quad\text{for all $i\in\mathbb{N}_{0}$,}$ (2.1)
where we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret negative exponents occurring for small values of $i$.
For a contradiction, assume that $M$ is not $G$-invariant, i.e. not a $\mathbb{Z}_{p}G$-submodule. Then there are elements $\alpha_{1},\ldots,\alpha_{e}\in\mathbb{Z}_{p}G$ and $c_{2}\in\mathbb{N}$ such that $y=\sum_{k=1}^{e}x_{k}\alpha_{k}\in N\smallsetminus p^{c_{2}}N$. For every $i\in\mathbb{N}$, we deduce from $\xi_{e}=\max\{\xi_{1},\ldots,\xi_{e}\}$ that $p^{\lfloor i\xi_{e}\rfloor+c_{1}}M\subseteq L_{i}$ and hence
$p^{\lfloor i\xi_{e}\rfloor+c_{1}}y\in(L_{i}\cap N)\smallsetminus p^{\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}}N.$
Thus (2.1) yields $\lfloor i\xi_{e}\rfloor+c_{1}+c_{2}>\lfloor i\xi_{e+1}\rfloor-c_{1}$. Since $\xi_{e}<\xi_{e+1}$ this yields the desired contradiction for sufficiently large $i\in\mathbb{N}$.
It remains to prove that the lower $p$-series $M_{i}=\lambda_{i}(M)$, $i\in\mathbb{N}_{0}$, of $M$ has a stratification with the frame $(x_{1},\ldots,x_{e})$ and growth rate $(\xi_{1},\ldots,\xi_{e})$. As before we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret any negative powers of $p$, whose occurrence does not cause any serious problems. We put
$b=c_{1}+b_{1}\qquad\text{with}\qquad b_{1}=\left\lceil\frac{3c_{1}+3}{\xi_{e+1}-\xi_{e}}\right\rceil.$
From the choice of $c_{1}$ and from $b\geq c_{1}$ we deduce that, for all $i\in\mathbb{N}_{0}$,
$M_{i}\subseteq L_{i}\cap M\subseteq\left(\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}\,x_{k}\right)\cap M\subseteq p^{-b}\,\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}.$
Thus it suffices to show that, for all $i\in\mathbb{N}_{0}$ with $i\geq b_{1}$,
$p^{b}\,\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}\subseteq M_{i}.$
Let $i\in\mathbb{N}_{0}$ with $i\geq b_{1}$. From $p\in\mathfrak{a}_{G}$ it follows that $p^{b_{1}}M_{i-b_{1}}\subseteq M_{i-b_{1}}.\mathfrak{a}_{G}^{\,b_{1}}=M_{i}$, and it suffices to establish that $p^{c_{1}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}\subseteq M_{i-b_{1}}$. There is no harm in working modulo $p\big{(}p^{c_{1}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}\big{)}$ and hence it is enough to show that
$\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}}\mathbb{Z}_{p}\,x_{k}\subseteq M_{i-b_{1}}+\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}+1}\mathbb{Z}_{p}\,x_{k}.$
Recall that $L=M\oplus N$ as a $\mathbb{Z}_{p}$-lattice and let $\tau\colon L\to M$ denote the projection along $N$. Recall further that $\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}}\mathbb{Z}_{p}\,x_{k}\subseteq L_{i}$. Hence it is enough to show that
$L_{i}\tau\subseteq M_{i-b_{1}}+\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor+c_{1}+1}\mathbb{Z}_{p}\,x_{k}.$ (2.2)
From $L_{b_{1}}\subseteq M+p^{\lfloor b_{1}\xi_{e+1}\rfloor-c_{1}}N$ we conclude that
$L_{i}\tau=\big{(}L_{b_{1}}.\mathfrak{a}_{G}^{\,i-b_{1}}\big{)}\tau\subseteq\big{(}\underbrace{M.\mathfrak{a}_{G}^{\,i-b_{1}}}_{=M_{i-b_{1}}}+p^{\lfloor b_{1}\xi_{e+1}\rfloor-c_{1}}\underbrace{N.\mathfrak{a}_{G}^{\,i-b_{1}}}_{\subseteq L_{i-b_{1}}}\big{)}\tau,$
and hence $L_{i-b_{1}}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor(i-b_{1})\xi_{k}\rfloor-c_{1}}\mathbb{Z}_{p}x_{k}$ gives
$L_{i}\tau$ $\subseteq M_{i-b_{1}}+\big{(}p^{\lfloor b_{1}\xi_{e+1}\rfloor-c_{1}}L_{i-b_{1}}\big{)}\tau$
$\subseteq M_{i-b_{1}}+\bigoplus_{k=1}^{e}p^{\lfloor b_{1}\xi_{e+1}\rfloor+\lfloor(i-b_{1})\xi_{k}\rfloor-2c_{1}}\mathbb{Z}_{p}\,x_{k}.$
Now our choice of $b_{1}$ guarantees that, for each $k\in\{1,\ldots,e\}$,
$\lfloor b_{1}\xi_{e+1}\rfloor+\lfloor(i-b_{1})\xi_{k}\rfloor-2c_{1}\geq b_{1}\xi_{e+1}+(i-b_{1})\xi_{k}-2c_{1}-2\\ =i\xi_{k}+c_{1}+b_{1}(\xi_{e+1}-\xi_{k})-(3c_{1}+2)\geq\lfloor i\xi_{k}\rfloor+c_{1}+1,$
and this yields (2.2). ∎
Corollary 2.5.
Let $G$ be a finitely generated pro-$p$ group, and let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a right $G$-action. Then every stratification of the lower $p$-series of $L$ is given by $G$-invariant sublattices.
Proof.
Suppose that the lower $p$-series of $L$ admits a stratification, with the frame $(x_{1},\ldots,x_{d})$ and growth rate $(\xi_{1},\ldots,\xi_{d})$. There are $\vartheta_{1}<\ldots<\vartheta_{r}$, for some $r\in\{1,\ldots,d\}$, such that $\{\xi_{k}\mid 1\leq k\leq d\}=\{\vartheta_{s}\mid 1\leq s\leq r\}$. For $1\leq s\leq r$ let $d_{s}$ denote the multiplicity of $\vartheta_{s}$ in $(\xi_{1},\ldots,\xi_{d})$, and for $0\leq s\leq r$ put $D(s)=d_{1}+\ldots+d_{s}$. Then the $\mathbb{Z}_{p}$-lattice $L$ decomposes as
$L=N_{1}\oplus\ldots\oplus N_{r},\quad\text{where}\quad N_{s}=\bigoplus_{k=D(s-1)+1}^{D(s)}\mathbb{Z}_{p}x_{k}\quad\text{for $1\leq s\leq r$,}$
and our stratification of the lower $p$-series of $L$ is realised by the series
$\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}N_{s},\quad i\in\mathbb{N}_{0}.$ (2.3)
For each $s\in\{1,\ldots,r\}$, Proposition 2.4 yields $N_{s}.\mathfrak{a}_{G}\subseteq N_{1}\oplus\ldots\oplus N_{s}$, because the $\mathbb{Z}_{p}$-lattice $N_{1}\oplus\ldots\oplus N_{s}$ is $G$-invariant. Recalling $\vartheta_{1}<\ldots<\vartheta_{r}$ we deduce that for every $i\in\mathbb{N}_{0}$,
$\Big{(}\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}N_{s}\Big{)}.\mathfrak{a}_{G}=\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}\big{(}N_{s}.\mathfrak{a}_{G}\big{)}\\ \subseteq\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}\big{(}N_{1}\oplus\ldots\oplus N_{s}\big{)}=\bigoplus_{s=1}^{r}p^{\lfloor i\vartheta_{s}\rfloor}N_{s}.$
Hence all terms of the series (2.3) are $G$-invariant. ∎
Lemma 2.6.
Let $G$ be a finitely generated pro-$p$ group, and let $L,\tilde{L}$ be $\mathbb{Z}_{p}$-lattices equipped with $G$-actions such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L\cong\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}\tilde{L}$ as $\mathbb{Q}_{p}G$-modules. Put $d=\dim_{\mathbb{Z}_{p}}L=\dim_{\mathbb{Z}_{p}}\tilde{L}$. If the lower $p$-series of $L$ admits a stratification with growth rate $(\xi_{1},\ldots,\xi_{d})$ then so does the lower $p$-series of $\tilde{L}$.
Proof.
Without loss of generality, $L$ and $\tilde{L}$ are $G$-invariant full $\mathbb{Z}_{p}$-lattices in one and the same $\mathbb{Q}_{p}G$-module $V=\mathbb{Q}_{p}L=\mathbb{Q}_{p}\tilde{L}$ and $d=\dim_{\mathbb{Q}_{p}}V$. Then there exists $c_{1}\in\mathbb{N}$ such that $\tilde{L}\subseteq p^{-c_{1}}L$. Since $p^{-c_{1}}L$ and $L$ are isomorphic as $\mathbb{Z}_{p}G$-modules, we may assume without loss of generality that $\tilde{L}\subseteq L$.
Furthermore, there is $c_{2}\in\mathbb{N}$ such that $p^{c_{2}}L\subseteq\tilde{L}$. This implies that $p^{c_{2}}\lambda_{i}(L)\subseteq\lambda_{i}(\tilde{L})\subseteq\lambda_{i}(L)$ for all $i\in\mathbb{N}$. Let $(x_{1},\ldots,x_{d})$ be the frame for a stratification of the lower $p$-series of $L$, with growth rate $(\xi_{1},\ldots,\xi_{d})$. Then there is $c_{3}\in\mathbb{N}$ such that, for $i\in\mathbb{N}$,
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor+c_{3}}\mathbb{Z}_{p}x_{k}\subseteq\lambda_{i}(L)\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c_{3}}\mathbb{Z}_{p}x_{k}.$
Below we produce a $\mathbb{Z}_{p}$-basis $y_{1},\ldots,y_{d}$ for $\tilde{L}$ and $c_{4}\in\mathbb{N}$ such that
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor+c_{4}}\mathbb{Z}_{p}x_{k}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}.$ (2.4)
With $c=c_{2}+c_{3}+c_{4}$, this yields for all $i\in\mathbb{N}_{0}$,
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor+c}\mathbb{Z}_{p}y_{k}\subseteq\lambda_{i}(\tilde{L})\subseteq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor-c}\mathbb{Z}_{p}y_{k}.$
Thus the lower $p$-series of $\tilde{L}$ admits a stratification with frame $(y_{1},\ldots,y_{d})$ and growth rate $(\xi_{1},\ldots,\xi_{d})$.
It remains to establish (2.4) for suitable $y_{1},\ldots,y_{d}$ and $c_{4}$. For each $e\in\{1,\ldots,d\}$ we choose
$y_{e}=a_{e,1}x_{1}+\ldots+a_{e,e-1}x_{e-1}+p^{b_{e}}x_{e}\in\tilde{L},$
where $a_{e,1},\ldots,a_{e,e-1}\in\mathbb{Z}_{p}$ denote suitable coefficients and $b_{e}\in\{0,1,\ldots,c_{2}\}$, in such a way that $\bigoplus_{k=1}^{e}\mathbb{Z}_{p}y_{k}=\tilde{L}\cap\bigoplus_{k=1}^{e}\mathbb{Z}_{p}x_{k}$. We observe that $y_{1},\ldots,y_{d}$ do form a $\mathbb{Z}_{p}$-basis for $\tilde{L}$ and that $b_{1}+\ldots+b_{d}=\log_{p}\lvert L:\tilde{L}\rvert$. Furthermore, we see that for each $i\in\mathbb{N}$ and each $e\in\{1,\ldots,d\}$,
$p^{\lfloor i\xi_{e}\rfloor}y_{e}=\left(\sum_{k=1}^{e-1}\big{(}p^{\lfloor i\xi_{e}\rfloor-\lfloor i\xi_{k}\rfloor}a_{e,k}\big{)}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}\big{)}\right)+p^{b_{e}}\big{(}p^{\lfloor i\xi_{e}\rfloor}x_{e}\big{)}.$
This implies that
$\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\leq\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$
with
$\log_{p}\left\lvert\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}:\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}y_{k}\right\rvert=b_{1}+\ldots+b_{d}=\log_{p}\lvert L:\tilde{L}\rvert.$
Hence (2.4) holds for $c_{4}=\log_{p}\lvert L:\tilde{L}\rvert$. ∎
We are now in a position to prove Theorem 1.1.
Proof of Theorem 1.1.
Let $L$ be a $\mathbb{Z}_{p}$-lattice equipped with a $G$-action, for a finitely generated pro-$p$ group $G$. In view of Lemma 2.1, Proposition 2.4 and Corollary 2.5, it remains to prove that the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ admits a stratification with rational growth rates in $[\nicefrac{{1}}{{d}},1]$, where $d=\dim_{\mathbb{Z}_{p}}L$.
We argue by induction on $d$. Of course, the claim holds trivially if $L=\{0\}$, and we now suppose that $d\geq 1$. Let $N$ be a $\mathbb{Z}_{p}G$-submodule of $L$ such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Choose $n\in\mathbb{N}$ such that $p^{n+1}L\cap N\subseteq pN$. Then $\tilde{L}=p^{n}L+N$ is a $G$-invariant open $\mathbb{Z}_{p}$-sublattice of $L$, and $N$ is a direct summand of the $\mathbb{Z}_{p}$-lattice $\tilde{L}$. By Lemma 2.6, there is no harm in replacing $L$ with $\tilde{L}$ and we may thus suppose that $L$ itself decomposes as $L=M\oplus N$ for a $\mathbb{Z}_{p}$-sublattice $M$.
Suppose that we can arrange for $M$ to be $G$-invariant. Then $L$ is a direct sum $M\oplus N$ of $\mathbb{Z}_{p}G$-modules and its lower $p$-series decomposes accordingly: $\lambda_{i}(L)=\lambda_{i}(M)\oplus\lambda_{i}(N)$ for $i\in\mathbb{N}_{0}$. If both $\dim_{\mathbb{Z}_{p}}M$ and $\dim_{\mathbb{Z}_{p}}N$ are less than $d$, our claim follows by induction, while in the remaining case, $L=N$ and $M=\{0\}$, the proof can be finished by appealing to Proposition 2.3. Using Lemma 2.6, the same argument can be applied if $N$ admits a $G$-invariant almost complement $M^{\prime}$ in $L$, i.e. a $G$-invariant $\mathbb{Z}_{p}$-sublattice $M^{\prime}$ which complements $N$ to yield an open sublattice $M^{\prime}\oplus N$ of $L$.
Now suppose that $N$ has no almost complement in $L$. Let us fix some additional notation. We write
$L_{i}=\lambda_{i}(L),\quad i\in\mathbb{N}_{0},$
for the lower $p$-series. Put $e=\dim_{\mathbb{Z}_{p}}M$. By Proposition 2.3 there is $\eta\in[\nicefrac{{1}}{{(d-e)}},1]\cap\mathbb{Q}$ such that the lower $p$-series of $N$ is equivalent to the series $p^{\lfloor i\eta\rfloor}N$, $i\in\mathbb{N}_{0}$. By induction on the dimension, the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L/N$, which is $(L_{i}+N)/N$, $i\in\mathbb{N}$, admits a stratification, based on a frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$, with growth rates $\xi_{1}\leq\ldots\leq\xi_{e}$ in $[\nicefrac{{1}}{{e}},1]\cap\mathbb{Q}$. Set $\xi_{0}=0$ for notational convenience.
We choose representatives $x_{1},\ldots,x_{e}\in M$ for the cosets $\mathbf{x}_{1},\ldots,\mathbf{x}_{e}\in L/N$, and pick $\tilde{e}\in\{0,1,\ldots,e-1\}$ as large as possible subject to the conditions:
[…]
In fact, we arrange that among all possible choices – for $M$ among all possible almost complements for $N$ in $L$ (with a subsequent reduction to $L=M\oplus N$), for the stratification based on the frame $(\mathbf{x}_{1},\ldots,\mathbf{x}_{e})$ and for the representatives $x_{1},\ldots,x_{e}$ in $L$ – we make one that renders $\tilde{e}$ as large as possible. In addition we choose $\hat{e}\in\{\tilde{e}+1,\ldots,e\}$ maximal with $\xi_{\hat{e}}=\xi_{\tilde{e}+1}$, and for short we put
$\vartheta=\xi_{\tilde{e}+1}=\xi_{\tilde{e}+2}=\ldots=\xi_{\hat{e}}\qquad\text{and}\qquad T=\bigoplus_{k=\tilde{e}+1}^{\hat{e}}\mathbb{Z}_{p}x_{k}.$
Our motivation here is that, by Proposition 2.4, the $\mathbb{Z}_{p}$-lattice $\tilde{M}\oplus T\oplus N$ is $G$-invariant, even though $\tilde{M}\oplus T$ is not.
We proceed to establish three auxiliary claims. As on previous occasions we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret terms involving negative powers of $p$ and lying outside $L$.
Claim 1. There are elements $w_{i,k}\in N$, for $i\in\mathbb{N}_{0}$ and $\tilde{e}<k\leq e$, and $n_{i}\in\mathbb{N}$ with $n_{i}\leq\lfloor i\eta\rfloor$, for $i\in\mathbb{N}_{0}$, such that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0}.$ (2.5)
Indeed, our construction already incorporates a stratification of the lower $p$-series modulo $N$ and accordingly there exists $c_{1}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{1}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq L_{i}+N\quad\text{and}\quad p^{c_{1}}L_{i}\subseteq\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;N.$
Moreover, the $G$-invariant $\mathbb{Z}_{p}$-lattice $\tilde{M}=\bigoplus_{k=1}^{\tilde{e}}\mathbb{Z}_{p}x_{k}$ embeds isomorphically as $(\tilde{M}+N)/N$ into $L/N$. By Proposition 2.4, applied to $L/N$, there exists $c_{2}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{2}}\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq\lambda_{i}(\tilde{M})\subseteq L_{i}.$
We put $c_{3}=\max\{c_{1},c_{2}\}$ and observe that, a fortiori, for all $i\in\mathbb{N}_{0}$,
$p^{c_{3}}p^{\lfloor i\xi_{k}\rfloor}x_{k}\in L_{i}\qquad\text{for $1\leq k\leq\tilde{e}$.}$
Choose $j\in\mathbb{N}$ such that $L_{j}\subseteq p^{c_{3}}L$. Then for all $i\in\mathbb{N}_{0}$ with $i\geq j$,
$p^{c_{3}}\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq(L_{i}+N)\cap p^{c_{3}}L=L_{i}+(N\cap p^{c_{3}}L)=L_{i}+p^{c_{3}}N$
allows us to select $w_{i,\tilde{e}+1},\ldots,w_{i,e}\in N$ such that
$p^{c_{3}}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\in L_{i}\qquad\text{for $\tilde{e}<k\leq e$.}$
For $i\in\mathbb{N}_{0}$ with $i<j$ we set $w_{i,\tilde{e}+1}=\ldots=w_{i,e}=0$; small values of $i$ play an insignificant role and this simple choice has no relevance beyond the fact that the series in (2.5) is properly defined. To streamline the notation, we also set
$w_{i,1}=\ldots=w_{i,\tilde{e}}=0\qquad\text{for all $i\in\mathbb{N}_{0}$.}$ (2.6)
For $i\in\mathbb{N}_{0}$ with $i\geq j$ we deduce directly from our definitions that
$p^{c_{3}}\Big{(}\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;(L_{i}\cap N)\Big{)}\subseteq L_{i},$
and next we observe that
$p^{c_{1}}L_{i}\subseteq\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;N=\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;N$
implies
$p^{c_{1}+c_{3}}L_{i}=p^{c_{3}}\big{(}p^{c_{1}}L_{i}\big{)}\subseteq\Big{(}\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\underbrace{p^{c_{3}}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}}_{\in L_{i}}\Big{)}\;\oplus\;(L_{i}\cap N).$
We have established that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the filtration series
$\bigoplus_{k=1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;(L_{i}\cap N),\quad i\in\mathbb{N}_{0}.$
By Proposition 2.2, the filtration series $L_{i}\cap N$, $i\in\mathbb{N}_{0}$, of $N$ is equivalent to a filtration series of the form $p^{n_{i}}N$, $i\in\mathbb{N}_{0}$, and our choice of $\eta$, in connection with the lower $p$-series of $N$, yields $c_{4}\in\mathbb{N}$ such that $p^{\lfloor i\eta\rfloor+c_{4}}N\subseteq\lambda_{i}(N)\subseteq L_{i}$ for $i\in\mathbb{N}_{0}$. This allows us to arrange that $n_{i}\in\mathbb{N}$ satisfies $n_{i}\leq\lfloor i\eta\rfloor$ for all $i\in\mathbb{N}_{0}$. Recalling the notational ‘nullnummer’ (2.6), we deduce that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the filtration series in (2.5), as claimed. $\diamond$
Claim 2. There exists $b\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$L_{i}\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N.$ (2.7)
Clearly, the $\mathbb{Z}_{p}$-lattice $L$ decomposes as a direct sum
$L=\tilde{M}\oplus K\oplus N$
of its sublattices $\tilde{M}$, $K=\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}x_{k}$ and $N$. Furthermore, the lattices $\tilde{M}$ and $N$ are $G$-invariant, whereas $K$ is not.
Recall that $\mathfrak{a}_{G}\trianglelefteq\mathbb{Z}_{p}G$ denotes the $p$-augmentation ideal and yields the description $L_{i}=L.\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the lower $p$-series of $L$, and similarly for $N$. According to our set-up, there exists $c_{1}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{1}}\big{(}K.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{c_{1}}L_{i}\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor}K\oplus N\qquad\text{and}\qquad p^{c_{1}}\big{(}N.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{\lfloor i\eta\rfloor}N.$
By decomposing elements into their $\tilde{M}$-, $K$- and $N$-coordinates and tracing their ‘trajectory’ as we descend along the lower $p$-series, we can make the inclusion recorded above more precise. We choose to work modulo $\tilde{M}$ and can thus focus on the $K$- and $N$-coordinates. Of significance are the $K$-coordinates, because they can give rise to non-trivial $K$- and $N$-coordinates at the next step. We introduce a parameter $j$, which indicates that we follow the $K$-coordinates for $j$ steps, then switch to $N$-coordinates and continue to follow the latter for the remaining $i-j$ steps. Concretely, we deduce that for all $i\in\mathbb{N}_{0}$,
$K.\mathfrak{a}_{G}^{\,i}$ $\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor-c_{1}}N.\mathfrak{a}_{G}^{\,i-1-j}$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor+\lfloor(i-1-j)\eta\rfloor-2c_{1}}N$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus p^{\min\{\lfloor(i-1)\vartheta\rfloor,\lfloor(i-1)\eta\rfloor\}-2c_{1}-2}N.$
This yields, for $i\in\mathbb{N}_{0}$,
$L_{i}=\tilde{M}.\mathfrak{a}_{G}^{\,i}+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\subseteq M+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\\ \subseteq M+K.\mathfrak{a}_{G}^{\,i}+p^{\lfloor i\eta\rfloor-c_{1}}N\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-2c_{1}-3}N.$
Thus (2.7) holds for $b=2c_{1}+3$. $\diamond$
Claim 3. Let $\mathcal{V}=\big{(}p^{-b}N\big{)}^{\hat{e}-\tilde{e}}$, with $b$ as in Claim 2. Then there exists $m\in\mathbb{N}_{0}$ such that for every choice of $\mathbf{v}=(v_{\tilde{e}+1},\ldots,v_{\hat{e}})\in\mathcal{V}$, we have
$p^{m}N\subseteq\tilde{M}+\sum_{k=\tilde{e}+1}^{\hat{e}}(x_{k}+v_{k}).\mathbb{Z}_{p}G.$ (2.8)
For each $\mathbf{v}=(v_{\tilde{e}+1},\ldots,v_{\hat{e}})\in\mathcal{V}$, we think of $\tilde{T}_{\mathbf{v}}=\sum_{k=\tilde{e}+1}^{\hat{e}}\mathbb{Z}_{p}(x_{k}+v_{k})$ as a deformation of $T$ inside $M\oplus p^{-b}N\subseteq\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$. Recall that, by our choice of $\tilde{e}$, the $\mathbb{Z}_{p}$-lattice $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\subseteq L$ is not $G$-invariant, whereas $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\oplus N=\tilde{M}\oplus p^{b}T\oplus N$ is $G$-invariant by Proposition 2.4. This implies that
$\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N\neq\{0\}.$
Recall that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is a simple $\mathbb{Q}_{p}G$-module. Hence the non-zero $G$-invariant $\mathbb{Z}_{p}$-lattice $\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N$ is open in $N$ of lower level $l_{\mathbf{v}}\in\mathbb{N}_{0}$, say; compare with Proposition 2.2 and the paragraph above it.
Clearly, the function $\mathcal{V}\to\mathbb{N}_{0}$, $\mathbf{v}\mapsto l_{\mathbf{v}}$ is locally constant and, since the domain $\mathcal{V}$ is compact, the maximum $m=\max\{l_{\mathbf{v}}\mid\mathbf{v}\in\mathcal{V}\}$ exists and has the property described in (2.8). $\diamond$
Now we just need to put the pieces together. Combining the assertions of Claim 1 and Claim 2, we see that there exists $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the filtration series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0},$
where
$w_{i,k}\in p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N\quad\text{and}\quad\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b\leq n_{i}\leq\lfloor i\eta\rfloor$
for all relevant indices $i,k$.
First consider the case $\eta\leq\vartheta$. In this situation
$w_{i,k}\in p^{\lfloor i\eta\rfloor-b}N\quad\text{and}\quad\lfloor i\eta\rfloor-b\leq n_{i}\leq\lfloor i\eta\rfloor$
for all relevant indices $i,k$ and it follows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+c_{1})$-equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}p^{\lfloor i\xi_{k}\rfloor}x_{k}\;\oplus\;p^{\lfloor i\eta\rfloor}N,\quad i\in\mathbb{N}_{0},$
We choose any $\mathbb{Z}_{p}$-basis $y_{1},\ldots,y_{d-e}$ for $N$. Then $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to
$\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=1}^{d-e}p^{\lfloor i\eta\rfloor}\mathbb{Z}_{p}y_{k}$
and, if $k\in\{1,\ldots,\tilde{e}\}$ is minimal subject to $\eta\leq\xi_{k}$, this yields a stratification with frame
$(x_{1},x_{2},\ldots,x_{k-1},y_{1},y_{2},\ldots,y_{d-e},x_{k},x_{k+1},\ldots,x_{e})$
and growth rate
$(\xi_{1},\xi_{2},\ldots,\xi_{k-1},\underbrace{\eta,\eta,\ldots,\eta}_{\text{$d-e$ entries}},\xi_{k},\xi_{k+1},\ldots,\xi_{e})\in\big{(}[\nicefrac{{1}}{{\max\{e,d-e\}}},1]\cap\mathbb{Q}\big{)}^{d}.$
Finally, suppose that $\vartheta<\eta$. In this situation
$w_{i,k}=p^{\lfloor i\vartheta\rfloor}v_{i,k}\quad\text{with}\quad v_{i,k}\in p^{-b}N\qquad\text{and}\qquad\lfloor i\vartheta\rfloor-b\leq n_{i}\leq\lfloor i\eta\rfloor.$
Using Claim 3, we deduce that for $i\in\mathbb{N}_{0}$,
$L_{i}\supseteq p^{c_{1}}\Big{(}p^{\lfloor i\vartheta\rfloor}\tilde{M}+\bigoplus_{k=\tilde{e}+1}^{\hat{e}}p^{\lfloor i\vartheta\rfloor}\mathbb{Z}_{p}(x_{k}+v_{i,k})\Big{)}.\mathbb{Z}_{p}G\\ \supseteq p^{c_{1}+\lfloor i\vartheta\rfloor}\Big{(}\tilde{M}+\sum_{k=\tilde{e}+1}^{\hat{e}}(x_{k}+v_{i,k}).\mathbb{Z}_{p}G\Big{)}\supseteq p^{c_{1}+m+\lfloor i\vartheta\rfloor}N$
and hence
$p^{n_{i}}N\supseteq p^{c_{1}}L_{i}\cap N\supseteq p^{2c_{1}+m+\lfloor i\vartheta\rfloor}N.$
It follows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+2c_{1}+m)$-equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;p^{\lfloor i\vartheta\rfloor}N,\quad i\in\mathbb{N}_{0}.$
We choose any $\mathbb{Z}_{p}$-basis $y_{1},\ldots,y_{d-e}$ for $N$. Then $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to
$\bigoplus_{k=1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=1}^{d-e}p^{\lfloor i\vartheta\rfloor}\mathbb{Z}_{p}y_{k}$
and this yields a stratification with frame
$(x_{1},x_{2},\ldots,x_{\tilde{e}},y_{1},y_{2},\ldots,y_{d-e},x_{\tilde{e}+1},x_{\tilde{e}+2},\ldots,x_{e})$
and growth rate
$(\xi_{1},\xi_{2},\ldots,\xi_{\tilde{e}},\underbrace{\vartheta,\vartheta,\ldots,\vartheta}_{\text{$d-e$ entries}},\xi_{\tilde{e}+1},\xi_{\tilde{e}+2},\ldots,\xi_{e})\in\big{(}[\nicefrac{{1}}{{\max\{e,d-e\}}},1]\cap\mathbb{Q}\big{)}^{d}.\qed$
Next we use $p$-adic Lie theory to derive Theorem 1.2.
Proof of Theorem 1.2.
Being $p$-adic analytic, the pro-$p$ group $G$ has finite rank. By [5, Prop. 3.9], there exists $j\in\mathbb{N}$ such that $U=P_{j}(G)$ is uniformly powerful of rank $\mathrm{rk}(U)=\dim(G)=d$ and such that $P_{i}(G)$ is powerfully embedded in $U$ for all $i\in\mathbb{N}$ with $i\geq j$. We recall that there is an explicit isomorphism of categories translating between uniformly powerful pro-$p$ groups and powerful $\mathbb{Z}_{p}$-Lie lattices; see [5, Sec. 4 and 9] and [15]. This means that the underlying set of the pro-$p$ group $U$ can be equipped in a canonical way with the structure of a $\mathbb{Z}_{p}$-Lie lattice $L$ carrying the same topology and incorporating essentially all information about $U$. In particular, the construction is such that exponentiation in $U$ corresponds to scalar multiplication in $L$ and that the conjugation action of $G$ on $U$ translates to a continuous $\mathbb{Z}_{p}$-linear action of $G$ on $L$. In order to distinguish between the two parallel points of view, it is convenient to write $\underline{x}$ for $x\in U$ when it features as an element of the Lie lattice $L$ rather than as a group element. Using this notation, the earlier statements can be phrased as follows: the (identity) map $U\to L$, $x\mapsto\underline{x}$ is a homeomorphism; we have $\underline{x^{a}}=a\underline{x}$ for all $x\in U$, $a\in\mathbb{Z}_{p}$; and conjugation, encapsulated in the canonical homomorphism $G\to\operatorname{Inn}(G)$, induces a continuous homomorphism
$G\to\operatorname{Aut}_{\mathbb{Z}_{p}}(L),\quad g\mapsto\alpha_{g},\qquad\text{where $\underline{x}\alpha_{g}=\underline{x^{g}}$ for all $x\in U$, $g\in G$.}$
Furthermore, for every $i\in\mathbb{N}$ with $i\geq j$, the powerfully embedded subgroup $P_{i}(G)\trianglelefteq_{\mathrm{o}}U$ gives, on the Lie lattice side, a powerfully embedded Lie sublattice $L_{i^{*}}\trianglelefteq_{\mathrm{o}}L$, where we write $i^{*}=i-j$ to ease notationally a recurrent constant shift by $j$ in the index; see [15, Thm. 1.4].
Clearly, $j^{*}=0$ gives $L_{j^{*}}=L=\lambda_{j^{*}}(L)$, and we contend that, in fact, $L_{i^{*}}=\lambda_{i^{*}}(L)$ for all $i\geq j$. Let $i\in\mathbb{N}$ with $i>j$. Since $P_{i-1}(G)$ is powerful, its Frattini subgroup equals $P_{i-1}(G)^{p}$ and corresponds to the Lie sublattice $pL_{i^{*}-1}$. Multiplicative cosets of $P_{i-1}(G)^{p}$ in $P_{i-1}(G)$ are the same as additive cosets of $pL_{i^{*}-1}$ in $L_{i^{*}-1}$, and multiplication in $P_{i-1}(G)/P_{i-1}(G)^{p}$ is the same operation as addition in $L_{i^{*}-1}/pL_{i^{*}-1}$; see [5, Cor. 4.15]. Since
$\nicefrac{{\displaystyle P_{i-1}(G)}}{{\displaystyle P_{i}(G)}}\cong\nicefrac{{\left(\displaystyle\nicefrac{{\displaystyle P_{i-1}(G)}}{{\displaystyle P_{i-1}(G)^{p}}}\right)}}{{\left(\displaystyle\nicefrac{{\displaystyle P_{i}(G)}}{{\displaystyle P_{i-1}(G)^{p}}}\right)}}$
is the largest $G$-central quotient of the multiplicative group $P_{i-1}(G)/P_{i-1}(G)^{p}$, the corresponding term
$\nicefrac{{\displaystyle L_{i^{*}-1}}}{{\displaystyle L_{i^{*}}}}\cong\nicefrac{{\left(\displaystyle\nicefrac{{\displaystyle L_{i^{*}-1}}}{{\displaystyle pL_{i^{*}-1}}}\right)}}{{\left(\displaystyle\nicefrac{{\displaystyle L_{i^{*}}}}{{\displaystyle pL_{i^{*}-1}}}\right)}}$
is the largest $G$-central quotient of the additive group $L_{i^{*}-1}/pL_{i^{*}-1}$. By induction, $L_{i^{*}-1}$ equals $\lambda_{i^{*}-1}(L)$ and thus its largest $G$-central quotient of exponent $p$ is $L_{i^{*}-1}/\lambda_{i^{*}}(L)$. This implies that $L_{i^{*}}$ is indeed equal to $\lambda_{i^{*}}(L)$.
Theorem 1.1 supplies us with a stratification of the lower $p$-series $L_{i^{*}}$, $i\in\mathbb{N}$ with $i\geq j$, based on a frame $(\underline{x_{1}},\ldots,\underline{x_{d}})$ and with a growth rate $(\xi_{1},\ldots,\xi_{d})$, say, where $\nicefrac{{1}}{{d}}\leq\xi_{1}\leq\ldots\leq\xi_{d}\leq 1$ are rational numbers. This means in particular that $\underline{x_{1}},\ldots,\underline{x_{d}}$ form a $\mathbb{Z}_{p}$-basis for $L$ and that there exists $c\in\mathbb{N}_{0}$ such that, for all $i\geq j$,
$\check{M}_{i^{*}}\subseteq L_{i^{*}}\subseteq\hat{M}_{i^{*}}$
for
$\check{M}_{i^{*}}=\bigoplus_{k=1}^{d}p^{\lfloor i^{*}\xi_{k}\rfloor+c}\mathbb{Z}_{p}\underline{x_{k}}\quad\text{and}\quad\hat{M}_{i^{*}}=\bigoplus_{k=1}^{d}p^{\lfloor i^{*}\xi_{k}\rfloor-c}\mathbb{Z}_{p}\underline{x_{k}},$
where once more we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to resolve unambiguously multiplication by negative powers of $p$. By Corollary 2.5, the lattices $\check{M}_{i^{*}}$ and $\hat{M}_{i^{*}}$ are $G$-invariant for all $i\geq j$.
We consider $i\in\mathbb{N}$ with $i\geq\max\{j+2cd,6cd\}$. We see that
$[P_{i-2cd}(G),P_{i-2cd}(G)]\subseteq P_{2i-4cd}(G)\subseteq P_{i+2cd}(G).$
In conjunction with [6, Cor. 3.5], the description of the Lie bracket in terms of the group multiplication, as described in [5, Sec. 4] and recorded in [15, Equ. (2.1)] yields the corresponding result on the Lie lattice side,
$[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}}\subseteq L_{i^{*}+2cd},$
where we write $[\cdot,\cdot]_{\mathrm{Lie}}$ for the Lie bracket to distinguish it from group commutators. Similar to the simpler situation discussed above, the finite abelian group section $P_{i-2cd}(G)/P_{i+2cd}(G)$ of $G$ corresponds to a finite abelian Lie ring section $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ of the $\mathbb{Z}_{p}$-Lie lattice $L$: the multiplicative cosets are the same as the additive cosets, multiplication and addition are the same and the Lie bracket is trivial; compare [8, Thm. 4.5].
Indeed, we already observed that the Lie bracket on $L_{i^{*}-2cd}$ is trivial modulo $L_{i^{*}+2cd}$, and the Lie ring $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ is abelian. In conjunction with [15, Lem. B.2], the description of group multiplication via the Hausdorff series $\Phi$, namely
$\underline{xy}=\Phi(\underline{x},\underline{y})=\underline{x}+\underline{y}+\tfrac{1}{2}[\underline{x},\underline{y}]_{\mathrm{Lie}}+\sum_{n=3}^{\infty}u_{n}(\underline{x},\underline{y})\qquad\text{for $x,y\in U$}$
as described in [15], yields that multiplicative cosets of $P_{i+2cd}(G)$ in $P_{i-2cd}(G)$ and additive cosets of $L_{i^{*}+2cd}$ in $L_{i^{*}-2cd}$ agree with one another and that multiplication modulo $P_{i+2cd}(G)$ is the same operation as addition modulo $L_{i^{*}+2cd}$. Here the details are more tricky to work out and we indicate how to proceed when $p\geq 3$; the case $p=2$ is very similar, but would need notational refinements as in [5, 15] which we want to forgo. If $x,y\in P_{i-2cd}(G)$, then on the Lie lattice side we obtain
$u_{2}(\underline{x},\underline{y})=\tfrac{1}{2}[\underline{x},\underline{y}]_{\mathrm{Lie}}\in[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}}\subseteq L_{i^{*}+2cd}$
and, for $n\geq 3$,
$u_{n}(\underline{x},\underline{y})\in p^{-\lfloor(n-1)/(p-1)\rfloor}\big{[}\!\!\cdot\!\cdot\!\cdot\!\!\big{[}[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}},\underbrace{L_{i^{*}-2cd}]_{\mathrm{Lie}},\ldots,L_{i^{*}-2cd}}_{\text{$n-2$ terms}}\big{]}_{\mathrm{Lie}}\\ \subseteq p^{-\lfloor(n-1)/(p-1)\rfloor}p^{n-2}L_{i^{*}+2cd}\subseteq L_{i^{*}+2cd}.$
This shows that all $x,y\in P_{i-2cd}(G)$ satisfy
$\underline{xy}=\Phi(\underline{x},\underline{y})\equiv\underline{x}+\underline{y}\quad\text{modulo $L_{i^{*}+2cd}$.}$ (2.9)
In particular, the elements $x$ and $y$ are multiplicative inverses of one another modulo $P_{i+2cd}(G)$ if and only if $\underline{x}$ and $\underline{y}$ are additive inverses of one another modulo $L_{i^{*}+2cd}$. By considering inverses of inverses, we conclude that multiplicative cosets of $P_{i+2cd}(G)$ are the same as additive cosets of $L_{i^{*}+2cd}$. Now (2.9) shows that multiplication modulo $P_{i+2cd}(G)$ is the same operation as addition modulo $L_{i^{*}+2cd}$: we have established the isomorphism of abelian groups
$\nicefrac{{\displaystyle P_{i-2cd}(G)}}{{\displaystyle P_{i+2cd}(G)}}\cong\nicefrac{{\displaystyle L_{i^{*}-2cd}}}{{\displaystyle L_{i^{*}+2cd}}}\qquad\text{via}\quad xP_{i+2cd}(G)\mapsto\underline{x}+L_{i^{*}+2cd}.$
Since $G$ acts unipotently on the finite $p$-group $P_{i}(G)/P_{i}(G)^{p^{2c}}$ of order $p^{2cd}$, we conclude that $P_{i+2cd}(G)\subseteq P_{i}(G)^{p^{2c}}$ and likewise $P_{i}(G)\subseteq P_{i-2cd}(G)^{p^{2c}}$. On the Lie lattice side these inclusions translate into
$L_{i^{*}+2cd}\subseteq p^{2c}L_{i^{*}}\subseteq\check{M}_{i^{*}}\subseteq L_{i^{*}}\subseteq\hat{M}_{i^{*}}\subseteq p^{-2c}L_{i^{*}}\subseteq L_{i^{*}-2cd}.$
Recalling that $\check{M}_{i^{*}}$ and $\hat{M}_{i^{*}}$ are $G$-invariant and satisfy $p^{2c}\hat{M}_{i^{*}}=\check{M}_{i^{*}}$, we conclude that they yield open normal subgroups $\check{U}_{i},\hat{U}_{i}\trianglelefteq_{\mathrm{o}}G$ such that
$\check{U}_{i}\subseteq P_{i}(G)\subseteq\hat{U}_{i}$
and such that the abelian factor group $\hat{U}_{i}/\check{U}_{i}\cong\hat{M}_{i^{*}}/\check{M}_{i^{*}}$ is homocyclic of exponent $p^{2c}$.
It remains to factorise $\check{U}_{i},\hat{U}_{i}$ into products of procyclic subgroups. We consider $\check{U}_{i}$; the proof for $\hat{U}_{i}$ works in the same way. Since $\check{U}_{i}=\check{M}_{i^{*}}$ as sets, we certainly have $x_{k}^{\,p^{\lfloor i^{*}\xi_{k}\rfloor+c}}\in\check{U}_{i}$ for $1\leq k\leq d$. This shows that
$\check{V}_{i}=\overline{\langle x_{1}^{p^{\lfloor i^{*}\xi_{1}\rfloor+c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rangle}\subseteq_{\mathrm{c}}U$
is a subset of $\check{U}_{i}$. Furthermore, [5, Prop. 3.7 and Sec. 4.2] show that $\check{V}_{i}$ is the disjoint union of $p^{\sum_{k=1}^{d}(\lfloor i^{*}\xi_{d}\rfloor-\lfloor i^{*}\xi_{k}\rfloor)}$ cosets of the open normal subgroup
$\overline{\langle x_{1}^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rangle}\,\cdots\,\overline{\langle x_{d}^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rangle}=U^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\trianglelefteq_{\mathrm{o}}U$
of index $\lvert U:U^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\rvert=p^{(\lfloor i^{*}\xi_{d}\rfloor+c)d}$. This yields
$\mu_{U}(\check{V}_{i})=p^{\sum_{k=1}^{d}(\lfloor i^{*}\xi_{d}\rfloor-\lfloor i^{*}\xi_{k}\rfloor)}\big{\lvert}U:U^{p^{\lfloor i^{*}\xi_{d}\rfloor+c}}\big{\rvert}^{-1}=p^{-cd-\sum_{k=1}^{d}\lfloor i^{*}\xi_{k}\rfloor},$
where $\mu_{U}$ denotes the normalised Haar measure on the group $U$. We recall that the normalised Haar measures $\mu_{U}$ and $\mu_{L}$ on the compact space $U=L$ are the same; see [15, Prop. A.2] and [4, Lem. 1.9.4]. This gives
$\mu_{U}(\check{U}_{i})=\mu_{L}(\check{M}_{i^{*}})=\lvert L:\check{M}_{i^{*}}\rvert^{-1}=p^{-cd-\sum_{k=1}^{d}\lfloor i^{*}\xi_{k}\rfloor}=\mu_{U}(\check{V}_{i}),$
and hence $\check{U}_{i}=\check{V}_{i}$ as wanted. ∎
It does not take much extra effort to prove Corollaries 1.3 and 1.4.
Proof of Corollary 1.3.
We continue to use the notation established in the proof of Theorem 1.2. Let $i\in\mathbb{N}$ with $i\geq\max\{j+2cd,6cd\}$. In the special case $c=0$ everything is easier and the following argument, which implicitly uses $c\geq 1$, can be skipped. Recall that $A_{i}=\hat{U}_{i}/\check{U}_{i}$ is a homocyclic group of rank $d$ and exponent $p^{2c}$; it has a unique homocyclic subgroup of rank $d$ and exponent $p^{c}$, namely $B_{i}=A_{i}^{\,p^{c}}$, whose pre-image in $\hat{U}_{i}$ is
$G_{i}=\overline{\langle x_{1}^{\,p^{\lfloor(i-j)\xi_{1}\rfloor}},\ldots,x_{d}^{\,p^{\lfloor(i-j)\xi_{d}\rfloor}}\rangle}\trianglelefteq_{\mathrm{o}}U.$
The group $P_{i}(G)$ contains $\check{U}_{i}$ and maps, modulo $\check{U}_{i}$, onto a subgroup $C_{i}$ of $A_{i}$. Thus $C_{i}\cong\prod_{k=1}^{d}C_{p^{m_{i,k}}}$ for suitable parameters $0\leq m_{i,1}\leq\ldots\leq m_{i,d}\leq 2c$, and
$\bigl{\lvert}\log_{p}\lvert A_{i}:C_{i}\rvert-\log_{p}\lvert A_{i}:B_{i}\rvert\bigr{\rvert}\leq\sum_{k=1}^{d}\lvert m_{i,k}-c\rvert\leq cd.$
This yields
$\bigl{\lvert}\log_{p}\lvert G:P_{i}(G)\rvert-\log_{p}\lvert G:G_{i}\rvert\bigr{\rvert}\leq cd,$
and clearly
$\log_{p}\lvert G:G_{i}\rvert=\log_{p}\lvert G:P_{j}(G)\rvert+\sum_{k=1}^{d}\lfloor(i-j)\xi_{k}\rfloor.$
Thus it suffices to observe that, with the notation $\sigma=\sum_{k=1}^{d}\xi_{k}$,
$\lfloor(i-j)\sigma\rfloor-d\leq\sum_{k=1}^{d}\lfloor(i-j)\xi_{k}\rfloor\leq\lfloor(i-j)\sigma\rfloor.\qed$
Proof of Corollary 1.4.
The uniqueness of the parameters $\xi_{1},\ldots,\xi_{d}$ follows by looking at the elementary divisors of increasing sections, similar to the proof of Lemma 2.1. Indeed, suppose that $j$, $U$, $\xi_{1},\ldots,\xi_{d}$, $c$ and $\check{U}_{i}$ satisfy the assertions of Theorem 1.2. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$. By [5, Prop. 3.9], there is $j^{\prime}\in\mathbb{N}$ with $j^{\prime}\geq j$ such that, for all $i\in\mathbb{N}$ with $i\geq j^{\prime}$, the subgroups $P_{i}(G),\check{U}_{i}\trianglelefteq_{\mathrm{o}}G$ are uniformly powerful and thus give rise to full $\mathbb{Z}_{p}$-sublattices $A_{i},\check{M}_{i}\leq_{\mathrm{o}}\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$, by the Lie correspondence.
For $i_{1},i_{2}\in\mathbb{N}$ with $j^{\prime}\leq i_{1}\leq i_{2}$ the finite abelian $p$-group $A_{i_{1}}/A_{i_{2}}$ decomposes as
$A_{i_{1}}/A_{i_{2}}\cong\mathbb{Z}_{p}/p^{m_{i_{1},i_{2},1}}\mathbb{Z}_{p}\,\oplus\ldots\oplus\,\mathbb{Z}_{p}/p^{m_{i_{1},i_{2},d}}\mathbb{Z}_{p}$
with parameters $0\leq m_{i_{1},i_{2},1}\leq\ldots\leq m_{i_{1},i_{2},d}$ reflecting the elementary divisors. As in the proof of Lemma 2.1, we conclude that
$\xi_{k}=\lim_{\begin{subarray}{c}j^{\prime}\leq i_{1}\leq i_{2}\\ \text{s.t.}\ i_{2}-i_{1}\to\infty\end{subarray}}\frac{m_{i_{1},i_{2},k}}{i_{2}-i_{1}}\qquad\text{for $1\leq k\leq d$.}\qed$
3. Hausdorff spectra with respect to the lower $p$-series
Our next task is to use the stratification results from Section 2 to pin down the Hausdorff dimensions of closed subgroups in $p$-adic analytic pro-$p$ groups with respect to the lower $p$-series.
Theorem 3.1.
Let $L$ be a non-zero $\mathbb{Z}_{p}$-lattice of dimension $d=\dim_{\mathbb{Z}_{p}}(L)$. Let $\mathcal{S}$ be a filtration series of $L$ which admits a stratification with growth rate $(\xi_{1},\ldots,\xi_{d})$. Then the Hausdorff spectrum of the additive group $L$, with respect to $\mathcal{S}$, is
$\operatorname{hspec}^{\mathcal{S}}(L)=\left\{\frac{\varepsilon_{1}\xi_{1}+\ldots+\varepsilon_{d}\xi_{d}}{\xi_{1}+\ldots+\xi_{d}}\mid\varepsilon_{1},\ldots,\varepsilon_{d}\in\{0,1\}\right\}.$
In particular, this shows that $\lvert\operatorname{hspec}^{\mathcal{S}}(L)\rvert\leq 2^{d}$.
Furthermore, every subgroup $H\leq_{\mathrm{c}}L$ has strong Hausdorff dimension with respect to $\mathcal{S}$.
Proof.
We write $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and choose a frame $(x_{1},\ldots,x_{d})$ for a stratification of $\mathcal{S}$. We also recall that $0<\xi_{1}\leq\xi_{2}\leq\ldots\leq\xi_{d}$. By [17, Lem. 2.2], any two equivalent filtration series define the same Hausdorff dimension on closed subgroups of $L$. Hence we may suppose that
$L_{i}=\bigoplus_{k=1}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}\,x_{k}.$
Let $H\leq_{\mathrm{c}}L$ be a $\mathbb{Z}_{p}$-sublattice of dimension $\dim_{\mathbb{Z}_{p}}(H)=r$, say. Since $\mathbb{Z}_{p}$ is a principal ideal domain, $H$ has a $\mathbb{Z}_{p}$-basis $(y_{1},\ldots,y_{r})$ such that the parameters
$j(s)=\max\big{(}\{0\}\cup\{k\mid 1\leq k\leq d\text{ and }a_{s,k}\neq 0\}\big{)},\quad 1\leq s\leq r,$
where we write $y_{s}=\sum_{k=1}^{d}a_{s,k}x_{k}$ with suitable coefficients $a_{s,k}\in\mathbb{Z}_{p}$, satisfy $j(1)<\ldots<j(r)$. We contend that $H$ has strong Hausdorff dimension
$\operatorname{hdim}_{L}^{\mathcal{S}}(H)=\frac{\xi_{j(1)}+\ldots+\xi_{j(r)}}{\xi_{1}+\ldots+\xi_{d}}.$
If $r=0$, the lattice $H$ equals $\{0\}$ and indeed $\operatorname{hdim}_{L}^{\mathcal{S}}(H)=0$. Now suppose that $r\geq 1$ and proceed by induction. We write $H=H_{1}\oplus H_{2}$, where $H_{1}=\bigoplus_{s=1}^{r-1}\mathbb{Z}_{p}y_{s}$ and $H_{2}=\mathbb{Z}_{p}y_{r}$. By induction, we have
$\lim_{i\to\infty}\frac{\log_{p}\lvert H_{1}+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}=\operatorname{hdim}_{L}^{\mathcal{S}}(H_{1})=\frac{\xi_{j(1)}+\ldots+\xi_{j(r-1)}}{\xi_{1}+\ldots+\xi_{d}}.$ (3.1)
Now consider $H_{2}=\mathbb{Z}_{p}y_{r}$ and recall that $y_{r}=\sum_{k=1}^{j(r)}a_{r,k}x_{k}$ with $a_{r,j(r)}\neq 0$. Let $v_{p}\colon\mathbb{Z}_{p}\to\mathbb{N}_{0}\cup\{\infty\}$ denote the $p$-adic valuation map, and put
$m=\min\{v_{p}(a_{r,k})\mid 1\leq k\leq j(r)\text{ with }\xi_{k}=\xi_{j(r)}\}\in\mathbb{N}_{0}.$
For each $i\in\mathbb{N}$ the order of the finite cyclic group $(H_{2}+L_{i})/L_{i}$ equals the order of $y_{r}$ modulo $L_{i}$. Thus we see that, for all sufficiently large $i\in\mathbb{N}$,
$\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert=\max\{\lfloor i\xi_{k}\rfloor-v_{p}(a_{r,k})\mid 1\leq k\leq j(r)\}=\lfloor i\xi_{j(r)}\rfloor-m$
and thus
$\lim_{i\to\infty}\frac{\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert}{\log_{p}\lvert L:L_{i}\rvert}=\lim_{i\to\infty}\frac{\lfloor i\xi_{j(r)}\rfloor-m}{\lfloor i\xi_{1}\rfloor+\ldots+\lfloor i\xi_{d}\rfloor}=\frac{\xi_{j(r)}}{\xi_{1}+\ldots+\xi_{d}}.$ (3.2)
For each $i\in\mathbb{N}_{0}$, we see that
$\log_{p}\lvert H+L_{i}:L_{i}\rvert+\log_{p}\lvert(H_{1}+L_{i})\cap(H_{2}+L_{i}):L_{i}\rvert\\ =\log_{p}\lvert H_{1}+L_{i}:L_{i}\rvert+\log_{p}\lvert H_{2}+L_{i}:L_{i}\rvert,$
and combining this observation with (3.1) and (3.2) we arrive at
[…]
It suffices to prove that the upper limit on the right-hand side equals 00; in particular, this also shows that it is a proper limit. Since $H_{2}=\mathbb{Z}_{p}y_{r}$, we find that for each $i\in\mathbb{N}$ there exists $n_{i}\in\mathbb{N}_{0}$ such that
$(H_{1}+L_{i})\cap(H_{2}+L_{i})=p^{n_{i}}\mathbb{Z}_{p}\,y_{r}+L_{i}.$
Moreover,
$p^{n_{i}}y_{r}\in H_{1}+L_{i}\subseteq\bigoplus_{k=1}^{j(r)-1}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=j(r)}^{d}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}$
shows that $n_{i}+v_{p}(a_{r,j(r)})\geq\lfloor i\xi_{j(r)}\rfloor$ for all $i\in\mathbb{N}_{0}$. This implies that indeed
[…]
We are now in position to prove Theorem 1.5.
Proof of Theorem 1.5.
As in the proof of Theorem 1.2 we find $j\in\mathbb{N}$ such that $U=P_{j}(G)\leq_{\mathrm{o}}G$ is uniformly powerful and forms the starting point for the analysis carried out there. In particular, there are matching filtration series
$\mathcal{S}_{U}\colon U_{i^{*}}=P_{i}(G),\;i\in\mathbb{N}_{\geq j},\quad\text{and}\quad\mathcal{S}_{L}\colon L_{i^{*}}=\lambda_{i^{*}}(L),\;i\in\mathbb{N}_{\geq j},$
of the group $U$ and of the powerful $\mathbb{Z}_{p}$-Lie lattice $L$ associated to $U$, where we write $i^{*}=i-j$ as before.
Hausdorff dimensions of subgroups in $G$ with respect to the lower $p$-series $\mathcal{L}$ are closely linked to Hausdorff dimensions of subgroups in $U$ with respect to $\mathcal{S}_{U}$. Indeed, for every $H\leq_{\mathrm{c}}G$ and every open subgroup $W\leq_{\mathrm{o}}H$ with $W\subseteq U$,
$\operatorname{hdim}_{G}^{\mathcal{L}}(H)=\operatorname{hdim}_{U}^{\mathcal{S}_{U}}(W),$
and $H$ has strong Hausdorff dimension in $G$ if and only if $W$ does in $U$.
Moreover, we claim that, if $W$ is chosen so that it is powerful (using [5, Prop. 3.9]) with corresponding Lie sublattice $M\leq_{\mathrm{o}}L$, then
$\operatorname{hdim}_{U}^{\mathcal{S}_{U}}(W)=\operatorname{hdim}_{L}^{\mathcal{S}_{L}}(M),$ (3.3)
and again one of them is strong if and only if the other one is. The proof then concludes by the application of Theorems 1.1 and 3.1.
To see (3.3), let $i\in\mathbb{N}$ with $i\geq j$ and choose $m\in\mathbb{N}$ such that $U^{p^{m}}\leq U_{i^{*}}$, and accordingly $p^{m}L\leq L_{i^{*}}$. By [5, Cor. 4.15], multiplicative cosets of $U^{p^{m}}$ in $U$ correspond to additive cosets of $p^{m}L$ in $L$. Consequently, $U_{i^{*}}$ is the union of as many $U^{p^{m}}$-cosets as $L_{i^{*}}$ is the union of $p^{m}L$-cosets, hence
$\lvert U_{i^{*}}:U^{p^{m}}\rvert=\lvert L_{i^{*}}:p^{m}L\rvert.$
Since $W$ is powerful and $U_{i^{*}}$ is powerfully embedded in $U$, the group $WU_{i^{*}}$ is powerful and corresponds to the Lie lattice $M+L_{i^{*}}$. Again we obtain
$\lvert WU_{i^{*}}:U^{p^{m}}\rvert=\lvert M+L_{i^{*}}:p^{m}L\rvert.$
In combination this gives
$\frac{\log_{p}\lvert WU_{i^{*}}:U_{i^{*}}\rvert}{\log_{p}\lvert U:U_{i^{*}}\rvert}=\frac{\log_{p}\lvert WU_{i^{*}}:U^{p^{m}}\rvert-\log_{p}\lvert U_{i^{*}}:U^{p^{m}}\rvert}{\log_{p}\lvert U:U^{p^{m}}\rvert-\log_{p}\lvert U_{i^{*}}:U^{p^{m}}\rvert}\\ =\frac{\log_{p}\lvert M+L_{i^{*}}:p^{m}L\rvert-\log_{p}\lvert L_{i^{*}}:p^{m}L\rvert}{\log_{p}\lvert L:p^{m}L\rvert-\log_{p}\lvert L_{i^{*}}:p^{m}L\rvert}=\frac{\log_{p}\lvert M+L_{i^{*}}:L_{i^{*}}\rvert}{\log_{p}\lvert L:L_{i^{*}}\rvert},$
which shows that the Hausdorff dimension of $W$ in $U$ and the one of $M$ in $L$ are given by the lower limit of the same sequence. ∎
As indicated in the introduction, it is natural to look for sharp bounds for the size of the Hausdorff spectrum $\operatorname{hspec}^{\mathcal{L}}(G)$ of a $p$-adic analytic group $G$, with respect to its lower $p$-series, in terms of the dimension $\dim(G)$.
Example 3.2.
For $m\in\mathbb{N}$, let $q_{1},q_{2},\ldots,q_{m}$ denote the first $m$ consecutive primes, starting at $q_{1}=2$. We consider the metabelian pro-$p$ groups
$G_{m}=C\ltimes(A_{1}\times\ldots\times A_{m}),$
where $C=\langle y\rangle\cong\mathbb{Z}_{p}$ is infinite procyclic and, for each $j\in\{1,\ldots,m\}$, the group $A_{j}=\langle x_{j,1},\ldots,x_{j,q_{j}}\rangle\cong\mathbb{Z}_{p}^{\,q_{j}}$ is torsion-free abelian of rank $q_{j}$ and normal in $G_{m}$, the action of $C$ on $A_{j}$ being determined by
$[x_{j,k},y]=\begin{cases}x_{j,k+1}&\text{if $1\leq k\leq q_{j}-1$,}\\ x_{j,1}^{\,p}&\text{if $k=q_{j}$.}\end{cases}$
Clearly, the groups $G_{m}$ are $p$-adic analytic. Below we determine asymptotically the size of their Hausdorff spectra, with respect to the lower $p$-series. In particular, the estimate yields
$\frac{\log_{p}\bigl{\lvert}\operatorname{hspec}^{\mathcal{L}}(G_{m})\bigr{\rvert}}{\sqrt{\dim(G_{m})}}\to\infty\qquad\text{as $m\to\infty$.}$ (3.4)
We fix $m\in\mathbb{N}$ and set $G=G_{m}$. It is straightforward to check that $G$ has dimension
$\dim(G)=1+\sum_{j=1}^{m}q_{j}.$
Moreover, its lower $p$-series admits a stratification with growth rate
$(1,\,\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}},\,\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\,\ldots,\,\underbrace{\nicefrac{{1}}{{q_{m}}},\ldots,\nicefrac{{1}}{{q_{m}}}}_{\text{$q_{m}$ entries}}).$
Theorem 1.5 readily provides an upper bound for the size of the Hausdorff spectrum, namely
$\bigl{\lvert}\operatorname{hspec}^{\mathcal{L}}(G)\bigr{\rvert}\leq 2\prod_{j=1}^{m}(q_{j}+1).$
We arrive at a lower bound as follows. For $H\leq_{\mathrm{c}}G$ of the form $H=\prod_{j=1}^{m}B_{j}$ with $B_{j}\leq A_{j}$ and $\dim(B_{j})=d_{j}$, say, a direct analysis, similar to the one in the proof of Theorem 3.1, yields
$\operatorname{hdim}^{\mathcal{L}}_{G}(H)=\frac{\nicefrac{{d_{1}}}{{q_{1}}}+\nicefrac{{d_{2}}}{{q_{2}}}+\ldots+\nicefrac{{d_{m}}}{{q_{m}}}}{m+1}.$
We observe that, upon restricting to $0\leq d_{j}<q_{j}$ for $j\in\{1,\ldots,m\}$, the formula yields
$\prod_{j=1}^{m}q_{j}$
distinct values in the Hausdorff spectrum of $G$ with respect to $\mathcal{L}$.
The Prime Number Theorem provides the asymptotic equivalences
$\log\Big{(}2\prod_{j=1}^{m}(q_{j}+1)\Big{)}\sim\log\Big{(}\prod_{j=1}^{m}q_{j}\Big{)}\sim q_{m}\quad\text{and}\quad\sum_{j=1}^{m}q_{j}\sim\frac{q_{m}^{\,2}}{2\log q_{m}},$
as $m$ tends to infinity; compare with [20, Cor. 1] (and [2] for a more careful approximation). These estimates describe the size of the Hausdorff spectrum asymptotically and imply (3.4). ∎